### v1

# Fedor V. Fomin

Abstract
The connection between Hamiltonicity and the independence numbers of graphs has been a fundamental aspect of Graph Theory since the seminal works of the 1960s. This paper presents a novel algorithmic perspective on these classical problems. Our contributions are twofold.
First, we establish that a wide array of problems in undirected graphs, encompassing problems such as Hamiltonian Path and Cycle, Path Cover, Largest Linkage, and Topological Minor Containment are fixed-parameter tractable ($\operatorname{{\sf FPT}}$) parameterized by the independence number of a graph. To the best of our knowledge, these results mark the first instances of $\operatorname{{\sf FPT}}$ problems for such parameterization.
Second, we extend the algorithmic scope of the Gallai-Milgram theorem. The original theorem by Gallai and Milgram, asserts that for a graph $G$ with the independence number $\alpha(G)$, the vertex set of $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths. We show that determining whether a graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is $\operatorname{{\sf FPT}}$ parameterized by $k$.
Notably, the independence number parameterization, which describes graph’s density, departs from the typical flow of research in parameterized complexity, which focuses on parameters describing graph’s sparsity, like treewidth or vertex cover.
1 Introduction
A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph. Such parameters include treewidth, treedepth, or minimum vertex cover size. However, the parameterized complexity of many problems needs to be understood more when it comes to the parameters describing the graph’s density, such as its independence number, that is, the size of the maximum independent set.
The initial step in exploring algorithmic properties of graphs with bounded independence numbers was taken by Ovetsky Fradkin and Seymour in their work [FS15]. They demonstrated that a generalization of the edge-disjoint path problem, which they call “fusion”, can be solved in polynomial time on digraphs with a constant independence number. The work of Ovetsky Fradkin and Seymour, as well as the follow-up research [${}^{+}$20, MSSZ23], indicates that the class of digraphs with bounded independence numbers exhibits sufficient structure for algorithmic exploitation.
In this paper, we leverage the structure of undirected graphs with bounded independence numbers to develop parameterized algorithms. Our approach is applicable to a wide range of combinatorial problems, including Hamiltonicity and path covers in undirected graphs. The interplay between Hamiltonicity and independence numbers is a fundamental theme in Graph Theory, stemming from the classic theorem of Gallai–Milgram [GM60]. Surprisingly, the computational complexity of the Graph Hamiltonicity problem remained unresolved, even for graphs with an independence number of $\alpha(G)=3$, until the recent work by Jedlicková and Kratochvíl [JK23]. Jedlicková and Kratochvíl have demonstrated that not only Graph Hamiltonicity but also more general $-\ell$-Linkage problems (which involve connecting $\ell$ pairs of vertices with vertex-disjoint paths covering all the graph’s vertices) can be solved in polynomial time when the independence number $\alpha(G)$ of a graph is constant.
The algorithms proposed by Jedlicková and Kratochvíl for Graph Hamiltonicity is an XP algorithm for parameter $\alpha(G)$. This means that on an $n$-vertex graph $G$, the running times of these algorithms are the form of $n^{f(\alpha(G))}$ for some function $f$. To the best of our knowledge, no fundamental NP-hard problem has been found to be $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$. (The only trivial exception is coloring a graph $G$ in a constant number of colors $q$ because every yes-instance has at most $q\cdot\alpha(G)$ vertices.) This leads Jedlicková and Kratochvíl in  [JK23] to an intriguing question: What is the parameterized complexity of the Hamiltonian Path problem (and many other related problems) parameterized by $\alpha(G)$?
We find the situation concerning parameterization by $\alpha(G)$ unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$. For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning. First, for an input integer $k$, there are (either exact or approximation) algorithms that either correctly decide that the treewidth of the input graph is more than $k$ or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [Bod96, Kor21, RS86]. Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$. In contrast, determining whether $\alpha(G)\geq k$ is W[1]-hard with parameter $k$ [DF99]. Moreover, there is compelling evidence that the problem does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22]. To overcome the intractability of the independence number, we design algorithms with the following properties. For a given parameter $k$, the algorithm either solves an optimization problem at hand (Hamiltonicity, Path Cover, etc.) in $\operatorname{{\sf FPT}}$ time with respect to $k$, or it correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$. Such algorithms are what Raghavan and Spinrad referred to as robust [RS03].
Our contribution. To state formally our first main result, we need the following definitions.
Definition 1 (TM-model and TM-embedding).
Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)\subseteq V(G)$, such that all vertices of $M\setminus T$ have degree $2$. We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$. Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent. We further say that for the injective mapping $f:V(H)\to V(G)$ induced by the isomorphism above, $(M,f)$ is a topological minor embedding (TM-embedding) of $H$ in $G$. The size of a TM-model $(M,T)$ and of a TM-embedding $(M,f)$ is the number of vertices in $M$.
In other words, in a TM-embedding, the vertices of $H$ are mapped to distinct vertices of $G$ such that the edges of $H$ could be mapped to internally vertex-disjoint paths of $G$ connecting the corresponding pairs of vertices of $G$. Moreover, the mapping from edges of $H$ to paths in $G$ can be “read off” from the corresponding model subgraph $M$.
We also consider a more general variant of TM-embedding. In this variant, every vertex of $H$ is assigned a list of vertices of $G$ where it could be mapped.
Definition 2 (List TM-embedding).
For graphs $H$, $G$, and a mapping $L:V(H)\to 2^{V(G)}$, we say that TM-embedding $M$, $f:V(H)\to V(G)$ is a list topological minor embedding (list TM-embedding), if for each $h\in V(H)$, $f(h)\in L(h)$.
Our first result is the theorem establishing that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
Theorem 1.
There is an algorithm that, given graphs $H,G$, a list assignment $L:V(H)\to 2^{V(G)}$ and an integer parameter $k\geq 1$, in time $2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$
  • either correctly reports that there is no list TM-embedding of $H$ in $G$, or
  • computes a TM-embedding of $H$ in $G$ of the maximum size, where the maximum is taken over all list TM-embeddings of $H$ in $G$, or
  • outputs an independent set in $G$ of size $k$.
List TM-embedding encompasses several vital problems in graph algorithms and Theorem 1 implies several corollaries.
  • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$. If this path contains $|V(G)|$ vertices, this is a Hamiltonian path. For $H=C_{3}$, we get the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$. By Theorem 1, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$. This positively answers the open question from  [JK23].
  • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths. By Gallai–Milgram’s theorem [GM60], if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$. We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$. Thus, Theorem 1 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem. For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time of $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
  • Jedlicková and Kratochvíl in  [JK23] defined the following generalization of Path Cover. In the $\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$. The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph. For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path. $\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$. Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$. Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$. By Theorem 1, $\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$, resolving another open question of Jedlicková and Kratochvíl. By playing with different graphs $H$ in the List TM-embedding and Theorem 1, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
  • Another interesting scenario is the maximization variant of the $T$-Cycle problem. In the Longest $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals. The task is to decide whether there is a cycle passing through all terminals [BHT12, Kaw08, Wah13]. We can reduce this problem to List TM-embedding by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution. For each guessed ordering, we apply Theorem 1 for $H$ being a cycle on $|T|$ vertices such that each vertex of $H$ is assigned to precisely one terminal vertex, and the cyclic ordering of $H$ corresponds to the guess. This brings us to the $f(\alpha(G)+|T|)\cdot|G|^{\mathcal{O}(1)}$ algorithm computing a cycle of maximum length containing all $T$ terminals. Similar arguments hold for a more general colored variant of the problem. In this variant, we are given a graph $G$, and some of the vertices of $G$ are assigned colors from $1$ to $L$. Then, the task is to find a longest cycle containing the maximum number of colors [${}^{+}$23]. Note that an optimal cycle can contain many vertices of the same color. By guessing the colors participating in a solution cycle $C$ and the ordering of their first appearance in some cyclic ordering of $C$, we construct a cycle $H$ on at most $L$ vertices. Every vertex of $H$ is assigned the list of vertices of $G$ corresponding to the guessed color. This brings us to the algorithm of running time $f(\alpha(G)+L)\cdot|G|^{\mathcal{O}(1)}$.
The Gallai–Milgram theorem from 1960 [GM60] states that the vertex set of every (directed or undirected) graph can be covered by at most $\alpha(G)$ vertex-disjoint paths. Known constructive proofs of the Gallai–Milgram theorem can be used to construct a graph path cover with a size of at most $\alpha(G)$. Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
One of the corollaries of Theorem 1 is that deciding whether a graph $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$, parameterized by $\alpha(G)$. It is worth noting that Theorem 2 provides another non-trivial extension of this result as we solve this problem by taking $k=\alpha(G)-1$ in the theorem.
Theorem 2.
There is an algorithm that, given an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, in time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$ outputs a path cover $\mathcal{P}$ of $G$ and, furthermore,
  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
Related Work. The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms. Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved. Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16]. This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, ${}^{+}$20].
This progress also concerns structural parameterized complexity, especially in graphs of bounded treewidth [${}^{+}$11, BCKN13, CKN13, FLPS16, BST23]. However, the success of treewidth parameterization does not extend to more general graph width parameters. For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10]. For a comprehensive overview and further details on the algorithms for finding long cycles and paths in graphs, the reader may refer to Chapters 10, 11, and 12 of the book by Cygan et al. [${}^{+}$15].
The interplay of the independence number, connectivity, Hamiltonicity, and path cover is a popular theme in Graph Theory, with a historical foundation dating back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72]. An excellent introduction to this aspect of Graph Theory can be found in the survey by Jackson and Ordaz [JO90]. It is surprising that in graph algorithms, despite tremendous success in the study of Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth, which are sparse graphs, there has been almost no progress on graphs of bounded independence number. The notable exceptions are the works of Ovetsky Fradkin and Seymour [FS15] and Jedlicková and Kratochvíl [JK23]. The first work deals with the edge-disjoint path problem, and the second, most relevant to our first theorem, focuses on Graph Hamiltonicity and $-\ell$-Linkage problems. In the end of this section, we explain the most crucial differences between the approaches taken in our work and the work of Jedlicková and Kratochvíl. Both papers by Ovetsky Fradkin and Seymour and Jedlicková and Kratochvíl provide XP algorithms that run in polynomial time when the independence number of a graph is constant.
Theorem 1 could also be treated as a contribution to the rapidly developing area of algorithms on graphs, excluding some induced subgraphs like $P_{t}$ or $C_{t}$ [${}^{+}$21, ${}^{+}$21]. (Every graph with an independence number of $k$ excludes the graph $kK_{1}$ as an induced subgraph.) However, all results in this area concern the computation of the maximum independent set or, more generally, induced subgraphs of bounded treewidth. Hamiltonicity is a more complex problem; it becomes NP-complete already on $P_{5}$-free graphs (a superclass of split graphs) [BLS99].
Theorem 2 contributes to another popular trend in parameterized algorithms that explores the algorithmic extensions of classic theorems from extremal graph theory. Examples of such works include the algorithmic extensions of Dirac’s theorem by Fomin, Golovach, Sagunov, and Simonov [FGSS20] or Hajnal–Szemerédi’s theorem by Gan, Han and Hu [GHH23].
1.1 Proof overview
Theorem 1: We explain here the main steps of the proof only for $-\ell$-linkages. An extension to arbitrary list TM-embeddings is more technical but builds on similar ideas.
Let us start by sketching the $\operatorname{{\sf X}P}$ algorithm of Jedlicková and Kratochvíl [JK23]. Their algorithm solves $\ell$-Linkage and runs in $n^{f(\alpha(G)+\ell)}$ time. Since Theorem 1 is a strict improvement of the $\operatorname{{\sf X}P}$ algorithm, we highlight crucial differences that allow us to achieve fixed-parameter tractability for the more general problem. An essential step in the algorithm of Jedlicková and Kratochvíl is the following lemma: If a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked, meaning that a path cover of size $\ell$ exists with any given endpoints of the paths.
The algorithm of [JK23] is recursive and requires $k\geq\alpha(G)$ to be given in the input. (This requirement is not a problem for an $\operatorname{{\sf X}P}$ algorithm because $k$ could be computed by a brute force in time $\mathcal{O}(n^{\alpha(G)})$.) When the input graph $G$ is $\Omega(k\cdot\ell)$-connected, the algorithm correctly reports that $G$ is $-\ell$-linked. Otherwise, when the connectivity is smaller, a minimum vertex separator $X$ of $G$ has $\mathcal{O}(k\cdot\ell)$ vertices. Every connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component. For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$. For every connected component of $G-X$ and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component. Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices. By treating each set of entry and exit points as new terminal vertices, the algorithm is called recursively on each component. By restricting the depth of the recursion by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths.
Since the approach of Jedlicková and Kratochvíl exploits small separators separating highly connected components, the first natural direction to explore is whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm. The recursive understanding technique introduced by Chitnis et al. [${}^{+}$16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al. [LRSZ18]. Unfortunately, recursive understanding is not of much help here due to the following reasons. To apply recursive understanding, one has to be able to deal with the primary case of so-called $(p,q)$-unbreakable graphs. However, for our problem, this primary case is not more straightforward than the case of general graphs.
This is why we employ a different method of decomposing the graph. To prevent the $\operatorname{{\sf X}P}$ running time, our algorithm avoids recursive calls to itself. Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either output an independent set of $G$ of size $k$ or find a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of $G-S$ is $\Omega(k\cdot(\ell+|S|))$-connected. Set $S$ breaks solution paths into at most $\ell+2|S|$ parts. Some parts belong to $G[S]$, while others belong to $G-S$. Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time. However, guessing the entry and exit points of these parts, how it is done by Jedlicková and Kratochvíl, would be too time-consuming. Here is precisely the moment when the property of the decomposition comes into play. Each connected component of $G-S$ is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked. Therefore, for any selection of the paths’ endpoints, the required collection of vertex-disjoint paths spanning all component vertices always exists (unless empty). Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of $G-S$ without enumerating the entry-exit endpoints in $G-S$. The latter task reduces to computing a maximum matching in an auxiliary graph. This way, the whole enumeration can be done in $\operatorname{{\sf FPT}}$ time.
Another complication comes from the fact that we do not require the $\alpha(G)<k$ assumption and instead aim to find a solution or an independent set of size $k$. For that, a constructive version of the high connectivity case is required. That is, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we need to be able to construct a family of $\ell$ disjoint paths with the given endpoints spanning the graph or to find an independent set of size $k$. We employ the general result of Thomas and Wollan [TW05]: In a $10\ell$-connected graph, a collection of $\ell$ vertex-disjoint paths with the given endpoints always exists. However, the proof of Thomas and Wollan [TW05] is involved, and it is not apparent whether it could be made constructive. Instead, we employ the property that our graphs are of bounded independence number. We use a Ramsey-based argument, finding either an independent set of size $k$ or a large clique, and in the latter case, we can find the desired $\ell$ paths by connecting the endpoints to the clique via Menger’s theorem.
We remark here that the relation between connectivity and linkages in graphs is well-studied. In particular, a sequence of results, including the above-mentioned work of Thomas and Wollan [TW05], guarantees the existence of $\ell$ vertex-disjoint paths between given pairs of endpoints in a graph of connectivity $\Omega(\ell)$. However, there is a lack of simple and efficient algorithms that compute vertex-disjoint paths connecting the given $\ell$ pairs of endpoints, in an $\Omega(\ell)$-connected, or even $\ell^{\Omega(1)}$-connected graph. Our contributions in that direction can be of independent interest to the reader.
Finally, for Theorem 1, we additionally generalize the argument above from $-\ell$-linkages to arbitrary list TM-embeddings. Here, once more, the non-recursive structure of the algorithm proves helpful: With the same choice of the set $S$, we enumerate all possible behaviors of the target model on $S$ and between $S$ and the connected components of $G-S$. Again, we avoid enumerating concrete terminals and connecting vertices in $G-S$, fixing only the general connection pattern to $S$. While considerably more technical, the enumeration can still be done in $\operatorname{{\sf FPT}}$ time. Then, similarly to $-\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly-connecting component (or an independent set of size $k$), with arbitrary locations of the terminals.
Theorem 2: To sketch our second main result, Theorem 2, we recall the proof of Gallai-Milgram theorem for undirected graphs. It is based on the following observation: if there are more than $\alpha(G)$ paths in a path cover of $G$, then there is an edge between two endpoints of distinct paths. Then these two paths can be joined into a single one and the size of the path cover is reduced by one. This proof gives a polynomial-time algorithm constructing a path cover of $G$ consisting of at most $\alpha(G)$ paths. Note that the algorithm does not have to know or compute $\alpha(G)$ itself: it is enough to iteratively join a pair of paths with adjacent endpoints until it is no longer possible. The algorithm can start from a trivial path cover consisting of $|V(G)|$ empty paths.
The procedure above on its own does not provide any additional relation between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ except for $p\leq\alpha(G)$. We propose a way to deal with this issue. First, we distinguish the paths in $\mathcal{P}$ into usual and special. We call a path usual if its endpoints are distinct and there is no edge between them in $G$. Otherwise, we call a path special. We argue that having usual paths in $\mathcal{P}$ is beneficial. For example, if all paths in $\mathcal{P}$ are usual, then the endpoints of paths in $\mathcal{P}$ form an independent set of size $2p$ as, otherwise, we would join two paths. In particular, this implies that $p\leq\alpha(G)/2$. Even having $k$ usual paths in $\mathcal{P}$ is enough since this yields an independent set of size $p+k$. Together with $\mathcal{P}$, this independent set is a feasible output for the algorithm of Theorem 2.
While special paths are unsuitable for constructing a large enough independent set, they are suitable for joining. A special path is either a path of length zero or one, or a cycle of length at least three (if considered together with the edge between the endpoints). Then, an edge between two arbitrary vertices of two special paths is as helpful as an edge between endpoints: it allows us to join the two paths and cover the vertices of their union by a single path. Our algorithm exploits this property of special paths and joins them exhaustively until there is no edge between special paths.
This property of special paths, in fact, gives us even more insight into the structure of $G$. Since special paths are independent of each other, and there are less than $k$ usual paths in $\mathcal{P}$, taking just one vertex from each special path constitutes an independent set of size greater than $p-k$. Taking two non-adjacent vertices instead of one from $2k$ special paths gives an independent set of size at least $p+k$, solving the problem immediately. Hence, the remaining case is when all but at most $2k$ special paths have vertex sets inducing cliques in $G$.
However, this structural insight alone is only half of the solution. To proceed further, we need another and more sophisticated way of transforming the paths in $\mathcal{P}$. Instead of joining two paths as before, this transformation takes two special paths and one usual path, so there are two edges without common endpoints connecting the special paths to the usual path. The transformation produces three paths, and at least two of these paths are usual. Hence, this transformation reduces the number of special paths and does not increase $\mathcal{P}$.
The transformation above provides the remaining structural insight for the algorithm. When this transformation and all the transformations above cannot be applied, a vertex set $S$ of size less than $k$ separates all usual paths from at least $p-4k$ special paths with vertex sets inducing cliques. In an auxiliary Lemma 4, we prove that the small size of $S$ can be exploited and irrelevant cliques can be removed from $G$ with respect to the size of an optimal path cover and the independence number. After applying this polynomial-time routine, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
The proof outline has come to the final case when $p=\mathcal{O}(k^{2})$. In this case, with Theorem 1 in hand, we aim to find the optimal path cover of $G$ as discussed previously in this section. The parameter choice in Theorem 1 is essential, and we choose it to be $p+k$. Thus, if the application of Theorem 1 does not give an optimal path cover, it gives an independent set of size $p+k$. Either way, the algorithm produces a solution feasible for Theorem 2.
We finish the outline of the proof of Theorem 2 by noting that the algorithm described above takes polynomial time, except for the application of Theorem 1.
2 Proof of Theorem 1: List TM-embedding parameterized by independence number
In this section, we prove Theorem 1. The proof consists of three main steps, deferred to their respective subsections. The first step is the proof of the stronger version of the theorem for highly connected graphs which we call Spanning Lemma. It shows that for any fixed injection $f:V(H)\to V(G)$ the desired TM-embedding respecting $f$ or an independent set can be found; moreover, the found TM-embedding spans $V(G)$. The second step is the so-called Merging Lemma, that combines the solutions from several highly-connected components, provided that the size of the separator between them is small. Finally, to obtain the proof of Theorem 1 itself, we show an iterative procedure that either finds the desired separator, or an independent set.
2.1 Spanning Lemma
We first deal with highly-connected graphs. More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high. This step follows the approach of Jedlicková and Kratochvíl [JK23, Theorem 7], who proved that for every $k>1$ and $\ell>1$, if a graph $G$ satisfies $\alpha(G)<k$ and $G$ is $\max\{k\ell,10\ell\}$-connected, then $G$ is $-\ell$-linked. Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$. A graph is called $\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
To adapt the approach of Jedlicková and Kratochvíl for our purposes, we need to address two issues. The first issue is minor; we need to slightly strengthen the connectivity requirement to simplify the case analysis and allow for possible coinciding endpoints of the paths (which is not allowed in [JK23]). The second issue is more serious—the proof of Jedlicková and Kratochvíl is not constructive, meaning that it cannot be used to construct the linkage. However, we need a procedure constructing the linkage for computing a TM-embedding of $H$. The proof of Jedlicková and Kratochvíl builds on the following result of Thomas and Wollan [TW05], which is also non-constructive.
Proposition 1 (Corollary 1.3, [TW05]).
If $G$ is $10\ell$-connected then $G$ is $\ell$-linked.
We do not know how to make the proof of Proposition 1 constructive in general. However, since the graphs of our interest have bounded independence number, we can use this property to construct linkages in highly connected graphs. The algorithm in the following lemma serves as a key starting subroutine in the proof of our result for highly-connected instances.
Lemma 1.
There is an algorithm that, given integers $k,\ell\geq 1$, a $10\ell$-connected graph $G$, and a family of vertex pairs $(s_{1},t_{1}),\ldots,(s_{\ell},t_{\ell})$ such that $s_{i}\neq t_{i}$ for each $i\in[\ell]$, in time $2^{(k+\ell)^{\mathcal{O}(k)}}+|G|^{\mathcal{O}(1)}$ outputs either an independent set of size $k$ or a family of $\ell$ internally vertex-disjoint $s_{i}t_{i}$-paths $P_{i}$ for $i\in[\ell]$, that is, the $i$-th path connects $s_{i}$ and $t_{i}$ and does not contain any other $s_{j}$ or $t_{j}$ as an internal vertex.
Proof.
Let $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ be the set of vertices in the input pairs. We call the vertices in $T$ terminals. For simplicity, our algorithm first ensures that $s_{1},t_{1},s_{2},t_{2},\ldots,s_{\ell},t_{\ell}$ are $2\ell$ distinct vertices, i.e., $|T|=2\ell$. For each vertex $v\in T$ that appears in $p>1$ input pairs, the algorithm replaces $v$ by $p$ true twins, i.e., vertices that contain the neighborhood of $v$ and are also adjacent to each other. Then, each repeating occurrence of $v$ in the input pairs is replaced by a distinct copy. This transformation results in $2\ell$ distinct terminal vertices. The requirement on $10\ell$-connectivity of the graph still holds, and the modification does not increase the maximum size of an independent set. Furthermore, given an independent set in the transformed graph, replacing each copy of a terminal vertex with the corresponding original vertex of $G$ gives an independent set in the original graph. The same is true for paths avoiding terminal vertices: a path between copies is equivalent to the path between original vertices. Therefore, the transformed input is equivalent to the original input, and the respective solutions can be transformed in polynomial time. Henceforth, the algorithm works with the transformed input, meaning that all vertices in the input pairs are distinct.
Observe that by Proposition 1, a family of vertex-disjoint $s_{i}t_{i}$-paths always exists. Hence, our task is to construct such paths. First, we show that if $G$ contains a clique of size $2\ell$, then the paths can be constructed in polynomial time.
Claim 1.
Let $C$ be a given clique of size $2\ell$ in $G$. Then a family of internally vertex-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$, where for each $i\in[\ell]$, $P_{i}$ connects $s_{i}$ with $t_{i}$, can be constructed in polynomial time.
Proof of Claim 1. Since $G$ is $2\ell$-connected, by Menger’s theorem [Men27, Die17], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$. These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [FF56] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time. Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$. Note that it is possible that one or two of these paths are trivial, i.e., $s_{i}=c_{2i-1}$ or $t_{i}=c_{2i}$ can be true. As $C$ is a clique in $G$ and $c_{2i-1}c_{2i}\in E(G)$, these paths can be concatenated, resulting in the $s_{i}t_{i}$-path $P_{i}$. Since $c_{1},c_{2},\ldots,c_{2\ell}$ are pairwise distinct, any two paths among $P_{1},P_{2},\ldots,P_{\ell}$ do not share any common vertices. We obtain $\ell$ paths as required. ∎
When $G$ is of bounded size, we employ brute-force enumeration.
Claim 2.
If $G$ has less than $(k+2\ell)^{k}$ vertices, then paths $P_{1},P_{2},\ldots,P_{\ell}$ can be found in $\mathcal{O}(2^{(k+2\ell)^{2k}})$ time.
Proof of Claim 2. It is sufficient to enumerate all possible subsets of edges of $G$ contained in the desired family of paths. When a subset $S$ of edges is fixed by the algorithm, it is sufficient to check that for each $i\in[\ell]$, there is a connected component in the graph with the vertex set $V(G)$ and the edge set $S$, that is a path between $s_{i}$ and $t_{i}$. If $S$ satisfies this for each $i\in[\ell]$, the algorithm stops and reports the corresponding components as $P_{1},P_{2},\ldots,P_{\ell}$. ∎
It is left to explain how the algorithm constructs the clique $C$, given that $G$ has at least $(k+2\ell)^{k}$ vertices. The essential tool here is the classical Ramsey’s theorem [Ram30], which guarantees a large clique or independent set in a large enough graph. For completeness, we present the corresponding polynomial-time subroutine based on the Ramsey number bound given by Erdős and Szekeres in 1935 [ES09].
Claim 3.
There is a polynomial-time algorithm that, given a graph $G$ and two integers $r,s>0$ such that $|V(G)|\geq\binom{r+s-2}{r-1}$, outputs either an independent set of size $r$ or a clique of size $s$ in $G$.
Proof of Claim 3. Note that $G$ necessary has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [ES09]. The algorithm is recursive. Since $r,s>0$, $|V(G)|>1$ and $G$ has at least one vertex, denote it by $v$. If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops. For $r,s\geq 2$, consider vertex sets $A=N_{G}(v)$ and $B=V(G)\setminus\{v\}\setminus A$. Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$. If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size $s-1$. In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and $s-1>0$. If its output is a clique of size $s-1$, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$. Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required. The case of $|B|\geq\binom{r+s-3}{r-2}$ is symmetric, as $v$ can be appended to any independent set in $G[B]$. Because the values of $\binom{x+y-2}{y-1}=\binom{x+y-2}{x-1}$ for $x\leq r$ and $y\leq s$ can be computed in $\mathcal{O}(rs)$ time using the recursive formula for binomial coefficients, we obtain that the running time of the algorithm is $\mathcal{O}(n^{2})$ where $n=|V(G)|$. The proof of the claim is complete. ∎
Now, we are ready to complete the algorithm:
  • If $|V(G)|\leq(k+2\ell)^{k}$, the algorithm applies the subroutine of 2 to find the solution, outputs it and stops.
  • Otherwise, since $(k+2\ell)^{k}\geq\binom{k+2\ell-2}{k-1}$, the algorithm applies the subroutine of 3 with $r=k$ and $s=2\ell$.
  • If the output is an independent set of size $k$, the algorithm outputs it, and stops.
  • Otherwise, the output is a clique of size $2\ell$ in $G$. The algorithm uses the subroutine of 1 to find the solution, then outputs it and stops.
The correctness of the algorithm and the upper bound on its running time follow from the claims. The proof of the lemma is complete. ∎
We now proceed to present an algorithm that, for a highly connected graph $G$, a graph $H$, and a fixed mapping $f:V(H)\to V(G)$, either constructs a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or finds an independent set of size $k$ in $G$.
Lemma 2 (Spanning Lemma).
Let $H$, $G$ be graphs, such that $H$ is non-empty. Let $f:V(H)\to V(G)$ be an injective mapping, and let $k$ be an integer. Let $G$ additionally be $(\max\{k+2,10\}\cdot h)$-connected, where $h=|V(H)|+|E(H)|$. There is an algorithm with running time $2^{(h+k)^{\mathcal{O}(k)}}+|G|^{\mathcal{O}(1)}$ that computes either a subgraph $M$ of $G$ so that $(M,f)$ is a TM-embedding of $H$ in $G$ spanning $V(G)$, or an independent set of size $k$ in $G$.
Proof.
First, the algorithm reduces to the case where $H$ has no isolated vertices. Since each isolated vertex $v$ in $H$ has a fixed image $f(v)$ in $G$, the algorithm simply removes $v$ from $H$ and $f(v)$ from $G$. In the following steps of the algorithm, if a spanning TM-embedding of $H$ is produced, the algorithm extends this embedding with the removed images of isolated vertices. Such an embedding remains spanning for $G$, while all described procedures are performed in polynomial time. Note that removing an equal number of vertices from both $G$ and $H$ preserves the connectivity constraint in the statement. From now on, every vertex of $H$ has at least one incident edge, while $G$ is ($\max\{k+2,10\}\cdot|E(H)|$)-connected. In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ vertex pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$. To see that the reformulation is equivalent, put $\ell=|E(H)|$ and $(s_{i},t_{i})=(f(x_{i}),f(y_{i}))$, where $x_{i},y_{i}$ are the endpoints of the $i$-th edge of $H$. Then, given the paths $P_{i}$, it is straightforward to construct the subgraph $M$ by taking the union of the paths; $(M,f)$ then is the desired TM-embedding of $H$ in $G$. The reverse is also possible: given a TM-embedding $(M,f)$ of $H$ in $G$ spanning $V(G)$, the respective path can be found by traversing $M$ from $f(x_{i})$ to $f(y_{i})$. From now on, we do not refer to TM-embeddings, but work with the equivalent disjoint-path statement above. The first step of the algorithm is to find the initial family of paths $P_{1},P_{2},\ldots P_{\ell}$, that satisfies all constraints except for spanning $V(G)$. That is, $P_{i}$ is an $s_{i}t_{i}$-path, and for each $i\neq j$, $P_{i}$ and $P_{j}$ have no common vertices except for possibly their endpoints. For this, the algorithm invokes the algorithm of Lemma 1 as a subroutine. The call is valid since $G$ is $10\ell$-connected, therefore in time $2^{(k+\ell)^{\mathcal{O}(k)}}+|G|^{\mathcal{O}(1)}$ the initial sequence $P_{1},P_{2},\ldots,P_{\ell}$ is produced. The only step left is to ensure that the spanning constraint is satisfied by the path sequence. The algorithm achieves this incrementally. That is, the algorithm enlarges one of the paths in the sequence without violating any other constraints. If we show how this can be done exhaustively and in polynomial time, it is clearly sufficient to find the desired TM-embedding.
We describe the enlargement process, akin to the method outlined in [JK23], with a key distinction. Unlike the scenario in [JK23], we lack assurance that $G$ satisfies $\alpha(G)<k$, making the procedure potentially prone to failure. Therefore, we ensure that (a) the enlargement is completed within polynomial time and (b) if the enlargement fails, an independent set of size $k$ is identified in $G$.
The procedure works as follows: If $P_{1},P_{2},\ldots,P_{\ell}$ span all vertices of $G$, the algorithm achieves the desired outcome, reports the solution, and halts. Otherwise, let $S=\bigcup_{i=1}^{\ell}V(P_{i})$ denote the span of the path sequence, with the assumption that $S\neq V(G)$. The algorithm selects any $x\in V(G)\setminus S$. By the $(k+2)\ell$-connectivity of $G$, Menger’s theorem guarantees the existence of $\min\{(k+2)\ell,|S|\}$ paths. Each path goes from $x$ to $S$, and no two paths share any common vertex (including endpoints in $S$) but $x$. One can find such paths by utilizing the Ford-Fulkerson algorithm [FF56] to find a flow of size at most $(k+2)\ell$ in $\mathcal{O}(k\ell\cdot|G|)$ running time.
If $|S|\leq(k+2)\ell$, then two of these paths end in $s_{1}$ and $v$, where $v$ is the neighbor of $s_{1}$ in $P_{1}$. Denote these paths by $Q_{1}$ and $Q_{2}$. These two paths are internally disjoint, both starting in $x$ and ending in $s_{1}$ and $v$, respectively. The algorithm enlarges $P_{1}$ by replacing it with the path $(s_{1}Q_{1}x)\circ(xQ_{2}v)\circ(vP_{1}t_{1})$, that is, $x$ becomes embedded in $P_{1}$ between $s_{1}$ and $v$. Here and next, for a path $P$ and vertices $u$, $v$ on $P$ we denote by $uPv$ the subpath of $P$ from $u$ to $v$, and for two paths $P_{1}$ and $P_{2}$ sharing an endpoint, we denote by $P_{1}\circ P_{2}$ their concatenation. This operation strictly increases the span of the path sequence. It is left for the algorithm to process the case $(k+2)\ell<|S|$. There are precisely $(k+2)\ell$ paths that were found by the Ford-Fulkerson algorithm, so at least $k\ell$ paths end in internal vertices of $P_{1},P_{2},\ldots,P_{\ell}$. Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$. The algorithm takes arbitrary $k$ of them. Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$. For each $a\in[k]$, let $u_{a}$ be the successor of $v_{a}$ in $P_{i}$ if one follows vertices on $P_{i}$ from $s_{i}$ to $t_{i}$. If the set $\{u_{1},u_{2},\ldots,u_{k}\}$ is an independent set in $G$, then the algorithm reports this independent set and stops. Otherwise, there is an edge $u_{a}u_{b}$ in $G$ for some $a,b\in[k]$. Without loss of generality, $v_{a}$ goes before $v_{b}$ on $P_{i}$ in the order from $s_{i}$ to $t_{i}$. The algorithm then enlarges $P_{i}$ by replacing it with $(s_{i}P_{i}v_{a})\circ(v_{a}Q_{a}x)\circ(xQ_{b}v_{b})\circ(v_{b}P_{i}u_{a})\circ u_{a}u_{b}\circ(u_{b}P_{i}s_{i})$. Again, $x$ is embedded into $P_{i}$ while no original vertices of $P_{i}$ are lost, increasing the total span. The algorithm repeats the described procedure until paths $P_{1},P_{2},\ldots,P_{\ell}$ span the whole vertex set of $G$, or an independent set of size $k$ is encountered. This procedure is repeated at most $|V(G)|$ times, so the total running time of this part is polynomial in the size of the instance. The correctness of the algorithm follows from the discussion. The proof of the lemma is complete. ∎
2.2 Merging Lemma
With the highly-connected case at hand, we now move to the second key ingredient in the proof of Theorem 1. We show that if the graph is not necessarily highly-connected, but contains a small-sized vertex subset such that after its removal the remaining components are highly-connected, then we can solve the instance in $\operatorname{{\sf FPT}}$ time. Here, the parameter is the the size of the subset to remove plus the size of $H$ plus $k$. In essense, the proof boils down to reconstructing the global solution from the solutions obtained on the highly-connected components via Lemma 2, therefore we dub the result Merging Lemma, stated next.
Lemma 3 (Merging Lemma).
Let $G$ be a graph, let $S\subseteq V(G)$ be a vertex subset in $G$, and let $C_{1}$, …, $C_{t}$ be the connected components of $G-S$. Let $H$ be a graph, let $k$ be a parameter, and $L:V(H)\to 2^{V(G)}$ be a list assignment. Assume that each component $C_{i}$ is at least $\max\{k+2,10\}\cdot(3h+3s)$-connected, where $h=|V(H)|+|E(H)|$, $s=|S|$. There is an algorithm that either returns a maximum-size list TM-embedding of $H$ in $G$ respecting the list assignment $L$, or correctly reports that none exists, or returns an independent set in $G$ of size $k$. The running time of the algorithm is $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$.
Proof.
The main idea of the proof is to enumerate all possible interactions of the desired TM-model with $S$. We encapsulate the choice of the interaction in an object we call a cut descriptor, to be formally defined next. Intuitively, a cut descriptor captures the behavior of a TM-model of $H$ in $G$ in the subset $S$ and its neighborhood: Given a TM-model $(M,T)$ of $H$ in $G$, its respective cut descriptor is obtained from $M$ by dissolving all non-terminals outside of $S$, where vertices not in $S$ are represented implicitly by the index of their component. The latter part is crucial to keep the number of choices the function of the parameter. As we will see later, such an implicit representation is sufficient since Lemma 2 always provides a spanning TM-model (or an independent set) and allows to fix an arbitrary injection of the terminals.
Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$. Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $W_{i}=\{w\in W:\xi(w)\in C_{i}\}$. We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
  1. CD1
    $(M_{W},f_{W})$ is a TM-embedding of $H$ in $M_{W}$;
  2. CD2
    every $w\in W$ is in $T_{W}\cup N_{M_{W}}(S)$;
  3. CD3
    for each $w,w^{\prime}\in W$ with $\xi(w)\neq\xi(w^{\prime})$, $w$ and $w^{\prime}$ are not adjacent in $M_{W}$;
  4. CD4
    $M_{W}[S]$ is a subgraph of $G$, and for each $s\in S,h\in V(H)$ with $f_{W}(h)=s$, $s\in L(h)$;
  5. CD5
    there exists an injective assignment $\tau:W\to V(G)$ such that i) $\tau(w)\in C_{\xi(w)}$ ii) if there exists $h\in V(H)$ with $f_{W}(h)=w$, then $\tau(w)\in L(h)$ iii) $N_{M_{W}}(w)\cap S\subseteq N_{G}(\tau(w))$.
We first observe an upper bound on the number of extra vertices in a cut descriptor.
Claim 4.
If $(W,M_{W},\xi,f_{W})$ is a cut descriptor, then $|W|$ is at most $h+2s$.
Proof of Claim 4. We call a vertex $w\in W$ a terminal if $w\in T_{W}$, and a connector otherwise. From CD2, every connector in $w$ is adjacent to a vertex of $S$ in $M_{W}$. Since $f_{W}$ is injective, number of terminals in $W$ is at most $|V(H)|$, therefore it remains to bound the number of connectors in $W$ by $|E(H)|+2s$. Let $w\in W$ be a connector that is adjacent to a terminal in $S$, the number of such connectors is at most $\sum_{s\in S\cap T_{W}}|N_{M_{W}}(s)\cap W|\leq|E(H)|$, since by CD1 every $s\in S\cap T_{W}$ has degree $\deg_{H}(h)$, where $h$ is the unique pre-image of $s$ with respect to $f_{W}$, and no edge of $H$ is counted twice. Finally, the number of connectors in $W$ adjacent only to non-terminals of $S$ is at most $2s$, since every non-terminal has degree two in $M_{W}$ by CD1. ∎
We also argue that if any list TM-embedding of $H$ in $G$ exists, then there exists also a “matching” cut descriptor. For a subgraph $M$ of $G$ and a subset $C\in V(G)$, we say that $M$ hits $C$ if $M[C]$ contains an edge, $M$ touches $C$ if $M[C]$ contains no edges but at least one vertex, and $M$ avoids $C$ if $V(M)\cap C=\emptyset$.
Claim 5.
Let $(M,f)$ be a list TM-embedding of $H$ in $G$. Then there exist a cut descriptor $(W,M_{W},\xi,f_{W})$ such that $M_{W}[S]=M[S]$, and for each $i\in[t]$, $M$ hits $C_{i}$ if and only if $M_{W}$ hits $W_{i}$, touches $C_{i}$ if and only if $M_{W}$ touches $W_{i}$ and $|V(M)\cap C_{i}|=|W_{i}|$, and avoids $C_{i}$ if and only if $W_{i}=\emptyset$.
Proof of Claim 5. We construct the desired cut descriptor as follows. First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$. We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$. We set $f_{W}$ as the respective mapping of the terminals obtained from $f$. Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction. We set $\xi$ so that $W_{i}=W\cap C_{i}$ for each $i\in[t]$. We now proceed to verify the remaining properties of a cut descriptor.
For CD3, consider $w,w^{\prime}\in W$ with $\xi(w)\neq\xi(w^{\prime})$. Let $w\in W_{i},w^{\prime}\in W_{j}$ in $M_{W}$, $i\neq j$, and $w\in C_{i}$, $w^{\prime}\in C_{j}$ in $G$. Since $S$ separates $C_{i}$ and $C_{j}$ in $G$, and $M$ is a subgraph of $G$, $w$ and $w^{\prime}$ are not adjacent in $M$. By construction of $M_{W}$, only non-terminals outside of $S$ and $W$ are dissolved; every such non-terminal lies in $C_{i^{\prime}}$ together with its neighbors, for some $i^{\prime}\in[t]$, and dissolving it only changes edges between vertices of $C_{i^{\prime}}$. Therefore, $w$ and $w^{\prime}$ are also not adjacent in $M_{W}$.
We now move to CD4, by construction $M_{W}[S]=M[S]$ and is therefore a subgraph of $G[S]$. Moreover, $f\equiv f_{W}$ on $S$; let $s$ be such that $f(h)=f_{W}(h)=s$, then $s\in L(h)$ since $(M,f)$ is a list TM-embedding respecting $L$.
Finally, for CD5, the assigment $\tau$ is given directly by construction as an identity assignment on $W\subset V(G)$. For each $w\in W$, $\tau(w)\in C_{\xi(w)}$ by the construction of $\xi$. If $f_{W}(h)=w$, then $w=\tau(w)\in L(h)$ since $f(h)=w$ and $(M,f)$ is a list TM-embedding respecting $L$. By construction of $M_{W}$, the edges between $W$ and $S$ are never changed between $M$ and $M_{W}$, therefore $N_{M_{W}}(w)\cap S=N_{M}(w)\cap S$, and the latter is a subset of $N_{G}(w)$ since $M$ is a subgraph of $G$.
For the last claim of the lemma, we argue again by construction: since the dissolving operation only affects edges inside $C_{i}$, for some $i$, and dissolving a vertex in $C_{i}$ leaves $C_{i}$ with at least one edge inside, the status of the component (hit/touched/avoided) is always preserved between $M$ and $M_{W}$. Specifically, if an edge of $M$ was present in $C_{i}$, then after the dissolving, an edge is still present in $M_{W}$ in $W_{i}$. If no edge but some vertices of $M$ are present in $C_{i}$ then all these vertices are terminals or neighbors of $S$, and $|M\cap C_{i}|=|W_{i}|$, and if no vertices of $M$ are in $C_{i}$, then also $W_{i}$ is empty. ∎
Now, the algorithm proceeds as follows. First, it branches over the choice of a cut descriptor. Then, for a fixed cut descriptor, the algorithm picks an arbitrary assignment fulfilling CD5. Since the assignment fixes the vertices of $W$ in $G$, we run the algorithm of Lemma 2 on each $C_{i}$ to find a TM-model connecting the vertices of $W_{i}$ as prescribed by $M_{W}$. If any of the runs returns an independent set of size $k$, we output it and stop. Otherwise, we augment the model $(M_{W},T_{W})$ to a model $(M,T)$ of $H$ in $G$ via the returned model in each $C_{i}$. Out of all models of $H$ in $G$ obtained in the branches, the algorithm returns the one of maximum size. The above algorithm is shown in detail in Algorithm 1.
Input: Graphs $G$ and $H$, mapping $L:V(H)\to 2^{V(G)}$, integer $k$, vertex subset $S\subseteq V(G)$ splitting $G$ into components $C_{1}$, …, $C_{t}$ such that for each $i\in[t]$, $G[C_{i}]$ is $\max\{k+2,10\}(3h+3s)$-connected, where $h=|V(H)|+|E(H)|$, $s=|S|$.
Output: An independent set of size $k$ in $G$, or a list TM-embedding of $H$ in $G$ respecting $L$, or the output that no such TM-embedding exists.
1 $\mathcal{M}\leftarrow\emptyset$; for $|W|\leftarrow 0$ to $h+2s$ do
2       $W\leftarrow$vertex set of size $|W|$; foreach $M_{W}$ a graph on $S\cup W$, $\xi:W\to[t]$, $f_{W}:V(H)\to S\cup W$ do
3             verify CD1CD4, verify CD5 by a matching instance and pick arbitrary suitable $\tau$; if any of CD1CD5 fail then
4                   continue;
5             end if
6            $M\leftarrow M_{W}$, $f\leftarrow f_{W}$; for $i\leftarrow 1$ to $t$ do
7                   $T_{i}\leftarrow\{w\in W:\xi(w)=i\}$, $H_{i}=M_{W}[T_{i}]$; if $H_{i}$ contains no edges then
8                         $O\leftarrow$ empty graph over the vertex set $\tau(T_{i})$;
9                  else
10                         $O\leftarrow$ invoke Lemma 2 with $H_{i}$, $G[C_{i}]$, $\left.\tau\right|_{T_{i}}$, $k$;
11                   end if
12                  if $O$ is an independent set of size $k$ then
13                         return $O$.
14                  else
15                         $M_{i}\leftarrow O$; $M\leftarrow M$ where $M[T_{i}]$ is replaced by $M_{i}$;
16                         foreach $w\in(T_{i}\cap\operatorname{Im}f_{W})$ do
17                               $f(f_{W}^{-1}(w))\leftarrow\tau(w)$;
18
19                         end foreach
20
21                   end if
22
23             end for
24            $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
25       end foreach
26
27 end for
28if $\mathcal{M}\neq\emptyset$ then
29       return $(M,f)\in\mathcal{M}$ with $\max|M|$.
30else
31       return No solution.
32 end if
Algorithm 1 Algorithm of Lemma 3.
It remains to verify that Algorithm 1 is correct and to upper-bound its running time. The series of claims presented next verifies the correctness of the algorithm. While checking the conditions CD1CD5 for a fixed cut descriptor in Algorithm 1 is generally straightforward, we provide more details for the condition CD5 by constructing $\tau$ directly.
Claim 6.
In Algorithm 1, the condition CD5 can be verified, and a suitable assignment $\tau$ can be constructed via a call to a maximum matching algorithm on an auxiliary bipartite graph of polynomial size.
Proof of Claim 6. We construct an auxiliary bipartite graph $B$ where the two parts are $W$ and $V(G)$, and $w\in W$, $v\in V(G)$ are adjacent if setting $\tau(w)=v$ does not violate any of the conditions i–iii) of CD5. That is, $w$ and $v$ are adjacent if and only if
  • $v\notin S$ and $v\in C_{\xi(w)}$,
  • if $w\in\operatorname{Im}f_{W}$ with $f_{W}(h)=w$, then $v\in L(h)$,
  • $N_{M_{W}}\cap S\subseteq N_{G}(v)$.
Clearly, a suitable assignment $\tau$ exists if and only if a matching in $B$ that covers $W$ exists. ∎
Then, we show that invoking Lemma 2 with the given arguments is possible.
Claim 7.
The call to Lemma 2 in Algorithm 1 is valid.
Proof of Claim 7. By the condition in Algorithm 1, $H_{i}$ is not empty. The mapping $\left.\tau\right|_{T_{i}}$ is indeed an injective mapping from $T_{i}=V(H_{i})$ to $C_{i}$, where the latter is guaranteed by CD5, i). It remains to verify that $G[C_{i}]$ is $\max\{k+2,10\}\cdot(|V(H_{i})|+|E(H_{i})|)$-connected. By the statement of the lemma, $G[C_{i}]$ is $\max\{k+2,10\}\cdot(3h+3s)$-connected, so it suffices to argue that $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$.
We consider two kinds of edges in $H_{i}$ separately. First, there are at most $2|E(H)|$ edges incident to vertices in $V(H_{i})\cap T_{W}$, since for each $w\in T_{w}$, $\deg_{H_{i}}(w)\leq\deg_{M_{W}}(w)$, and $\sum_{w\in T_{W}}\deg_{M_{W}}(w)\leq 2|E(H)|$. Second, consider edges where both endpoints are connectors. By CD1CD2, each connector has degree two in $M_{W}$ and at least one neighbor in $S$. Therefore, two connectors adjacent in $H_{i}$ have in total two distinct edges into $S$. The size of the cut between $S$ and $W$ in $M_{W}$ is at most $2s+|E(H)|$, since the total degree of terminals in $S$ towards $W$ is at most $|E(H)|$, and every other vertex has degree two. Thus, the number of edges of the second type in $H_{i}$ is at most $s+|E(H)|/2$. The total number of edges in $H_{i}$ is then at most $\frac{3}{2}|E(H)|+s$.
For vertices, there are at most $|V(H)|$ vertices in $H_{i}$ that are in $T_{W}$, and every other vertex is a connector in $M_{W}$. As above, every connector has at least one neighbor in $S$, and the cut size between $S$ and $W$ is at most $2s+|E(H)|$. Therefore, the total number of vertices in $H_{i}$ is at most $|V(H)|+|E(H)|+2s$, and $|V(H_{i})|+|E(H_{i})|\leq|V(H)|+\frac{5}{2}|E(H)|+3s\leq 3h+3s$. ∎
We are now able to show that we arrive at a TM-embedding of $H_{i}$ in $G[C_{i}]$, either after invoking Lemma 2 such that the output is not an independent set or after taking the empty graph over $\tau(T_{i})$.
Claim 8.
In Algorithm 1, $M_{i}$ is such that $(M_{i},\tau|_{T_{i}})$ is a TM-embedding of $H_{i}$ in $G[C_{i}]$.
Proof of Claim 8. In case $H_{i}$ is empty, in Algorithm 1 the algorithm constructs the empty graph over $\tau(T_{i})$, which is later assigned to $M_{i}$. Clearly, this is the desired TM-embedding.
If $H_{i}$ is not empty, a call to Lemma 2 is made in Algorithm 1. By 7, the call is valid, and therefore $O$ is either an independent set of size $k$ in $G[C_{i}]$, or the subgraph $M_{i}$ in the desired TM-embedding. Since the condition statement before Algorithm 1 explicitly verifies that $O$ is not an independent set of size $k$, in Algorithm 1 the latter case holds. ∎
In case for each $i\in[t]$ we obtain the desired TM-embedding of $H_{i}$ in $G[C_{i}]$, we argue that the constructed $(M,f)$ is indeed a valid list TM-embedding of $H$ in $G$.
Claim 9.
At Algorithm 1, $(M,f)$ is a list TM-embedding of $H$ in $G$ respecting $L$.
Proof of Claim 9. The algorithm starts with the TM-model $M_{W}$ of $H$ in $M_{W}$, and replaces $M_{W}[T_{i}]$ by $M_{i}$ for each $i\in[t]$, where $M_{i}$ is the TM-model of $H_{i}$ in $G[C_{i}]$. The list condition for $M$ is fulfilled immediately since $f$ and $f_{W}$ coincide, $f_{W}$ respects $L$ on $S$ via CD4, and $\tau$ respects $L$ via CD5, ii). We now argue that $(M,f)$ is a TM-embedding of $H$ in $G$.
First, we argue that $M$ is a subgraph of $G$ by considering three types of edges in $M$. Edges with two endpoints in $S$ are not changed between $M_{W}$ and $M$, and therefore are present in $G$ since $M_{W}[S]$ is a subgraph of $G$ by CD4. Consider an edge between $s\in S$ and $v\in C_{i}$, $i\in[t]$, in $M$. By construction of $M$, there exists an edge between $s$ and $w\in W_{i}$ in $M_{W}$ with $\tau(w)=v$. By CD5, $N_{M_{W}}(w)\cap S\subseteq N_{G}(\tau(w))$, meaning that $s$ and $v=\tau(w)$ are adjacent in $G$. Finally, consider an edge of $M$ with both endpoints outside $S$. By construction, the edge is a part of $M_{i}$, for some $i\in[t]$, and is present in $G$ since $M_{i}$ is a subgraph of $G[C_{i}]$.
Second, we claim that all vertices in $V(M)\setminus\operatorname{Im}f$ have degree two, and each $v\in\operatorname{Im}f$ with $v=f(h_{v})$ has $\deg_{M}(v)=\deg_{H}(h_{v})$. Based on the above, the vertices of $S$ do not change the degree between $M_{W}$ and $M$, and $M_{W}$ is a TM-model of $H$. Consider a vertex $v\in V(M)\setminus\operatorname{Im}f\setminus S$, $v\in C_{i}$ for some $i\in[t]$. If $v\notin T_{i}$, then $v$ has degree two in $M_{i}$ since $M_{i}$ is a TM-model of $H_{i}$, and has the same neighborhood in $M$ by construction of $M$. If $v\in T_{i}$, then its degree coincides with the degree of $w$ in $M_{W}$ with $\tau(w)=v$; the degree is also two since $M_{W}$ is a TM-model of $H$ and $w\notin\operatorname{Im}f_{W}$. Finally, for a vertex $v\in\operatorname{Im}f\setminus S$, $v\in C_{i}$ for some $i\in[t]$, $v=\tau(w)$ for some $w\in T_{i}\cap\operatorname{Im}f_{W}$ and $\deg_{M_{i}}(v)=\deg_{H_{i}}(w)$ since $M_{i}$ is a TM-model of $H_{i}$. Similarly, $N_{M_{W}}(w)\cap S=N_{M}(v)\cap S$, therefore $\deg_{M}(v)=\deg_{M_{i}}(v)+|N_{M}(v)\cap S|=\deg_{H_{i}}(w)+|N_{M_{W}}(w)\cap S|=\deg_{H}(h)$, where $v=f(h),w=f_{W}(h)$, since $(M_{W},f_{W})$ is a TM-model of $H$ and all neighbors of $w$ in $M_{W}$ are either in $S$ or in $T_{i}$ by CD5.
We now argue that dissolving non-terminals in $M$ results in a graph isomorphic to $H$ via $f$. We call the graph obtained after dissolving by $H^{\prime}$ and identify $V(H^{\prime})$ with a subset of $V(M)$. For $h_{v},h_{u}\in V(H)$, we show that $f(h_{v})$, $f(h_{u})$ are adjacent in $H^{\prime}$. Denote by $v=f(h_{v})$, $u=f(h_{u})$. If $v\in S$, let $v^{\prime}=v$, otherwise $v^{\prime}\in W$ is such that $\tau(v^{\prime})=v$. We define $u^{\prime}$ analogously to $u$. By the choice of $v^{\prime},u^{\prime}$, we have that $f_{W}(h_{v})=v^{\prime}$, $f_{W}(h_{u})=u^{\prime}$. Also, since $(M_{W},f_{W})$ is a TM-embedding of $H$, $v^{\prime}$ and $u^{\prime}$ are connected by a path $P^{\prime}$ in $M_{W}$ whose all internal vertices are non-terminals.
Based on $P^{\prime}$, we construct a $vu$-path $P$ in $M$ whose all internal vertices are non-terminals. First, replace all vertices of $W$ in $P^{\prime}$ with their images given by $\tau$. Second, for each two consecutive vertices $w_{1},w_{2}$ in $P^{\prime}$, such that $w_{1},w_{2}\in W$, insert in $P$ between $\tau(w_{1})$, $\tau(w_{2})$ the $\tau(w_{1})\tau(w_{2})$-path in $M_{i}$, whose all internal vertices are not in $T_{i}$. Such a path exists because $w_{1},w_{2}\in T_{i}$ for some $i\in[t]$ by CD3, and because $M_{i}$ is a TM-model of $M_{W}[T_{i}]$ in $G[C_{i}]$. Since for any $s\in S$ and $w\in W$ that are neighbors in $M_{W}$, vertices $s$ and $\tau(w)$ are also neighbors in $G$ by CD5, iii), $P$ is indeed a $vu$-path in $G$. By construction, all internal vertices are not in $\operatorname{Im}f$. Hence $v$ and $u$ are adjacent in $H^{\prime}$. It remains to observe that $\deg_{H^{\prime}}(v)=\deg_{M}(v)=\deg_{H}(h_{v})$ for any $v,h$ with $v=f(h_{v})$. Therefore, $H^{\prime}$ is isomorphic to $H$ via $f$, completing the proof. ∎
Next, we show that trying each possible cut descriptor guarantees that the largest list TM-embedding is found.
Claim 10.
The TM-embedding $(M,f)$ returned on Algorithm 1 is of maximum size among all suitable list TM-embedding.
Proof of Claim 10. Consider the list TM-embedding $(M^{*},f^{*})$ of $H$ in $G$ of maximum size. By 5, there exists a cut descriptor $(W,M_{W},\xi,f_{W})$ with $M_{W}[S]=M^{*}[S]$, and $M^{*}$ hits (touches/avoids) $C_{i}$ if and only if $M_{W}$ hits (touches/avoids) $W_{i}$ for each $i\in[t]$. Consider the TM-embedding $(M,f)$ constructed by the algorithm at Algorithm 1 with the cut descriptor set to $(W,M_{W},\xi,f_{W})$. Note that since the algorithm reached Algorithm 1, all calls to Lemma 2 resulted in a TM-model, and $M$ is successfully constructed. By construction, $M[S]=M_{W}[S]=M^{*}[S]$. Observe also that if $M_{W}$ hits $W_{i}$, then the TM-model $M_{i}$ returned by Lemma 2 spans $C_{i}$, meaning $|M\cap C_{i}|\geq|M^{*}\cap C_{i}|$. If $M^{*}$ touches $C_{i}$, then $|M^{*}\cap C_{i}|=|W_{i}|$, and $M$ has at least as many vertices in $C_{i}$. If $M^{*}$ avoids $C_{i}$, $|M^{*}\cap C_{i}|=0$. Therefore, in each $S$, $C_{1}$, $C_{2}$, …, $C_{t}$, $M$ has at least as many vertices as $M^{*}$, meaning $|M|\geq|M^{*}|$, proving the claim. ∎
We also argue that if the algorithm reports that no suitable TM-model is found, then indeed none exists.
Claim 11.
If $\mathcal{M}$ is empty by the end of the algorithm, then no list TM-model of $H$ in $G$ respecting $L$ exists.
Proof of Claim 11. Assume a list TM-model of $H$ in $G$ exists. Then by 5, there exists also a cut descriptor. Consider the iteration of the algorithm with this cut descriptor. Since the algorithm reaches Algorithm 1, all calls to Lemma 2 resulted in a suitable TM-embedding being found, therefore the algorithm necessarily adds a TM-embedding $(M,f)$ to $\mathcal{M}$ at Algorithm 1. This contradicts the assumption that Algorithm 1 is reached. ∎
Finally, we estimate the running time. First note that all lines except Algorithm 1 and Algorithm 1 incur at most a polynomial factor. By Lemma 2, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation. This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 7. Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$. We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$. We now argue that with a fixed $f_{W}$, there are at most $(s+w)^{2(s+w)}$ choices for $M_{W}$ that satisfy CD1. Indeed, terminals are fixed by $f_{W}$, therefore, it only remains to set the edges where at least one endpoint is a non-terminal. Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total. This part can also be upper-bounded by $(s+h)^{\mathcal{O}(s+h)}=2^{(s+h)^{1+o(1)}}$. The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$. ∎
2.3 Proof of Theorem 1
We are now ready to show the main result of this section, which we restate for convenience.
See 1
Proof.
Input: Graphs $G$ and $H$, mapping $L:V(H)\to 2^{V(G)}$, integer $k$.
Output: An independent set of size $k$ in $G$, or a list TM-embedding of $H$ in $G$ respecting $L$, or the output that no such TM-embedding exists.
1 $S\leftarrow\emptyset$; for $j\leftarrow 1$ to $\infty$ do
2       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$; if $t\geq k$ then
3             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$; return $I$.
4       end if
5      $\kappa_{j}\leftarrow 3\cdot\max\{k+2,10\}\cdot(|V(H)+|E(H)|+|S|)$; if for each $i\in[t]$, $C_{i}$ is $\kappa_{j}$-connected then
6             break;
7      else
8             pick $i$ s.t. $C_{i}$ is not $\kappa_{j}$-connected; find minimum separator $S^{\prime}$ in $G[C_{i}]$; $S\leftarrow S\cup S^{\prime}$;
9       end if
10
11 end for
$O\leftarrow$ invoke Lemma 3 on $G$, $H$, $L$, $k$, $S$; return $O$.
Algorithm 2 Algorithm of Theorem 1.
We iteratively construct the vertex set $S$ suitable for Lemma 3, then invoke Lemma 3. We start by setting $S=\emptyset$, then iteratively find a separator in one of the connected components of $G-S$, and add it to $S$. This happens until the connected components of $G-S$ are all $\kappa$-connected, where $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. If, at some point, there are at least $k$ connected components, we return the independent set constructed by taking a vertex from each connected component. Otherwise, when the connectivity condition is fulfilled, we run Lemma 3 on the given input with the computed $S$, and return the output. See Algorithm 2 for a detailed algorithm description.
We now argue that the algorithm is correct. If the algorithm returns at Algorithm 2, then the output is an independent set in $G$ of size $k$. Otherwise, the algorithm eventually reaches Algorithm 2, since at every iteration of the main loop, the size of $S$ increases by at least one. It is left to verify that the call to Lemma 3 is valid. For that, we need that every connected component of $G-S$ is $\kappa$-connected, for $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$. However, this is exactly the condition at Algorithm 2, and this is the only option for the algorithm to break out of the main loop while reaching Algorithm 2. Therefore, the algorithm’s output is correct by Lemma 3.
It remains to upper-bound the running time of the algorithm. For that, it is crucial to upper-bound the size of $S$ when the algorithm reaches Algorithm 2. First, we observe that the main loop performs at most $k$ iterations before it stops either at Algorithm 2 or at Algorithm 2. Specifically, at the start of the $j$-th iteration, the number of connected components in $G-S$ is at least $j$, which implies that at iteration $k$, the loop necessarily stops at Algorithm 2. This holds since each iteration increases the number of connected components by at least one, since a separator of one of the former components is added into $S$, while the remaining components are unchanged.
We now also observe that on the $j$-th iteration, the size of $S$ increases by at most $\kappa_{j}$, since we add to $S$ the minimum separator $S^{\prime}$ of some $G[C_{i}]$, and $G[C_{i}]$ is not $\kappa_{j}$-connected. Let $s_{j}$ be the size of $S$ at the start of the $j$-th iteration, and let $\alpha=3\cdot\max\{k+2,10\}$. We show the following bound.
Claim 12.
It holds that $s_{j}\leq(\alpha+2)^{j}\cdot h$, where $h=|V(H)|+|E(H)|$.
Proof of Claim 12. We show the claim by induction. The claim holds for $j=1$ since $s_{1}=0$. Assume the claim holds for $s_{j}$, consider $s_{j+1}$:
$s_{j+1}\leq s_{j}+\kappa_{j}=s_{j}+\alpha\cdot(s_{j}+h)=(\alpha+1)\cdot s_{j}+\alpha\cdot h\\ \leq(\alpha+1)\cdot(\alpha+2)^{j}\cdot h+\alpha\cdot h\leq(\alpha+2)^{j+1}\cdot h,$
where the last inequality holds since $\alpha\cdot h\leq(\alpha+2)^{j}\cdot h$ for $j\geq 1$. ∎
Since the number of iterations is at most $k$, the size of $S$ by Algorithm 2 is then at most $s_{k}=k^{\mathcal{O}(k)}\cdot h$. We now upper-bound the running time of Algorithm 1 by the running time bound of Lemma 3, since clearly up to Algorithm 2 the algorithm takes polynomial time. By Lemma 3, the total running time is thus at most $2^{(hk^{\mathcal{O}(k)})^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}=2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$. ∎
3 Proof of Theorem 2: Path cover bellow Gallai-Milgram
In this section, we prove Theorem 2. Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths. Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$. A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$. When we combine this theorem with Theorem 1, we obtain an algorithm solving Path Cover in time time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$. Our contribution in Theorem 2 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem. We restate Theorem 2 here for the reader’s convenience.
See 2
We require two auxiliary results to prove Theorem 2. The first result is dealing with small vertex separators of $G$. It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
Lemma 4.
Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of $G-S$. Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
Proof.
The construction of $X$ is based on the following observation. We say that a path $P$ in a path cover $\mathcal{P}$ of $G$ is degenerate if $V(P)=C_{i}$ for some $i\in[t]$.
Claim 13.
In any minimum path cover $\mathcal{P}$ of $G$, at most $2|S|$ of the sets $C_{1},C_{2},\ldots,C_{t}$ contain vertices of non-degenerate paths of $\mathcal{P}$.
Proof of Claim 13. Notice that if $\mathcal{P}$ has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path by the minimality of $\mathcal{P}$. Hence, if $C_{i}$ contains a vertex of a non-degenerate path then $\mathcal{P}$ has a path $P$ with an edge with one endpoint in $S$ and the other in $C_{i}$. Each vertex in $S$ is incident to at most two edges in the paths from $\mathcal{P}$. Thus, at most $2|S|$ edges are in the path of $\mathcal{P}$ with one endpoint in $S$ and the other in one of $C_{1},C_{2},\ldots,C_{t}$. Hence, at most $2|S|$ of the sets can contain vertices of non-degenerate paths. ∎ To construct $X$, for each vertex $v\in S$, we mark at most $2|S|$ cliques from $\{C_{i}\}_{i\in[t]}$ that contain neighbors of $v$. If there are at most $2|S|$ such cliques, then we mark all of them. Otherwise, we mark arbitrary $2|S|$ cliques containing neighbors of $v$. Then $X$ is defined as the set of indices of the marked cliques. It is straightforward that $|X|\leq 2|S|^{2}$ and $X$ can be constructed in polynomial time. We claim that there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$. In other words, each unmarked clique is covered by a degenerate path. Given a path cover $\mathcal{P}$ of $G$, we say that an edge $uv$ of a path $P\in\mathcal{P}$ is bad if $u\in S$ and $v$ is a vertex of an unmarked clique. We show that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges. The proof is by contradiction. We choose $\mathcal{P}$ as a minimum path cover of $G$ that minimizes the total number of bad edges in the paths. Suppose that there is $P\in\mathcal{P}$ that contain a bad edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$. We have that $C_{i}$ is unmarked and has a neighbor of $u$. Then, by the definition of the marking procedure, there are at least $2|S|$ marked cliques with neighbors of $u$. Because $C_{i}$ contains a vertex of the non-degenerate path $P$, we obtain that there is a marked clique $C_{j}$ for some $j\in[t]$ with a neighbor $w$ of $u$ such that $C_{j}$ is covered by a degenerate path $P^{\prime}\in\mathcal{P}$ by 13. Because $C_{j}$ is a clique, we can assume without loss of generality that $w$ is an end-vertex of $P^{\prime}$. Deleting the edge $uv$ separates $P$ into two paths $Q$ and $Q^{\prime}$ having end-vertices $v$ and $u$, respectively. We use the fact that $uw\in E(G)$ and construct the path $\hat{Q}$ from $Q^{\prime}$ and $P^{\prime}$ by joining their end-vertices $u$ and $w$ by the edge $uw$. Notice that $Q$ and $\hat{Q}$ are vertex-disjoint and $V(Q)\cup V(\hat{Q})=V(P)\cup V(P^{\prime})$. This allows us to construct the path cover $\mathcal{P}^{\prime}$ by replacing $P$ and $P^{\prime}$ in $\mathcal{P}$ by $Q$ and $\hat{Q}$. Because the number of paths in $\mathcal{P}^{\prime}$ is the same as in $\mathcal{P}$, $\mathcal{P}^{\prime}$ is a minimum path cover. However, the number of bad edges in the path of $\mathcal{P}^{\prime}$ is less than the number of bad edges in the paths of $\mathcal{P}$, contradicting the choice of $\mathcal{P}$. This proves that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges. To complete the proof, we remind that if a minimum path cover has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path. Thus, there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$. ∎
Our following auxiliary result is a “lighter version” of Theorem 2. It deals with the particular case of $\operatorname{pc}(G)<k^{C}$.
Lemma 5.
There is an algorithm with running time $2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$ that, given an $n$-vertex graph $G$, integer $k\geq 0$ and a path cover of $G$ with at most $\mathcal{O}(k^{C})$ paths for constant $C\geq 1$, outputs a path cover $\mathcal{P}$ of $G$ and
  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
Proof.
We present an algorithm $\mathcal{A}$. The algorithm assumes that $G$ contains at least one edge as, otherwise, $pc(G)=n$ and the minimum-size path cover consists of trivial paths. Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$. To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$. On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching. Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial. The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 1 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$. The subroutine works in
$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}}\cdot 2^{(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
running time. If the outcome of the subroutine is an independent set $I$ in $G$ of size $p^{\prime}+k$, then the algorithm outputs $\mathcal{P}^{\prime}$ and $I$ as the solution. Since $|\mathcal{P}^{\prime}|=p^{\prime}$, $|I|\geq|\mathcal{P}^{\prime}|+k$ is satisfied, so this solution satisfies the lemma statement. If the subroutine decides that $H_{ij}$ does not have TM-model in $G$ or the maximum TM-model of $H_{ij}$ in $G$ is not spanning, then $G$ cannot be covered with $i$ paths and $\mathcal{A}$ continues to the next iteration. If the subroutine gives a spanning TM-model of $H_{ij}$ in $G$, then the minimum path cover size of $G$ equals $i$ and $\mathcal{A}$ transforms this TM-model into a path cover of $G$ with $i$ paths. Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops. Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$. The proof of the lemma is complete. ∎
We are ready to prove the main result of this section.
Proof of Theorem 2.
We describe an algorithm that starts from a trivial path cover of $G$ and then gradually transforms it. These transformations are indicated as reduction rules and are performed in polynomial time. We show that each such rule does not increase the number of paths in the path cover. Moreover, each reduction rule necessarily reduces either the total number of paths or the number of special (to be defined later) paths in the path cover. A series of consecutive applications of the reduction rules always lead the algorithm to cases when the algorithm can produce a solution satisfying theorem statement in $f(k)\cdot n^{\mathcal{O}(1)}$ time and stop. We indicate such cases and the corresponding behavior of the algorithm as solution subroutines. The algorithm runtime thus always ends with a solution subroutine run. In what follows, we present the reduction rules and solution subroutines. Each of the reduction rules reduces the number of paths in a path cover or reduces the number of some special paths. Solution subroutines output path covers and independent sets. The rules and subroutines are given in the order in which the algorithm checks their applicability. That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied. When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties. In this case, the final solution subroutine calls the algorithms of Lemmas 4 and 5.
The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 00. Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$. Let the current set of paths be $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{m}\}$. We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
The first reduction rule is straightforward and comes from the proof of the Gallai-Milgram theorem for undirected graphs.
Reduction Rule 1.
If there is an edge $uv$ in $G$ such that $u\in\{s_{i},t_{i}\}$ and $v\in\{s_{j},t_{j}\}$ and $i\neq j$, then join $P_{i}$ and $P_{j}$ into a single path via $uv$.
The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one. Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$. When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$. Special and usual paths. We call a path $P_{i}$ in the path cover special, if either $s_{i}=t_{i}$, or $P_{i}$ consists of only edge $s_{i}t_{i}$, or its endpoints, are connected by the edge $s_{i}t_{i}$. If a path $P_{i}$ is not special, we call it usual. We claim that sufficiently many usual paths give a solution to the problem.
Claim 14.
If at least $k$ paths are usual in the path cover $\mathcal{P}=P_{1},\ldots,P_{m}$ and 1 is not applicable, then $m\leq\alpha(G)-k$. Moreover, an independent set $I$ in $G$ of size $m+k$ is computable in polynomial time.
Proof of Claim 14. We show that $G$ has an independent set $I$ of size $m+k$. Initialize $I$ with $\{s_{1},s_{2},\ldots,s_{p}\}$. Since 1 is not applicable, $I$ is independent in $G$. Let $i_{1},i_{2},\ldots,i_{k}$ be $k$ distinct indices of usual paths. Since each $P_{i_{j}}$ is usual, $t_{i_{j}}\neq s_{i_{j}}$ and $s_{i_{j}}$ is not connected with $t_{i_{j}}$. Also, 1 is not applicable, so there is no edge between $t_{i_{j}}$ and any vertex in $I$, while $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}\}$ is also an independent set in $G$. Put $t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}$ in $I$. We obtain an independent set of size $m+k$ in $G$, consequently $\alpha(G)-k\geq m$. ∎ 14 yields the corresponding solution subroutine for the algorithm.
Solution Subroutine 1.
If there are at least $k$ usual paths among $\mathcal{P}=P_{1},P_{2},\ldots,P_{m}$, output $\mathcal{P}$ and $I$ as in 14.
From now on, we assume that $\mathcal{P}$ has less than $k$ usual paths. The remaining reduction rules deal primarily with special paths. The following rule reduces the total number of paths given that $G$ has an edge between any vertex of a special path and an endpoint of an arbitrary other path.
Reduction Rule 2.
If $P_{i}$ is a special path of length at least two and there is an edge in $G$ connecting some $u\in V(P_{i})$ with $s_{j}$ (or $t_{j}$) for some $j\neq i$, then
  1. 1.
    Add edge $s_{i}t_{i}$ to $P_{i}$, so it becomes a cycle;
  2. 2.
    Remove edge $uv$ from $P_{i}$, where $v$ is any neighbor of $u$ in $P_{i}$, so it becomes an $uv$-path;
  3. 3.
    Join $P_{i}$ and $P_{j}$ via $us_{j}$ (or $ut_{j}$).
2 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, so $\mathcal{P}$ remains a path cover implying that the rule is correct. The following reduction rule deals with an edge between two vertices of distinct special paths. It is demonstrated in Figure 1a.
Reduction Rule 3.
If $P_{i}$ and $P_{j}$ are special paths of length at least two for $i\neq j$, and there is an edge in $G$ connecting $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$, then
  1. 1.
    Add edge $s_{j}t_{j}$ to $P_{j}$, so it becomes a cycle;
  2. 2.
    Remove edge $u_{j}v_{j}$ from $P_{j}$, so it becomes an $u_{j}v_{j}$-path.
  3. 3.
    Apply 2 to $P_{i}$, $P_{j}$ and $u_{i}u_{j}$.
Again, 3 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, thus reducing the number of paths in $\mathcal{P}$. We claim the following.
Claim 15.
If Reduction Rules 1,2, and 3 are not applicable, then for each pair of distinct special paths $P_{i},P_{j}$ in $\mathcal{P}$, there is no edge between $V(P_{i})$ and $V(P_{j})$ in $G$.
Proof of Claim 15. Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$. If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable. Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$. If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction. Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$. In this case, 1 is applicable, leading us to the final contradiction. ∎ The claim is the basis for the next solution subroutine, which the algorithm applies next. We explain how it works by proving the following claim.
Claim 16.
If 1 and Reduction Rules 1,2, and 3 are not applicable, and there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then $m<\alpha(G)-k$. Moreover, an independent set $I$ in $G$ of size more than $m+k$ is computable in polynomial time.
Proof of Claim 16. We construct an independent set $I$ consisting of more than $m+k$ vertices. Since 1 is not applicable, there are more than $m-k$ special paths. For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$. For each other special path, put its arbitrary vertex into $I$. By 15, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎ The solution subroutine based on 16 is given below. The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
Solution Subroutine 2.
If there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then output $\mathcal{P}$ and an $I$ as in 16.
The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$. The following reduction rule necessary reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
a $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$ b $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$$s_{\ell}$$t_{\ell}$$x_{i}$$x_{j}$
Figure 1: Illustration of transformations performed in (a) 3 and (b) 4.. Paths before the transformations are horizontal. Paths after the transformations are highlighted in bold and have pairwise-distinct colors. Dashed lines represent absence of an edge. Snake lines represent paths instead of single edges.
Reduction Rule 4.
If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$ and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
  1. 1.
    Make $u_{i}$ endpoint of $P_{i}$ and $u_{j}$ endpoint of $P_{j}$ similarly to the first two steps of 2;
  2. 2.
    Break $P_{\ell}$ into three (or two) parts: the prefix part $s_{\ell}P_{\ell}x_{i}$, the suffix part $x_{j}P_{\ell}t_{\ell}$, and (if exists) the remaining middle part;
  3. 3.
    Join $P_{i}$ with the prefix part via $u_{i}x_{i}$;
  4. 4.
    Join $P_{j}$ with the suffix part via $u_{j}x_{j}$.
Figure 1b illustrates the result of application of 4. The following claim addresses its correctness.
Claim 17.
If Reduction Rules 1, 2 are not applicable, then 4 reduces the number of special paths in $\mathcal{P}$ and does not increase the number of paths in $\mathcal{P}$.
Proof of Claim 17. The second part of the claim is trivial, since 4 produces three (or two, if $x_{i}x_{j}\in E(P_{\ell})$) paths by transforming three distinct paths $P_{i},P_{j},P_{\ell}$, so $|\mathcal{P}|$ cannot increase. To prove the first part, we show that the resulting $v_{i}s_{\ell}$-path and $v_{j}t_{\ell}$-path are usual. We denote these two paths by $Q_{i}$ and $Q_{j}$ correspondingly. We give the proof only for $Q_{i}$, since the proof for $Q_{j}$ is symmetric. We first show that $|V(Q_{i})|\geq 3$. Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$. The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly. Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold. But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$. Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$. This contradiction proves $|V(Q_{i})|\geq 3$. It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$. Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable. This contradiction proves the claim. ∎
The list of reduction rules of the algorithm is exhausted. Starting from this point, the algorithm exploits the structure of the graph, assuming that no reduction rule is applicable. In what follows, we explain these structural properties. We also give the last solution subroutine of the algorithm.
Claim 18.
If none of the Reduction Rules 1-5 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$. Moreover, such a set $S$ could be constructed in polynomial time.
Proof of Claim 18. Since 14 is not applicable, there are less than $k$ usual paths in $\mathcal{P}$. We construct $S$ by taking one or zero vertices from each usual path, so $|S|<k$ holds automatically. We now explain the choice of the vertex to put in $S$ for each usual path. Consider a usual path $P_{i}\in\mathcal{P}$. We call a vertex $u\in V(P_{i})$ a connector if there is a special path $P_{j}$ and a vertex $v\in V(P_{j})$ with $uv\in E(G)$. If $P_{i}$ has exactly one connector, then we put this connector in $S$. Otherwise, i.e. when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$. We show that at least $m-2k$ special paths that are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded. Assume now that $P_{j}$ is a special path that is not isolated in $G-S$. Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$. By 15, $P_{i}$ can only be usual. Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors. As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$. Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$. Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$. ∎
We are ready to present the final solution subroutine which combines all algorithmic results given previously in this section.
Solution Subroutine 3.
Perform the following steps:
  1. 1.
    Obtain $S$ according to 18 and put $h:=\max\{0,m-4k\}$;
  2. 2.
    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in $G-S$;
  3. 3.
    Apply the algorithm of Lemma 4 to $G$, $S$ and $V(P_{1}),V(P_{2}),\ldots,V(P_{h})$ and obtain a set $X$ of at most $2k^{2}$ indices, where each $i\in X$ satisfies $i\in[h]$;
  4. 4.
    Obtain $G^{\prime}$ and $\mathcal{P}^{\prime}$ by removing $V(P_{j})$ from $G$ and $P_{j}$ from $\mathcal{P}$, for each $j$ such that $j\in[h]$ and $j\notin X$;
  5. 5.
    Apply the algorithm of Lemma 5 as a subroutine to $G^{\prime},\mathcal{P}^{\prime}$ and $k^{\prime}:=2k$ and obtain a path cover $\mathcal{S}^{\prime}$ of $G^{\prime}$ and (possibly) an independent set $I^{\prime}$ in $G^{\prime}$;
  6. 6.
    Obtain a path cover $\mathcal{S}$ of $G$ by adding all paths in $\mathcal{P}\setminus\mathcal{P}^{\prime}$ to $\mathcal{S}^{\prime}$;
  7. 7.
    If $I^{\prime}$ was obtained, remove all vertices of $S$ from $I^{\prime}$ and for each path $P_{j}\in\mathcal{P}\setminus\mathcal{P}^{\prime}$, add one vertex of $P_{j}$ to $I^{\prime}$. Denote the resulting set by $I$;
  8. 8.
    Output $\mathcal{S}$ and (if obtained) $I$.
We conclude the proof with a claim certifying the correctness of the solution subroutine. In contrast with all previous rules and subroutines, 3 requires superpolynomial (but $\operatorname{{\sf FPT}}$ in $k$) running time.
Claim 19.
If Reduction Rules 1-5 and Solution Subroutines 1,2 are not applicable, then 3 works in $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$ time and produces a feasible solution.
Proof of Claim 19. First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 18 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques). We now move on to establish the upper bound on the running time. All steps, except for applying Lemma 5, are carried out in polynomial time. We have
$|\mathcal{P}^{\prime}|=|\mathcal{P}|-h+|X|=m-h+|X|\leq 4k+2k^{2}=\mathcal{O}(k^{2}).$
By applying Lemma 5 to $G^{\prime},k^{\prime}$ and $\mathcal{P}^{\prime}$, we obtain the running time $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$. Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$. Hence it is also a path cover. On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$. Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$. First, let us consider the scenario where $I$ is not produced. This can occur only when the algorithm of Lemma 5 does not produce $I^{\prime}$. Then Lemma 5 guarantees that $\mathcal{S^{\prime}}$ is an optimal vertex cover of $G^{\prime}$ and $|\mathcal{S}^{\prime}|=\operatorname{pc}(G^{\prime})$. By Lemma 4, removal of $V(P_{j})$ for any $j\in[h]\setminus X$ decreases $\operatorname{pc}(G)$ by one, so $\operatorname{pc}(G^{\prime})=\operatorname{pc}(G)-(h-|X|).$ Since $|\mathcal{P}\setminus\mathcal{P}^{\prime}|=h-|X|$, we have $|\mathcal{S}|=\operatorname{pc}(G)$. It means that $\mathcal{S}$ is an optimal path cover of $G$.
Finally, let us consider the situation where $I$ is produced. In this case, the algorithm in Lemma 5 provides $\mathcal{S^{\prime}}$ and $I^{\prime}$ with $|I^{\prime}-|\mathcal{S}^{\prime}|=k^{\prime}=2k$. We notice that
$|I|\geq|I^{\prime}|-|S|+|\mathcal{P}\setminus\mathcal{P}^{\prime}|>|I^{\prime}|-k+(h-|X|).$
As discussed in the previous paragraph, $|\mathcal{S}|-|\mathcal{S}^{\prime}|=h-|X|$. We conclude that
$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)+(h-|X|)=k.$
The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
This completes the proof of Theorem 2. ∎
4 Conclusion
In conclusion, this paper shows that Hamiltonian Path/Cycle and Path Cover are $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$. Furthermore, Path Cover is $\operatorname{{\sf FPT}}$ parameterized “below” the Gallai–Milgram bound.
We conclude with several open research questions. All our results concern undirected graphs. The complexity of all these problems on directed graphs of small independence number remains open. For example, we do not know whether Hamiltonian Cycle is $\operatorname{{\sf NP}}$-complete on directed graphs with an independence number $k$ for some fixed $k\geq 2$. But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$. Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [${}^{+}$11, ${}^{+}$19, FGLS10, LMS18]. However, all these reductions use gadgets with large independence numbers. Obtaining algorithmic lower bounds for parameterization with small independence number appears more challenging.
A similar set of questions is about the analogue of Theorem 2 for directed graphs. The theorem of Gallai and Milgram holds for directed graphs. However, whether it could be extended algorithmically or not is not known. The “simplest” question here is whether there is a polynomial-time algorithm for computing a path cover of size $\alpha(G)-1$?

### v2

@@
 Abstract
-The connection between Hamiltonicity and the independence numbers of graphs has been a fundamental aspect of Graph Theory since the seminal works of the 1960s.
-This paper presents a novel algorithmic perspective on these classical problems.
-Our contributions are twofold.
-First, we establish that a wide array of problems in undirected graphs, encompassing problems such as Hamiltonian Path and Cycle, Path Cover, Largest Linkage, and Topological Minor Containment are fixed-parameter tractable ($\operatorname{{\sf FPT}}$) parameterized by the independence number of a graph.
-To the best of our knowledge, these results mark the first instances of $\operatorname{{\sf FPT}}$ problems for such parameterization.
-Second, we extend the algorithmic scope of the Gallai-Milgram theorem.
-The original theorem by Gallai and Milgram, asserts that for a graph $G$ with the independence number $\alpha(G)$, the vertex set of $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths.
-We show that determining whether a graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is $\operatorname{{\sf FPT}}$ parameterized by $k$.
+The classic theorem of Gallai and Milgram (1960) asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$.
+The proof of Gallai–Milgram’s theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths.
+We prove the following algorithmic extension of Gallai–Milgram’s theorem for undirected graphs: determining whether an undirected graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is fixed-parameter tractable ($\mathsf{FPT}$) when parameterized by $k$.
+More precisely, we provide an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time
+$2^{k^{O(k^{4})}}\cdot n^{O(1)},$
+and outputs a path cover $\mathcal{P}$ of $G$. Furthermore, it:
+  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover,
+  • or outputs, together with $\mathcal{P}$, an independent set of size $\lvert\mathcal{P}\rvert+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+A key subroutine in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path.
+This result is of independent interest—prior to our work, no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most 3.
+Moreover, the techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment.
+We show that all these problems are $\mathsf{FPT}$ when parameterized by the independence number of the graph.
 Notably, the independence number parameterization, which describes graph’s density, departs from the typical flow of research in parameterized complexity, which focuses on parameters describing graph’s sparsity, like treewidth or vertex cover.
 1 Introduction
-A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph.
-Such parameters include treewidth, treedepth, or minimum vertex cover size.
-However, the parameterized complexity of many problems needs to be understood more when it comes to the parameters describing the graph’s density, such as its independence number, that is, the size of the maximum independent set.
-The initial step in exploring algorithmic properties of graphs with bounded independence numbers was taken by Ovetsky Fradkin and Seymour in their work [FS15].
-They demonstrated that a generalization of the edge-disjoint path problem, which they call “fusion”, can be solved in polynomial time on digraphs with a constant independence number.
-The work of Ovetsky Fradkin and Seymour, as well as the follow-up research [${}^{+}$20, MSSZ23], indicates that the class of digraphs with bounded independence numbers exhibits sufficient structure for algorithmic exploitation.
-In this paper, we leverage the structure of undirected graphs with bounded independence numbers to develop parameterized algorithms.
-Our approach is applicable to a wide range of combinatorial problems, including Hamiltonicity and path covers in undirected graphs.
-The interplay between Hamiltonicity and independence numbers is a fundamental theme in Graph Theory, stemming from the classic theorem of Gallai–Milgram [GM60].
-Surprisingly, the computational complexity of the Graph Hamiltonicity problem remained unresolved, even for graphs with an independence number of $\alpha(G)=3$, until the recent work by Jedlicková and Kratochvíl [JK23].
-Jedlicková and Kratochvíl have demonstrated that not only Graph Hamiltonicity but also more general $-\ell$-Linkage problems (which involve connecting $\ell$ pairs of vertices with vertex-disjoint paths covering all the graph’s vertices) can be solved in polynomial time when the independence number $\alpha(G)$ of a graph is constant.
-The algorithms proposed by Jedlicková and Kratochvíl for Graph Hamiltonicity is an XP algorithm for parameter $\alpha(G)$.
-This means that on an $n$-vertex graph $G$, the running times of these algorithms are the form of $n^{f(\alpha(G))}$ for some function $f$.
-To the best of our knowledge, no fundamental NP-hard problem has been found to be $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$.
-(The only trivial exception is coloring a graph $G$ in a constant number of colors $q$ because every yes-instance has at most $q\cdot\alpha(G)$ vertices.) This leads Jedlicková and Kratochvíl in  [JK23] to an intriguing question: What is the parameterized complexity of the Hamiltonian Path problem (and many other related problems) parameterized by $\alpha(G)$?
-We find the situation concerning parameterization by $\alpha(G)$ unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$.
+The Gallai–Milgram Theorem, published by Tibor Gallai and Arthur Milgram in 1960 [GM60], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s Theorem [Dil50], concerning the covering of graphs by simpler substructures.
+In its original form, the theorem states that every (directed or undirected) graph $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the size of a maximum independent set in $G$.
+Beyond its immediate implications, the Gallai–Milgram Theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence.
+This theme can be traced back to the seminal work of Rédei [Réd34], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path.
+Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [NW71] and Chvátal–Erdős [CE72], and have continued in later works (see, e.g., [AFG83, BBVR89, Jac87, Koi94, BAH88, BT07, HJ90, Tho01, FS09]).
+In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among the Gallai–Milgram theorem, Hamiltonicity, and the independence number.
+1.1 Our contribution.
+Known constructive proofs of the Gallai–Milgram theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths.
+Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
+Theorem 1.
+There is an algorithm that, given an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, in time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$ outputs a cover of $V(G)$ by a family $\mathcal{P}$ of vertex-disjoint paths and, furthermore,
+  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
+  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+Theorem 1 aligns with a broader trend in parameterized algorithms that seeks algorithmic analogues of classic theorems in extremal graph theory.
+Examples include algorithmic extensions of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem by Gan, Han, and Hu [GHH23].
+As with Theorem 1, each of these results provides an $\operatorname{{\sf FPT}}$ algorithm parameterized “above” or “beyond” a known combinatorial bound.
+A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [DF99] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
+Consequently, we cannot simply require $\alpha(G)$ to be part of the input, which makes the existence of an $\operatorname{{\sf FPT}}$ algorithm parameterized below $\alpha(G)$ all the more surprising.
+As a special case of Theorem 1, let $k=\alpha(G)-1$.
+In this scenario, the problem reduces to determining whether all vertices of $G$ can be covered by a single path—i.e., whether $G$ has a Hamiltonian path.
+Consequently, Theorem 1 implies that deciding whether $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$ when parameterized by $\alpha(G)$.
+The construction of such an $\operatorname{{\sf FPT}}$ algorithm for Hamiltonicity is a key component of the proof of Theorem 1.
+Moreover, the techniques developed to achieve this result extend well beyond Hamiltonicity and apply to a broader class of problems.
+Similar to Theorem 1, the situation concerning parameterization by $\alpha(G)$ is unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$.
 For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning.
@@
 Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$.
-In contrast, determining whether $\alpha(G)\geq k$ is W[1]-hard with parameter $k$ [DF99].
-Moreover, there is compelling evidence that the problem does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
-To overcome the intractability of the independence number, we design algorithms with the following properties.
-For a given parameter $k$, the algorithm either solves an optimization problem at hand (Hamiltonicity, Path Cover, etc.) in $\operatorname{{\sf FPT}}$ time with respect to $k$, or it correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$.
-Such algorithms are what Raghavan and Spinrad referred to as robust [RS03].
-Our contribution. To state formally our first main result, we need the following definitions.
+In contrast, this time we deal with the parameter that does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
+Our second theorem provides an $\operatorname{{\sf FPT}}$ algorithm for computing topological minors in a graph, parameterized by its independence number.
+Let us recall that a graph $H$ is called a topological minor of a graph $G$ if there exists a subgraph of $G$ that is isomorphic to a subdivision of $H$.
+A subdivision of a graph is formed by replacing one or more edges with simple paths (i.e., inserting additional vertices along those edges).
+Thus, saying that $H$ is a topological minor of $G$ means that $G$ contains a subgraph $M$ that represents a “subdivided” version of $H$, which we refer to as a topological minor model $M$ of $H$ in $G$.
+Many interesting problems can be expressed in terms of topological minors.
+For example, the problem of computing a longest cycle in a graph is equivalent to computing the largest model $M$ of a graph $H=C_{3}$, the cycle on three vertices.
+More formally,
 Definition 1 (TM-model and TM-embedding).
-Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)\subseteq V(G)$, such that all vertices of $M\setminus T$ have degree $2$.
+Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of $M\setminus T$ have degree $2$.
 We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$.
@@
 Moreover, the mapping from edges of $H$ to paths in $G$ can be “read off” from the corresponding model subgraph $M$.
-We also consider a more general variant of TM-embedding. In this variant, every vertex of $H$ is assigned a list of vertices of $G$ where it could be mapped.
+To capture topological minor models containing specific terminal vertices, like finding a longest cycle containing a specified terminal vertex set $T$, we also consider a more general variant of TM-embedding.
+In this variant, every vertex of $H$ is assigned a list of vertices of $G$ where it could be mapped.
 Definition 2 (List TM-embedding).
 For graphs $H$, $G$, and a mapping $L:V(H)\to 2^{V(G)}$, we say that TM-embedding $M$, $f:V(H)\to V(G)$ is a list topological minor embedding (list TM-embedding), if for each $h\in V(H)$, $f(h)\in L(h)$.
-Our first result is the theorem establishing that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
+The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
+The algorithm, for a given parameter $k$, either computes a TM-embedding of maximum size in $\operatorname{{\sf FPT}}$ time with respect to $k$, or correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$.
+Such algorithms are what Raghavan and Spinrad referred to as robust [RS03].
+More precisely,
 Theorem 2.
@@
   • outputs an independent set in $G$ of size $k$.
-List TM-embedding encompasses several vital problems in graph algorithms and Theorem 1 implies several corollaries.
+To our knowledge, neither list TM-embeddings nor the Maximum List TM-Embedding problem have been studied so far.
+However, this problem is a natural generalization of several vital problems in graph algorithms.
+In particular, Theorem 2 implies the following corollaries.
   • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$.
 If this path contains $|V(G)|$ vertices, this is a Hamiltonian path.
-For $H=C_{3}$, we get the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$.
+For $H=C_{3}$, we have the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$.
 By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$.
-This positively answers the open question from  [JK23].
   • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths.
-By Gallai–Milgram’s theorem [GM60], if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$.
+By Gallai–Milgram’s theorem, if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$.
 We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$.
 Thus, Theorem 2 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem.
-For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time of $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
-  • Jedlicková and Kratochvíl in  [JK23] defined the following generalization of Path Cover.
-In the $\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$.
+Of course, it also follows from Theorem 1.
+For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
+  • Consider the following generalization of Path Cover.
+In the $-\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$.
 The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph.
 For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path.
-$\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$.
+$-\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$.
 Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$.
 Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$.
-By Theorem 1, $\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$, resolving another open question of Jedlicková and Kratochvíl.
+By Theorem 2, $-\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$.
 By playing with different graphs $H$ in the List TM-embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
   • Another interesting scenario is the maximization variant of the $T$-Cycle problem.
-In the Longest $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals.
-The task is to decide whether there is a cycle passing through all terminals [BHT12, Kaw08, Wah13].
-We can reduce this problem to List TM-embedding by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution.
+In the $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals.
+The task is to decide whether there is a cycle passing through all terminals [BHT12, Kaw08, Wah13] (in [BHT12] authors look for the minimum-length cycle).
+We can reduce the maximization variant to finding a list TM-embedding of maximum size by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution.
 For each guessed ordering, we apply Theorem 2 for $H$ being a cycle on $|T|$ vertices such that each vertex of $H$ is assigned to precisely one terminal vertex, and the cyclic ordering of $H$ corresponds to the guess.
@@
 In this variant, we are given a graph $G$, and some of the vertices of $G$ are assigned colors from $1$ to $L$.
-Then, the task is to find a longest cycle containing the maximum number of colors [${}^{+}$23].
+Then, the task is to find a longest cycle containing the maximum number of colors [FGK+23].
 Note that an optimal cycle can contain many vertices of the same color.
@@
 This brings us to the algorithm of running time $f(\alpha(G)+L)\cdot|G|^{\mathcal{O}(1)}$.
-The Gallai–Milgram theorem from 1960 [GM60] states that the vertex set of every (directed or undirected) graph can be covered by at most $\alpha(G)$ vertex-disjoint paths.
-Known constructive proofs of the Gallai–Milgram theorem can be used to construct a graph path cover with a size of at most $\alpha(G)$.
-Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
-One of the corollaries of Theorem 1 is that deciding whether a graph $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$, parameterized by $\alpha(G)$.
-It is worth noting that Theorem 2 provides another non-trivial extension of this result as we solve this problem by taking $k=\alpha(G)-1$ in the theorem.
-Theorem 2.
-There is an algorithm that, given an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, in time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$ outputs a path cover $\mathcal{P}$ of $G$ and, furthermore,
-  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
-  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
-Related Work.
-The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms.
-Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved.
-Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16].
-This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, ${}^{+}$20].
-This progress also concerns structural parameterized complexity, especially in graphs of bounded treewidth [${}^{+}$11, BCKN13, CKN13, FLPS16, BST23].
-However, the success of treewidth parameterization does not extend to more general graph width parameters.
-For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10].
-For a comprehensive overview and further details on the algorithms for finding long cycles and paths in graphs, the reader may refer to Chapters 10, 11, and 12 of the book by Cygan et al.
-[${}^{+}$15].
-The interplay of the independence number, connectivity, Hamiltonicity, and path cover is a popular theme in Graph Theory, with a historical foundation dating back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72].
-An excellent introduction to this aspect of Graph Theory can be found in the survey by Jackson and Ordaz [JO90].
-It is surprising that in graph algorithms, despite tremendous success in the study of Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth, which are sparse graphs, there has been almost no progress on graphs of bounded independence number.
-The notable exceptions are the works of Ovetsky Fradkin and Seymour [FS15] and Jedlicková and Kratochvíl [JK23].
-The first work deals with the edge-disjoint path problem, and the second, most relevant to our first theorem, focuses on Graph Hamiltonicity and $-\ell$-Linkage problems.
-In the end of this section, we explain the most crucial differences between the approaches taken in our work and the work of Jedlicková and Kratochvíl.
-Both papers by Ovetsky Fradkin and Seymour and Jedlicková and Kratochvíl provide XP algorithms that run in polynomial time when the independence number of a graph is constant.
-Theorem 1 could also be treated as a contribution to the rapidly developing area of algorithms on graphs, excluding some induced subgraphs like $P_{t}$ or $C_{t}$ [${}^{+}$21, ${}^{+}$21].
-(Every graph with an independence number of $k$ excludes the graph $kK_{1}$ as an induced subgraph.) However, all results in this area concern the computation of the maximum independent set or, more generally, induced subgraphs of bounded treewidth.
-Hamiltonicity is a more complex problem; it becomes NP-complete already on $P_{5}$-free graphs (a superclass of split graphs) [BLS99].
-Theorem 2 contributes to another popular trend in parameterized algorithms that explores the algorithmic extensions of classic theorems from extremal graph theory.
-Examples of such works include the algorithmic extensions of Dirac’s theorem by Fomin, Golovach, Sagunov, and Simonov [FGSS20] or Hajnal–Szemerédi’s theorem by Gan, Han and Hu [GHH23].
-1.1 Proof overview
-Theorem 1: We explain here the main steps of the proof only for $-\ell$-linkages. An extension to arbitrary list TM-embeddings is more technical but builds on similar ideas.
-Let us start by sketching the $\operatorname{{\sf X}P}$ algorithm of Jedlicková and Kratochvíl [JK23].
-Their algorithm solves $\ell$-Linkage and runs in $n^{f(\alpha(G)+\ell)}$ time.
-Since Theorem 1 is a strict improvement of the $\operatorname{{\sf X}P}$ algorithm, we highlight crucial differences that allow us to achieve fixed-parameter tractability for the more general problem.
-An essential step in the algorithm of Jedlicková and Kratochvíl is the following lemma: If a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked, meaning that a path cover of size $\ell$ exists with any given endpoints of the paths.
-The algorithm of [JK23] is recursive and requires $k\geq\alpha(G)$ to be given in the input.
-(This requirement is not a problem for an $\operatorname{{\sf X}P}$ algorithm because $k$ could be computed by a brute force in time $\mathcal{O}(n^{\alpha(G)})$.) When the input graph $G$ is $\Omega(k\cdot\ell)$-connected, the algorithm correctly reports that $G$ is $-\ell$-linked.
-Otherwise, when the connectivity is smaller, a minimum vertex separator $X$ of $G$ has $\mathcal{O}(k\cdot\ell)$ vertices.
-Every connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component.
+1.2 Overview of Theorems 1 and 2
+The proof of Theorem 1 uses Theorem 2 as a black box. We start with sketching the proof of Theorem 1.
+Overview of the proof of Theorem 1: To sketch the proof of Theorem 1, we first recall the standard proof of the Gallai–Milgram theorem for undirected graphs.
+This proof hinges on a simple observation: if a path cover of $G$ consists of more than $\alpha(G)$ paths, then there is an edge connecting the endpoints of two different paths.
+Merging these two paths into one reduces the cover by one path.
+Reiterating this step yields a polynomial-time algorithm that produces a path cover of size at most $\alpha(G)$.
+Notably, the algorithm does not need to know $\alpha(G)$ in advance; it merely keeps merging pairs of paths with adjacent endpoints until no such pair remains.
+The procedure can begin from the trivial cover formed by $\lvert V(G)\rvert$ single-vertex paths.
+The procedure above does not, by itself, provide any additional relationship between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ beyond $p\leq\alpha(G)$.
+To address this issue, we introduce the notions of usual and special paths.
+A path is called usual if its endpoints are distinct and no edge connects them in $G$; otherwise, the path is special.
+We argue that having usual paths in $\mathcal{P}$ is advantageous.
+For instance, if all paths in $\mathcal{P}$ are usual, then the endpoints of these paths form an independent set of size $2p$.
+(Otherwise, we could merge two paths whose endpoints are adjacent.) In particular, this implies $p\leq\alpha(G)/2$.
+Even having $k$ usual paths is sufficient, since their endpoints would then form an independent set of size $p+k$.
+Alongside $\mathcal{P}$, this independent set would constitute a feasible output for the algorithm in Theorem 1.
+While special paths are not helpful for constructing a large enough independent set, they are convenient for merging.
+A special path is either a path of length zero or one, or it can be viewed (along with the edge between its endpoints) as a cycle of length at least three.
+Then, an edge between any two vertices of different special paths is as effective as an edge between the endpoints: it enables us to merge those two paths and cover all their vertices with a single path.
+Our algorithm takes advantage of this property by exhaustively merging special paths until no edges remain between them.
+(a)  Refer to caption      (b)  Refer to caption      (c)  Refer to caption
+Figure 1: Example of a three-path transformation.
+In part (a), $P_{2}$ is a usual path, while $P_{1}$ and $P_{3}$ are special paths.
+Dashed lines represent some of the absent edges.
+Part (b) represents the same paths connected by edges required for the transformation.
+Part (c) shows how new paths are formed.
+$P_{1}$ and $P_{3}$ become usual paths, while $P_{2}$ can become special or even disappear completely.
+This property of special paths offers additional insight into the structure of $G$.
+Because there are no edges with endpoints in distinct special paths, and there are fewer than $k$ usual paths in $\mathcal{P}$, selecting one vertex from each special path already forms an independent set of size exceeding $p-k$.
+Moreover, if we choose two non-adjacent vertices from each of $2k$ special paths, we obtain an independent set of size at least $p+k$, immediately resolving the problem.
+Thus, the only remaining scenario is that, except for at most $2k$ special paths, vertex set of each of the special paths is a clique in $G$.
+However, this structural insight alone addresses only part of the problem.
+To progress further, we require a more sophisticated transformation of the paths in $\mathcal{P}$.
+Rather than merging two paths, we now consider two special paths and one usual path such that there are two edges (with no shared endpoints) connecting the special paths to the usual path.
+Using these edges, the transformation creates three new paths, at least two of which become usual.
+As a result, the number of special paths in $\mathcal{P}$ decreases without increasing the total number of paths (see Figure 1).
+If all these transformations cannot be applied, we arrive at the following situation.
+There must exist a vertex set $S$ of size smaller than $k$ that separates all usual paths from at least $p-4k$ special paths, each of which induces a clique in $G$.
+In the auxiliary Lemma 1, we demonstrate how this small separator $S$ can be exploited to remove irrelevant cliques from $G$ without affecting the size of an optimal path cover or the independence number.
+Applying this polynomial-time procedure ensures that, in the end, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
+By applying transformations and irrelevant cliques removals, we end up in the situation, where $p=\mathcal{O}(k^{2})$.
+In this case, we invoke Theorem 2 to compute an optimal path cover of $G$, using $p+k$ as its parameter.
+If Theorem 2 does not return an optimal path cover, it instead provides an independent set of size $p+k$.
+In either outcome, the algorithm obtains a valid solution for Theorem 1.
+All steps of the algorithm are performed in polynomial time, except for the call to Theorem 2, resulting in running time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$.
+Overview of the proof of Theorem 2: We focus on outlining the main ideas of the proof specifically for $-\ell$-linkages.
+The more general extension to arbitrary list TM-embeddings is similar but requires additional technical details.
+To begin, we sketch an $\operatorname{{\sf X}P}$ algorithm that solves $-\ell$-Linkage in time $n^{f(\alpha(G)+\ell)}$.
+Although this algorithm does not achieve the desired complexity, it is instructive for understanding the proof.
+Suppose we are given a parameter $k\geq\alpha(G)$.
+(This assumption is acceptable for an $\operatorname{{\sf X}P}$ algorithm because $k$ can be computed by brute force in $\mathcal{O}(n^{\alpha(G)})$ time.) The $\operatorname{{\sf X}P}$ algorithm is recursive and is inspired by the Chvátal-Erdős’s theorem on highly connected graphs, see Proposition 1.
+Informally, Chvátal-Erdős’s theorem states that if a graph lacks a large independent set but is highly connected, it necessarily contains a Hamiltonian path.
+If its connectivity is low, we find a small vertex cut, delete it, and then recursively process the connected components of the resulting graph.
+Because a Hamiltonian path might traverse each component multiple times, we find it more convenient to work with $-\ell$-linkages, that is, linkages consisting of vertex-disjoint paths connecting specified pairs of vertices and covering all vertices of the graph.
+To handle such cases, one can strengthen Chvátal-Erdős theorem to apply to Hamiltonian-linkages.
+Following the foundational work of Bollobás and Thomason [BT96], there have been numerous advances in understanding how linkages relate to connectivity.
+However, our setting additionally requires that the linkages are Hamiltonian.
+While no existing results directly address this requirement, we adapt the result of Thomas and Wollan [TW05] on linkages in highly connected graphs to show that if a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked.
+In other words, there exists a path cover of size $\ell$ that links any prescribed set of endpoints.
+The algorithm proceeds as follows.
+If the input graph $G$ is $\Omega(k\cdot\ell)$-connected, then the algorithm correctly reports that $G$ is $-\ell$-linked.
+Otherwise, if the connectivity is lower, there exists a minimum vertex separator $X$ of $G$ with $\mathcal{O}(k\cdot\ell)$ vertices.
+Each connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component.
 For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$.
@@
 Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices.
-By treating each set of entry and exit points as new terminal vertices, the algorithm is called recursively on each component.
-By restricting the depth of the recursion by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths.
-Since the approach of Jedlicková and Kratochvíl exploits small separators separating highly connected components, the first natural direction to explore is whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm.
+The algorithm then treats these points as new terminals and makes a recursive call for each connected component.
+Restricting the recursion depth by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths.
+Overall, this yields a running time of $n^{f(\alpha(G)+\ell)}$.
+Since the $\operatorname{{\sf X}P}$ algorithm exploits small separators separating highly connected components, the first natural idea that comes to mind is to check whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm.
 The recursive understanding technique introduced by Chitnis et al.
-[${}^{+}$16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al.
+[CCH+16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al.
 [LRSZ18].
-Unfortunately, recursive understanding is not of much help here due to the following reasons.
-To apply recursive understanding, one has to be able to deal with the primary case of so-called $(p,q)$-unbreakable graphs.
-However, for our problem, this primary case is not more straightforward than the case of general graphs.
-This is why we employ a different method of decomposing the graph.
-To prevent the $\operatorname{{\sf X}P}$ running time, our algorithm avoids recursive calls to itself.
+Unfortunately, recursive understanding does not appear to be particularly helpful in our setting for the following reason: in order to apply recursive understanding, one must handle the so-called $(p,q)$-unbreakable graphs as the “base case.” However, for our problem, addressing $(p,q)$-unbreakable graphs is not simpler than handling the general case.
+Thus, we need a different method of decomposing the graph.
+To prevent the $\operatorname{{\sf X}P}$ running time, the $\operatorname{{\sf FPT}}$ algorithm avoids recursive calls to itself.
 Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either output an independent set of $G$ of size $k$ or find a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of $G-S$ is $\Omega(k\cdot(\ell+|S|))$-connected.
@@
 Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time.
-However, guessing the entry and exit points of these parts, how it is done by Jedlicková and Kratochvíl, would be too time-consuming.
+However, guessing the entry and exit points of these parts, the way we describe it for the $\operatorname{{\sf X}P}$ algorithm, is too time-consuming.
 Here is precisely the moment when the property of the decomposition comes into play.
@@
 This way, the whole enumeration can be done in $\operatorname{{\sf FPT}}$ time.
-Another complication comes from the fact that we do not require the $\alpha(G)<k$ assumption and instead aim to find a solution or an independent set of size $k$.
-For that, a constructive version of the high connectivity case is required.
-That is, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we need to be able to construct a family of $\ell$ disjoint paths with the given endpoints spanning the graph or to find an independent set of size $k$.
-We employ the general result of Thomas and Wollan [TW05]: In a $10\ell$-connected graph, a collection of $\ell$ vertex-disjoint paths with the given endpoints always exists.
-However, the proof of Thomas and Wollan [TW05] is involved, and it is not apparent whether it could be made constructive.
-Instead, we employ the property that our graphs are of bounded independence number.
-We use a Ramsey-based argument, finding either an independent set of size $k$ or a large clique, and in the latter case, we can find the desired $\ell$ paths by connecting the endpoints to the clique via Menger’s theorem.
-We remark here that the relation between connectivity and linkages in graphs is well-studied.
-In particular, a sequence of results, including the above-mentioned work of Thomas and Wollan [TW05], guarantees the existence of $\ell$ vertex-disjoint paths between given pairs of endpoints in a graph of connectivity $\Omega(\ell)$.
-However, there is a lack of simple and efficient algorithms that compute vertex-disjoint paths connecting the given $\ell$ pairs of endpoints, in an $\Omega(\ell)$-connected, or even $\ell^{\Omega(1)}$-connected graph.
-Our contributions in that direction can be of independent interest to the reader.
+Another complication arises because we do not assume $\alpha(G)<k$.
+(Recall that we cannot compute $\alpha(G)$ in $\operatorname{{\sf FPT}}$ time.) Instead, we aim either to find a desired solution or an independent set of size $k$.
+This requirement calls for a constructive variant of the high-connectivity case.
+Specifically, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we must be able to construct $\ell$ disjoint paths connecting these endpoints and covering all vertices of the graph, or else find an independent set of size $k$.
+For this purpose, we rely on a combinatorial result by Thomas and Wollan [TW05], which states that in a $10\ell$-connected graph, one can always find $\ell$ vertex-disjoint paths connecting any given pairs of endpoints.
+However, the proof in [TW05] is non-constructive—it builds on certain properties of a minimal counterexample, and it is not obvious how to transform that argument into an algorithm that yields the required paths.
+To circumvent this issue, we make use of the fact that our graphs have bounded independence number.
+Through a Ramsey-based argument, we either locate an independent set of size $k$ or find a large clique.
+In the latter scenario, we can apply Menger’s theorem to connect the endpoints to the clique, thereby constructing the required $\ell$ paths.
+As a side remark, several combinatorial results—including the aforementioned work of Thomas and Wollan [TW05]—guarantee the existence of $\ell$ vertex-disjoint paths between specified pairs of endpoints in a graph with connectivity $\Omega(\ell)$.
+However, there is a notable shortage of simple and efficient algorithms to compute such vertex-disjoint paths in $\Omega(\ell)$-connected (or even $\ell^{\Omega(1)}$-connected) graphs.
+Our contribution in this direction may therefore be of independent interest.
 Finally, for Theorem 2, we additionally generalize the argument above from $-\ell$-linkages to arbitrary list TM-embeddings.
@@
 Then, similarly to $-\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly-connecting component (or an independent set of size $k$), with arbitrary locations of the terminals.
-Theorem 2: To sketch our second main result, Theorem 2, we recall the proof of Gallai-Milgram theorem for undirected graphs.
-It is based on the following observation: if there are more than $\alpha(G)$ paths in a path cover of $G$, then there is an edge between two endpoints of distinct paths.
-Then these two paths can be joined into a single one and the size of the path cover is reduced by one.
-This proof gives a polynomial-time algorithm constructing a path cover of $G$ consisting of at most $\alpha(G)$ paths.
-Note that the algorithm does not have to know or compute $\alpha(G)$ itself: it is enough to iteratively join a pair of paths with adjacent endpoints until it is no longer possible.
-The algorithm can start from a trivial path cover consisting of $|V(G)|$ empty paths.
-The procedure above on its own does not provide any additional relation between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ except for $p\leq\alpha(G)$.
-We propose a way to deal with this issue.
-First, we distinguish the paths in $\mathcal{P}$ into usual and special.
-We call a path usual if its endpoints are distinct and there is no edge between them in $G$.
-Otherwise, we call a path special.
-We argue that having usual paths in $\mathcal{P}$ is beneficial.
-For example, if all paths in $\mathcal{P}$ are usual, then the endpoints of paths in $\mathcal{P}$ form an independent set of size $2p$ as, otherwise, we would join two paths.
-In particular, this implies that $p\leq\alpha(G)/2$.
-Even having $k$ usual paths in $\mathcal{P}$ is enough since this yields an independent set of size $p+k$.
-Together with $\mathcal{P}$, this independent set is a feasible output for the algorithm of Theorem 2.
-While special paths are unsuitable for constructing a large enough independent set, they are suitable for joining.
-A special path is either a path of length zero or one, or a cycle of length at least three (if considered together with the edge between the endpoints).
-Then, an edge between two arbitrary vertices of two special paths is as helpful as an edge between endpoints: it allows us to join the two paths and cover the vertices of their union by a single path.
-Our algorithm exploits this property of special paths and joins them exhaustively until there is no edge between special paths.
-This property of special paths, in fact, gives us even more insight into the structure of $G$.
-Since special paths are independent of each other, and there are less than $k$ usual paths in $\mathcal{P}$, taking just one vertex from each special path constitutes an independent set of size greater than $p-k$.
-Taking two non-adjacent vertices instead of one from $2k$ special paths gives an independent set of size at least $p+k$, solving the problem immediately.
-Hence, the remaining case is when all but at most $2k$ special paths have vertex sets inducing cliques in $G$.
-However, this structural insight alone is only half of the solution.
-To proceed further, we need another and more sophisticated way of transforming the paths in $\mathcal{P}$.
-Instead of joining two paths as before, this transformation takes two special paths and one usual path, so there are two edges without common endpoints connecting the special paths to the usual path.
-The transformation produces three paths, and at least two of these paths are usual.
-Hence, this transformation reduces the number of special paths and does not increase $\mathcal{P}$.
-The transformation above provides the remaining structural insight for the algorithm.
-When this transformation and all the transformations above cannot be applied, a vertex set $S$ of size less than $k$ separates all usual paths from at least $p-4k$ special paths with vertex sets inducing cliques.
-In an auxiliary Lemma 4, we prove that the small size of $S$ can be exploited and irrelevant cliques can be removed from $G$ with respect to the size of an optimal path cover and the independence number.
-After applying this polynomial-time routine, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
-The proof outline has come to the final case when $p=\mathcal{O}(k^{2})$.
-In this case, with Theorem 1 in hand, we aim to find the optimal path cover of $G$ as discussed previously in this section.
-The parameter choice in Theorem 1 is essential, and we choose it to be $p+k$.
-Thus, if the application of Theorem 1 does not give an optimal path cover, it gives an independent set of size $p+k$.
-Either way, the algorithm produces a solution feasible for Theorem 2.
-We finish the outline of the proof of Theorem 2 by noting that the algorithm described above takes polynomial time, except for the application of Theorem 1.
+1.3 Related Work.
+The interplay among the independence number, connectivity, Hamiltonicity, and path covers has long been a popular theme in graph theory, with historical roots going back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72].
+Surveys by Jackson and Ordaz [JO90], Bondy [Bon95], and Kawarabayashi [Kaw01] offer excellent overviews of this area.
+However, most of the research thus far has been centered on establishing combinatorial conditions for Hamiltonicity (or related properties), rather than on algorithmic extensions.
+Theorem 1 contributes to a growing line of parameterized algorithms that extend classic theorems from extremal graph theory.
+Examples include algorithmic adaptations of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem [GHH23].
+Such algorithmic extensions can be seen as part of a broader trend in parameterized complexity called above (or below) guarantee parameterization, where one observes that: the natural parameterization of a maximization or minimization problem by the solution size is not satisfactory if there is a lower bound for the solution size that is sufficiently large.
+The term above guarantee parameterization was introduced by Mahajan and Raman [MRS09], and it has proved fruitful for studying numerous fundamental problems in parameterized complexity and kernelization [AGK+10, BCDF19, CJM+13, FGL+21, FGSS22b, GP16a, GKLM11, GvIMY12, GP16b, GRSY07, LNR+14, GM22].
+A feature distinguishing Theorem 1 from the previous work is that it provides “tractability below intractable parameter”.
+The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms.
+Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved.
+Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16].
+This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, FLP+20].
+A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph [CNP+11, BCKN13, CKN13, FLPS16, BST23], see also [CFK+15].
+Such parameters include treewidth, treedepth, or minimum vertex cover size.
+On the other hand, for more “general” parameters Hamiltonicity becomes intractable.
+For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10] and is paraNP parameterized by $\alpha$-treewidth [DMS24].
+The existence of a Hamiltonian cycle remains NP-complete on planar graphs [GJT76], circle graphs [Dam89] and for several other generalizations of interval graphs [BB86], for chordal bipartite graphs [Mül96] and for split (and therefore also for chordal) graphs [Gol04], but is solvable in linear time in interval graphs [Kei85] and in convex bipartite graphs [Mül96].
+The existence of a Hamiltonian path can be decided in polynomial time for cocomparability graphs [DDKS91] and for circular arc graphs [Dam93].
+Many of the above mentioned graph classes are hereditary, i.e., closed in the induced subgraph order, and as such can be described by collections of forbidden induced subgraphs.
+This has led to carefully examining $H$-free graphs, i.e., graph classes with a single forbidden induced subgraph.
+For many graphs $H$, the class of $H$-free graphs has nice structural properties (e.g., $P_{4}$-free graphs are exactly the cographs) or their structural properties can be used to design polynomial algorithms for various graph theory problems (as a recent example, cf.
+colouring of $P_{6}$-free graphs in [CSZ19]).
+For Hamiltonian-type problems, consider three-vertex graphs $H$.
+For $H=K_{3}$, both Hamiltonian cycle and Hamiltonian path are NP-complete on triangle-free graphs, since they are NP-complete on bipartite graphs [Mül96].
+For $H=3K_{1}$, the edgeless graph on three vertices, both Hamiltonian cycle and Hamiltonian path are polynomial time solvable [DGJ81].
+The remaining two graphs, $P_{3}$ and $K_{1}+K_{2}$, are induced subgraphs of $P_{4}$, thus the corresponding class of $H$-free graphs is a subclass of cographs, and as such a subclass of cocomparability graphs, in which the Hamiltonian path and Hamiltonian cycle problems are solvable in polynomial time [DDKS91].
+However, a complete characterization of graphs $H$ for which Hamiltonian path or Hamiltonian cycle problems are solvable in polynomial time (and for which they are NP-complete) is not in sight.
+Until now, the complexity has been open even for the next smallest edgeless graph $4K_{1}$.
+Theorem 2 answers this complexity question for all edgeless forbidden induced subgraphs.
+In a companion paper Jedličková and Kratochvíl [JK24] explicitly described situations in which a Hamiltonian path for $\alpha(G)=3,4$ and $5$ in which a Hamiltonian path does exist.
+The parameterized complexity of many graph problems remains insufficiently explored for parameters capturing a graph’s density—such as its independence number.
+It is surprising that, despite the substantial progress in understanding Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth (which are sparse), relatively little has been achieved for graphs of bounded independence number (which are typically dense).
+Only a handful of results ($\operatorname{{\sf FPT}}$ or $\operatorname{{\sf X}P}$) address parameterization by a graph’s independence number.
+These include the $\operatorname{{\sf X}P}$ algorithm of Ovetsky Fradkin and Seymour [FS15] for the edge-disjoint paths problem in directed graphs, the kernelization algorithms of Lochet et al.
+[LLM+20] for various cycle-hitting problems in directed graphs, the subexponential parameterized algorithms of Misra et al.
+[MSSZ23] for digraphs with constant independence number, and the recent work of Maalouly, Steiner, and Wulf [MSW23] on $\operatorname{{\sf FPT}}$ algorithms for exact matching.
+2 Preliminaries
+We consider simple undirected graphs without loops or multiple edges.
+The vertex set of a graph $G$ is denoted by $V(G)$, its edge set by $E(G)$.
+Edges are considered as two-element sets of vertices, thus we write $u\in e$ to express that a vertex $u$ is incident to an edge $e$.
+For the sake of brevity, we write $uv$ instead of $\{u,v\}$ for the edge containing vertices $u$ and $v$.
+We say that $u$ is adjacent to $v$ if $uv\in E(G)$.
+The degree of a vertex is the number of other vertices adjacent to it.
+The subgraph of $G$ induced by vertices $S\subseteq V(G)$ will be denoted by $G[S]$.
+We use $G-S$ for $G[V(G)\setminus S]$.
+The independence number of a graph $G$, denoted by $\alpha(G)$, is the order of the largest edgeless induced subgraph of $G$.
+With the standard notion of $K_{k}$ being the complete graph with $k$ vertices and $G+H$ being the disjoint union of graphs $G$ and $H$, $\alpha(G)$ is equal to the largest $k$ such that $kK_{1}$ is an induced subgraph of $G$.
+A graph is called $H$-free if it contains no induced subgraph isomorphic to $H$.
+A path in a graph $G$ is a sequence of distinct vertices such that any two consecutive ones are adjacent.
+The length of a path is the number of its edges.
+A cycle is formed by a path of length greater than 1 which connects two adjacent vertices.
+The path (cycle) is Hamiltonian if it contains all vertices of the graph.
+The path cover number of $G$, denoted by $\mathrm{pc}(G)$, is the smallest number of vertex disjoint paths that cover all vertices of $G$.
+A graph is connected if any two vertices are connected by a path.
+Since the problems we are interested in are either trivially infeasible on disconnected graphs, or can be reduced to studying the components of connectivity one by one, we only consider connected input graphs in the sequel.
+A vertex cut in a graph $G$ is a set $A\subset V(G)$ of vertices such that the graph $G-A=G[V(G)\setminus A]$ is disconnected.
+The vertex connectivity $c_{v}(G)$ of a graph $G$ is the order of a minimum cut in $G$, or $|V(G)|-1$ if $G$ is a complete graph.
+Since we will not consider edge connectivity, we will often omit the adjective when talking about the connectivity measure, we always have vertex connectivity in mind.
+Although we consider undirected graphs, when we talk about a path in a graph, the path itself is considered traversed in the direction from its starting vertex to the ending one.
+Formally, when we say that a path $P$ connects a vertex $x$ to a vertex $y$, then by $P^{-1}$ we denote the same path, but traversed from $y$ to $x$.
+This is important when creating a longer path by concatenating shorter ones.
+We build upon the following results of Chvátal and Erdős which we will later generalize to Hamiltonian linkages in Theorem 2.
+Proposition 1.
+[CE72] Let $G$ be an $s$-connected graph.
+  1. 1.
+    If $\alpha(G)<s+2$, then $G$ has a Hamiltonian path.
+  2. 2.
+    If $\alpha(G)<s+1$, then $G$ has a Hamiltonian cycle.
+  3. 3.
+    If $\alpha(G)<s$, then $G$ is Hamiltonian connected (i.e., every pair of vertices is joined by a Hamiltonian path).
+3 Proof of Theorem 1: Path cover bellow Gallai-Milgram
+In this section, we prove Theorem 1.
+Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths.
+Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$.
+A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$.
+When we combine this theorem with Theorem 2, we obtain an algorithm solving Path Cover in time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$.
+Our contribution in Theorem 1 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem.
+We restate Theorem 1 here for the reader’s convenience.
+See 1
+We require two auxiliary results to prove Theorem 1.
+3.1 Small separators: Reducing connected components
+The first result is dealing with small vertex separators of $G$.
+It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
+Lemma 1.
+Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of $G-S$.
+Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
+Proof.
+The construction of $X$ is based on the following observation. We say that a path $P$ in a path cover $\mathcal{P}$ of $G$ is degenerate if $V(P)=C_{i}$ for some $i\in[t]$.
+Claim 1.
+In any minimum path cover $\mathcal{P}$ of $G$, at most $2|S|$ of the sets $C_{1},C_{2},\ldots,C_{t}$ contain vertices of non-degenerate paths of $\mathcal{P}$.
+Proof of Claim 1.
+Notice that if $\mathcal{P}$ has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path by the minimality of $\mathcal{P}$.
+Hence, if $C_{i}$ contains a vertex of a non-degenerate path then $\mathcal{P}$ has a path $P$ with an edge with one endpoint in $S$ and the other in $C_{i}$.
+Each vertex in $S$ is incident to at most two edges in the paths from $\mathcal{P}$.
+Thus, at most $2|S|$ edges are in the path of $\mathcal{P}$ with one endpoint in $S$ and the other in one of $C_{1},C_{2},\ldots,C_{t}$.
+Hence, at most $2|S|$ of the sets can contain vertices of non-degenerate paths.
+∎
+To construct $X$, for each vertex $v\in S$, we mark at most $2|S|$ cliques from $\{C_{i}\}_{i\in[t]}$ that contain neighbors of $v$.
+If there are at most $2|S|$ such cliques, then we mark all of them.
+Otherwise, we mark arbitrary $2|S|$ cliques containing neighbors of $v$.
+Then $X$ is defined as the set of indices of the marked cliques.
+It is straightforward that $|X|\leq 2|S|^{2}$ and $X$ can be constructed in polynomial time.
+We claim that there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$.
+In other words, each unmarked clique is covered by a degenerate path.
+Given a path cover $\mathcal{P}$ of $G$, we say that an edge $uv$ of a path $P\in\mathcal{P}$ is bad if $u\in S$ and $v$ is a vertex of an unmarked clique.
+We show that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges.
+The proof is by contradiction.
+We choose $\mathcal{P}$ as a minimum path cover of $G$ that minimizes the total number of bad edges in the paths.
+Suppose that there is $P\in\mathcal{P}$ that contains a bad edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$.
+We have that $C_{i}$ is unmarked and has a neighbor of $u$.
+Then, by the definition of the marking procedure, there are at least $2|S|$ marked cliques with neighbors of $u$.
+Because $C_{i}$ contains a vertex of the non-degenerate path $P$, we obtain that there is a marked clique $C_{j}$ for some $j\in[t]$ with a neighbor $w$ of $u$ such that $C_{j}$ is covered by a degenerate path $P^{\prime}\in\mathcal{P}$ by 1.
+Because $C_{j}$ is a clique, we can assume without loss of generality that $w$ is an end-vertex of $P^{\prime}$.
+Deleting the edge $uv$ separates $P$ into two paths $Q$ and $Q^{\prime}$ having end-vertices $v$ and $u$, respectively.
+We use the fact that $uw\in E(G)$ and construct the path $\hat{Q}$ from $Q^{\prime}$ and $P^{\prime}$ by joining their end-vertices $u$ and $w$ by the edge $uw$.
+Notice that $Q$ and $\hat{Q}$ are vertex-disjoint and $V(Q)\cup V(\hat{Q})=V(P)\cup V(P^{\prime})$.
+This allows us to construct the path cover $\mathcal{P}^{\prime}$ by replacing $P$ and $P^{\prime}$ in $\mathcal{P}$ by $Q$ and $\hat{Q}$.
+Because the number of paths in $\mathcal{P}^{\prime}$ is the same as in $\mathcal{P}$, $\mathcal{P}^{\prime}$ is a minimum path cover.
+However, the number of bad edges in the path of $\mathcal{P}^{\prime}$ is less than the number of bad edges in the paths of $\mathcal{P}$, contradicting the choice of $\mathcal{P}$.
+This proves that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges.
+To complete the proof, we remind that if a minimum path cover has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path.
+Thus, there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$.
+∎
+3.2 The case of $\operatorname{pc}(G)<k^{C}$
+The second auxiliary result is a “lighter version” of Theorem 1. It deals with the particular case of $\operatorname{pc}(G)<k^{C}$.
+Lemma 2.
+There is an algorithm with running time $2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$ that, given an $n$-vertex graph $G$, integer $k\geq 0$ and a path cover of $G$ with at most $\mathcal{O}(k^{C})$ paths for constant $C\geq 1$, outputs a path cover $\mathcal{P}$ of $G$ and
+  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
+  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+Proof.
+We present an algorithm $\mathcal{A}$. The algorithm assumes that $G$ contains at least one edge as, otherwise, $pc(G)=n$ and the minimum-size path cover consists of trivial paths.
+Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$.
+To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$.
+On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching.
+Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial.
+The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 2 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$.
+The subroutine works in
+$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}}\cdot 2^{(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
+running time.
+If the outcome of the subroutine is an independent set $I$ in $G$ of size $p^{\prime}+k$, then the algorithm outputs $\mathcal{P}^{\prime}$ and $I$ as the solution.
+Since $|\mathcal{P}^{\prime}|=p^{\prime}$, $|I|\geq|\mathcal{P}^{\prime}|+k$ is satisfied, so this solution satisfies the lemma statement.
+If the subroutine decides that $H_{ij}$ does not have TM-model in $G$ or the maximum TM-model of $H_{ij}$ in $G$ is not spanning, then $G$ cannot be covered with $i$ paths and $\mathcal{A}$ continues to the next iteration.
+If the subroutine gives a spanning TM-model of $H_{ij}$ in $G$, then the minimum path cover size of $G$ equals $i$ and $\mathcal{A}$ transforms this TM-model into a path cover of $G$ with $i$ paths.
+Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops.
+Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$.
+The proof of the lemma is complete.
+∎
+3.3 Proof of Theorem 1
+We are ready to prove Theorem 1.
+We describe an algorithm that starts from a trivial path cover of $G$ and then gradually transforms it.
+These transformations are indicated as reduction rules and are performed in polynomial time.
+We show that each such rule does not increase the number of paths in the path cover.
+Moreover, each reduction rule necessarily reduces either the total number of paths or the number of special (to be defined later) paths in the path cover.
+A series of consecutive applications of the reduction rules always lead the algorithm to cases when the algorithm can produce a solution satisfying theorem statement in $f(k)\cdot n^{\mathcal{O}(1)}$ time and stop.
+We indicate such cases and the corresponding behavior of the algorithm as solution subroutines.
+The algorithm runtime thus always ends with a solution subroutine run.
+In what follows, we present the reduction rules and solution subroutines.
+Each of the reduction rules reduces the number of paths in a path cover or reduces the number of some special paths.
+Solution subroutines output path covers and independent sets.
+The rules and subroutines are given in the order in which the algorithm checks their applicability.
+That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied.
+When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties.
+In this case, the final solution subroutine calls the algorithms of Lemmas 1 and 2.
+The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 00.
+Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$.
+Let the current set of paths be $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{m}\}$.
+We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
+The first reduction rule is straightforward and comes from the proof of the Gallai-Milgram theorem for undirected graphs.
+Reduction Rule 1.
+If there is an edge $uv$ in $G$ such that $u\in\{s_{i},t_{i}\}$ and $v\in\{s_{j},t_{j}\}$ and $i\neq j$, then join $P_{i}$ and $P_{j}$ into a single path via $uv$.
+The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one.
+Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$.
+When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
+Special and usual paths.
+We call a path $P_{i}$ in the path cover special, if either $s_{i}=t_{i}$, or $P_{i}$ consists of only edge $s_{i}t_{i}$, or its endpoints are connected by the edge $s_{i}t_{i}$.
+If a path $P_{i}$ is not special, we call it usual.
+We claim that sufficiently many usual paths give a solution to the problem.
+Claim 2.
+If at least $k$ paths are usual in the path cover $\mathcal{P}=P_{1},\ldots,P_{m}$ and 1 is not applicable, then $m\leq\alpha(G)-k$.
+Moreover, an independent set $I$ in $G$ of size $m+k$ is computable in polynomial time.
+Proof of Claim 2. We show that $G$ has an independent set $I$ of size $m+k$. Initialize $I$ with $\{s_{1},s_{2},\ldots,s_{m}\}$. Since 1 is not applicable, $I$ is independent in $G$.
+Let $i_{1},i_{2},\ldots,i_{k}$ be $k$ distinct indices of usual paths.
+Since each $P_{i_{j}}$ is usual, $t_{i_{j}}\neq s_{i_{j}}$ and $s_{i_{j}}$ is not connected with $t_{i_{j}}$.
+Also, 1 is not applicable, so there is no edge between $t_{i_{j}}$ and any vertex in $I$, while $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}\}$ is also an independent set in $G$.
+Put $t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}$ in $I$. We obtain an independent set of size $m+k$ in $G$, consequently $\alpha(G)-k\geq m$. ∎
+2 yields the corresponding solution subroutine for the algorithm.
+Solution Subroutine 1.
+If there are at least $k$ usual paths among $\mathcal{P}=P_{1},P_{2},\ldots,P_{m}$, output $\mathcal{P}$ and $I$ as in 2.
+From now on, we assume that $\mathcal{P}$ has less than $k$ usual paths.
+The remaining reduction rules deal primarily with special paths.
+The following rule reduces the total number of paths given that $G$ has an edge between any vertex of a special path and an endpoint of an arbitrary other path.
+Reduction Rule 2.
+If $P_{i}$ is a special path of length at least two and there is an edge in $G$ connecting some $u\in V(P_{i})$ with $s_{j}$ (or $t_{j}$) for some $j\neq i$, then
+  1. 1.
+    Add edge $s_{i}t_{i}$ to $P_{i}$, so it becomes a cycle;
+  2. 2.
+    Remove edge $uv$ from $P_{i}$, where $v$ is any neighbor of $u$ in $P_{i}$, so it becomes an $uv$-path;
+  3. 3.
+    Join $P_{i}$ and $P_{j}$ via $us_{j}$ (or $ut_{j}$).
+2 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, so $\mathcal{P}$ remains a path cover implying that the rule is correct.
+The following reduction rule deals with an edge between two vertices of distinct special paths.
+It is demonstrated in Figure 2a.
+Reduction Rule 3.
+If $P_{i}$ and $P_{j}$ are special paths of length at least two for $i\neq j$, and there is an edge in $G$ connecting $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$, then
+  1. 1.
+    Add edge $s_{j}t_{j}$ to $P_{j}$, so it becomes a cycle;
+  2. 2.
+    Remove edge $u_{j}v_{j}$ from $P_{j}$, so it becomes an $u_{j}v_{j}$-path.
+  3. 3.
+    Apply 2 to $P_{i}$, $P_{j}$ and $u_{i}u_{j}$.
+Again, 3 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, thus reducing the number of paths in $\mathcal{P}$. We claim the following.
+Claim 3.
+If Reduction Rules 1, 2, and 3 are not applicable, then for each pair of distinct special paths $P_{i},P_{j}$ in $\mathcal{P}$, there is no edge between $V(P_{i})$ and $V(P_{j})$ in $G$.
+Proof of Claim 3.
+Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$.
+If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable.
+Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$.
+If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction.
+Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$.
+In this case, 1 is applicable, leading us to the final contradiction.
+∎
+The claim is the basis for the next solution subroutine, which the algorithm applies next. We explain how it works by proving the following claim.
+Claim 4.
+If 1 and Reduction Rules 1, 2, and 3 are not applicable, and there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then $m<\alpha(G)-k$.
+Moreover, an independent set $I$ in $G$ of size more than $m+k$ is computable in polynomial time.
+Proof of Claim 4.
+We construct an independent set $I$ consisting of more than $m+k$ vertices.
+Since 1 is not applicable, there are more than $m-k$ special paths.
+For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$.
+For each other special path, put its arbitrary vertex into $I$.
+By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎
+The solution subroutine based on 4 is given below. The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
+Solution Subroutine 2.
+If there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then output $\mathcal{P}$ and an $I$ as in 4.
+The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$.
+The following reduction rule necessary reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
+a $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$ b $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$$s_{\ell}$$t_{\ell}$$x_{i}$$x_{j}$
+Figure 2: Illustration of transformations performed in (a) 3 and (b) 4..
+Paths before the transformations are horizontal.
+Paths after the transformations are highlighted in bold and have pairwise-distinct colors.
+Dashed lines represent absence of an edge.
+Snake lines represent paths instead of single edges.
+Reduction Rule 4.
+If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$ and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
+  1. 1.
+    Make $u_{i}$ endpoint of $P_{i}$ and $u_{j}$ endpoint of $P_{j}$ similarly to the first two steps of 2;
+  2. 2.
+    Break $P_{\ell}$ into three (or two) parts: the prefix part $s_{\ell}P_{\ell}x_{i}$, the suffix part $x_{j}P_{\ell}t_{\ell}$, and (if exists) the remaining middle part;
+  3. 3.
+    Join $P_{i}$ with the prefix part via $u_{i}x_{i}$;
+  4. 4.
+    Join $P_{j}$ with the suffix part via $u_{j}x_{j}$.
+Figure 2b illustrates the result of application of 4. The following claim addresses its correctness.
+Claim 5.
+If Reduction Rules 1, 2 are not applicable, then 4 reduces the number of special paths in $\mathcal{P}$ and does not increase the number of paths in $\mathcal{P}$.
+Proof of Claim 5.
+The second part of the claim is trivial, since 4 produces three (or two, if $x_{i}x_{j}\in E(P_{\ell})$) paths by transforming three distinct paths $P_{i},P_{j},P_{\ell}$, so $|\mathcal{P}|$ cannot increase.
+To prove the first part, we show that the resulting $v_{i}s_{\ell}$-path and $v_{j}t_{\ell}$-path are usual.
+We denote these two paths by $Q_{i}$ and $Q_{j}$ correspondingly.
+We give the proof only for $Q_{i}$, since the proof for $Q_{j}$ is symmetric.
+We first show that $|V(Q_{i})|\geq 3$.
+Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$.
+The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly.
+Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold.
+But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$.
+Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$.
+This contradiction proves $|V(Q_{i})|\geq 3$.
+It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$.
+Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable.
+This contradiction proves the claim.
+∎
+The list of reduction rules of the algorithm is exhausted.
+Starting from this point, the algorithm exploits the structure of the graph, assuming that no reduction rule is applicable.
+In what follows, we explain these structural properties.
+We also give the last solution subroutine of the algorithm.
+Claim 6.
+If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$.
+Moreover, such a set $S$ could be constructed in polynomial time.
+Proof of Claim 6.
+Since 2 is not applicable, there are less than $k$ usual paths in $\mathcal{P}$.
+We construct $S$ by taking one or zero vertices from each usual path, so $|S|<k$ holds automatically.
+We now explain the choice of the vertex to put in $S$ for each usual path.
+Consider a usual path $P_{i}\in\mathcal{P}$.
+We call a vertex $u\in V(P_{i})$ a connector if there is a special path $P_{j}$ and a vertex $v\in V(P_{j})$ with $uv\in E(G)$.
+If $P_{i}$ has exactly one connector, then we put this connector in $S$.
+Otherwise, i.e.
+when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$.
+We show that at least $m-2k$ special paths are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded.
+Assume now that $P_{j}$ is a special path that is not isolated in $G-S$.
+Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$.
+By 3, $P_{i}$ can only be usual.
+Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors.
+As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$.
+Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$.
+Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$.
+∎
+We are ready to present the final solution subroutine which combines all algorithmic results given previously in this section.
+Solution Subroutine 3.
+Perform the following steps:
+  1. 1.
+    Obtain $S$ according to 6 and put $h:=\max\{0,m-4k\}$;
+  2. 2.
+    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in $G-S$;
+  3. 3.
+    Apply the algorithm of Lemma 1 to $G$, $S$ and $V(P_{1}),V(P_{2}),\ldots,V(P_{h})$ and obtain a set $X$ of at most $2k^{2}$ indices, where each $i\in X$ satisfies $i\in[h]$;
+  4. 4.
+    Obtain $G^{\prime}$ and $\mathcal{P}^{\prime}$ by removing $V(P_{j})$ from $G$ and $P_{j}$ from $\mathcal{P}$, for each $j$ such that $j\in[h]$ and $j\notin X$;
+  5. 5.
+    Apply the algorithm of Lemma 2 as a subroutine to $G^{\prime},\mathcal{P}^{\prime}$ and $k^{\prime}:=2k$ and obtain a path cover $\mathcal{S}^{\prime}$ of $G^{\prime}$ and (possibly) an independent set $I^{\prime}$ in $G^{\prime}$;
+  6. 6.
+    Obtain a path cover $\mathcal{S}$ of $G$ by adding all paths in $\mathcal{P}\setminus\mathcal{P}^{\prime}$ to $\mathcal{S}^{\prime}$;
+  7. 7.
+    If $I^{\prime}$ was obtained, remove all vertices of $S$ from $I^{\prime}$ and for each path $P_{j}\in\mathcal{P}\setminus\mathcal{P}^{\prime}$, add one vertex of $P_{j}$ to $I^{\prime}$.
+Denote the resulting set by $I$;
+  8. 8.
+    Output $\mathcal{S}$ and (if obtained) $I$.
+We conclude the proof with a claim certifying the correctness of the solution subroutine.
+In contrast with all previous rules and subroutines, 3 requires superpolynomial (but $\operatorname{{\sf FPT}}$ in $k$) running time.
+Claim 7.
+If Reduction Rules 1-4 and Solution Subroutines 1,2 are not applicable, then 3 works in $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$ time and produces a feasible solution.
+Proof of Claim 7.
+First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 6 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
+We now move on to establish the upper bound on the running time. All steps, except for applying Lemma 2, are carried out in polynomial time. We have
+$|\mathcal{P}^{\prime}|=|\mathcal{P}|-h+|X|=m-h+|X|\leq 4k+2k^{2}=\mathcal{O}(k^{2}).$
+By applying Lemma 2 to $G^{\prime},k^{\prime}$ and $\mathcal{P}^{\prime}$, we obtain the running time $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$.
+Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$.
+Hence it is also a path cover.
+On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$.
+Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$.
+First, let us consider the scenario where $I$ is not produced.
+This can occur only when the algorithm of Lemma 2 does not produce $I^{\prime}$.
+Then Lemma 2 guarantees that $\mathcal{S^{\prime}}$ is an optimal path cover of $G^{\prime}$ and $|\mathcal{S}^{\prime}|=\operatorname{pc}(G^{\prime})$.
+By Lemma 1, removal of $V(P_{j})$ for any $j\in[h]\setminus X$ decreases $\operatorname{pc}(G)$ by one, so $\operatorname{pc}(G^{\prime})=\operatorname{pc}(G)-(h-|X|).$ Since $|\mathcal{P}\setminus\mathcal{P}^{\prime}|=h-|X|$, we have $|\mathcal{S}|=\operatorname{pc}(G)$.
+It means that $\mathcal{S}$ is an optimal path cover of $G$.
+Finally, let us consider the situation where $I$ is produced.
+In this case, the algorithm in Lemma 2 provides $\mathcal{S^{\prime}}$ and $I^{\prime}$ with $|I^{\prime}|-|\mathcal{S}^{\prime}|=k^{\prime}=2k$.
+We notice that
+$|I|\geq|I^{\prime}|-|S|+|\mathcal{P}\setminus\mathcal{P}^{\prime}|>|I^{\prime}|-k+(h-|X|).$
+As discussed in the previous paragraph, $|\mathcal{S}|-|\mathcal{S}^{\prime}|=h-|X|$. We conclude that
+$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)+(h-|X|)=k.$
+The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
+This completes the proof of Theorem 1.
 4 Proof of Theorem 2: List TM-embedding parameterized by independence number
@@
 More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high.
-This step follows the approach of Jedlicková and Kratochvíl [JK23, Theorem 7], who proved that for every $k>1$ and $\ell>1$, if a graph $G$ satisfies $\alpha(G)<k$ and $G$ is $\max\{k\ell,10\ell\}$-connected, then $G$ is $-\ell$-linked.
 Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$.
-A graph is called $\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
-To adapt the approach of Jedlicková and Kratochvíl for our purposes, we need to address two issues.
-The first issue is minor; we need to slightly strengthen the connectivity requirement to simplify the case analysis and allow for possible coinciding endpoints of the paths (which is not allowed in [JK23]).
-The second issue is more serious—the proof of Jedlicková and Kratochvíl is not constructive, meaning that it cannot be used to construct the linkage.
-However, we need a procedure constructing the linkage for computing a TM-embedding of $H$.
-The proof of Jedlicková and Kratochvíl builds on the following result of Thomas and Wollan [TW05], which is also non-constructive.
+A graph is called $-\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
+We need a procedure constructing the linkage for computing a TM-embedding of $H$.
+For that we need a constructive version of the following theorem of Thomas and Wollan [TW05], which proof is non-constructive.
 Proposition 2 (Corollary 1.3, [TW05]).
@@
 Now, we are ready to complete the algorithm:
-  • If $|V(G)|\leq(k+2\ell)^{k}$, the algorithm applies the subroutine of 2 to find the solution, outputs it and stops.
-  • Otherwise, since $(k+2\ell)^{k}\geq\binom{k+2\ell-2}{k-1}$, the algorithm applies the subroutine of 3 with $r=k$ and $s=2\ell$.
+  • If $|V(G)|\leq(k+2\ell)^{k}$, the algorithm applies the subroutine of 9 to find the solution, outputs it and stops.
+  • Otherwise, since $(k+2\ell)^{k}\geq\binom{k+2\ell-2}{k-1}$, the algorithm applies the subroutine of 10 with $r=k$ and $s=2\ell$.
   • If the output is an independent set of size $k$, the algorithm outputs it, and stops.
-  • Otherwise, the output is a clique of size $2\ell$ in $G$. The algorithm uses the subroutine of 1 to find the solution, then outputs it and stops.
+  • Otherwise, the output is a clique of size $2\ell$ in $G$. The algorithm uses the subroutine of 8 to find the solution, then outputs it and stops.
 The correctness of the algorithm and the upper bound on its running time follow from the claims. The proof of the lemma is complete. ∎
@@
 Without loss of generality, $v_{a}$ goes before $v_{b}$ on $P_{i}$ in the order from $s_{i}$ to $t_{i}$.
-The algorithm then enlarges $P_{i}$ by replacing it with $(s_{i}P_{i}v_{a})\circ(v_{a}Q_{a}x)\circ(xQ_{b}v_{b})\circ(v_{b}P_{i}u_{a})\circ u_{a}u_{b}\circ(u_{b}P_{i}s_{i})$.
+The algorithm then enlarges $P_{i}$ by replacing it with $(s_{i}P_{i}v_{a})\circ(v_{a}Q_{a}x)\circ(xQ_{b}v_{b})\circ(v_{b}P_{i}u_{a})\circ u_{a}u_{b}\circ(u_{b}P_{i}t_{i})$.
 Again, $x$ is embedded into $P_{i}$ while no original vertices of $P_{i}$ are lost, increasing the total span.
@@
 This procedure is repeated at most $|V(G)|$ times, so the total running time of this part is polynomial in the size of the instance.
-The correctness of the algorithm follows from the discussion.
-The proof of the lemma is complete.
-∎
+The correctness of the algorithm follows from the discussion. The proof of the lemma is complete. ∎
 4.2 Merging Lemma
@@
 Proof.
+Figure 3: An illustration to the proof of Lemma 5: A graph $G$ with a TM-model of $H$, the graph $M_{W}$ of the cut descriptor corresponding to the model, and the graph $H$, the model of which is sought.
+The images of the vertices of $H$ in $G$ and $M_{W}$ that are fixed by $f$ and $f_{W}$, respectively, are indicated by the matching node shapes.
 The main idea of the proof is to enumerate all possible interactions of the desired TM-model with $S$.
@@
 As we will see later, such an implicit representation is sufficient since Lemma 4 always provides a spanning TM-model (or an independent set) and allows to fix an arbitrary injection of the terminals.
+See Figure3 for an illustration of the relationship between a TM-model of $H$ in $G$ and its cut descriptor.
 Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$.
-Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $W_{i}=\{w\in W:\xi(w)\in C_{i}\}$.
+Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)\in C_{i}\}$.
 We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@
 Let $(M,f)$ be a list TM-embedding of $H$ in $G$.
-Then there exist a cut descriptor $(W,M_{W},\xi,f_{W})$ such that $M_{W}[S]=M[S]$, and for each $i\in[t]$, $M$ hits $C_{i}$ if and only if $M_{W}$ hits $W_{i}$, touches $C_{i}$ if and only if $M_{W}$ touches $W_{i}$ and $|V(M)\cap C_{i}|=|W_{i}|$, and avoids $C_{i}$ if and only if $W_{i}=\emptyset$.
+Then there exist a cut descriptor $(W,M_{W},\xi,f_{W})$ such that $M_{W}[S]=M[S]$, and for each $i\in[t]$, $M$ hits $C_{i}$ if and only if $M_{W}$ hits $T_{i}$, touches $C_{i}$ if and only if $M_{W}$ touches $T_{i}$ and $|V(M)\cap C_{i}|=|T_{i}|$, and avoids $C_{i}$ if and only if $T_{i}=\emptyset$.
 Proof of Claim 12.
@@
 Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction.
-We set $\xi$ so that $W_{i}=W\cap C_{i}$ for each $i\in[t]$.
+We set $\xi$ so that $T_{i}=W\cap C_{i}$ for each $i\in[t]$.
 We now proceed to verify the remaining properties of a cut descriptor.
 For CD3, consider $w,w^{\prime}\in W$ with $\xi(w)\neq\xi(w^{\prime})$.
-Let $w\in W_{i},w^{\prime}\in W_{j}$ in $M_{W}$, $i\neq j$, and $w\in C_{i}$, $w^{\prime}\in C_{j}$ in $G$.
+Let $w\in T_{i},w^{\prime}\in W_{j}$ in $M_{W}$, $i\neq j$, and $w\in C_{i}$, $w^{\prime}\in C_{j}$ in $G$.
 Since $S$ separates $C_{i}$ and $C_{j}$ in $G$, and $M$ is a subgraph of $G$, $w$ and $w^{\prime}$ are not adjacent in $M$.
@@
 For the last claim of the lemma, we argue again by construction: since the dissolving operation only affects edges inside $C_{i}$, for some $i$, and dissolving a vertex in $C_{i}$ leaves $C_{i}$ with at least one edge inside, the status of the component (hit/touched/avoided) is always preserved between $M$ and $M_{W}$.
-Specifically, if an edge of $M$ was present in $C_{i}$, then after the dissolving, an edge is still present in $M_{W}$ in $W_{i}$.
-If no edge but some vertices of $M$ are present in $C_{i}$ then all these vertices are terminals or neighbors of $S$, and $|M\cap C_{i}|=|W_{i}|$, and if no vertices of $M$ are in $C_{i}$, then also $W_{i}$ is empty.
+Specifically, if an edge of $M$ was present in $C_{i}$, then after the dissolving, an edge is still present in $M_{W}$ in $T_{i}$.
+If no edge but some vertices of $M$ are present in $C_{i}$ then all these vertices are terminals or neighbors of $S$, and $|M\cap C_{i}|=|T_{i}|$, and if no vertices of $M$ are in $C_{i}$, then also $T_{i}$ is empty.
 ∎
@@
 Then, for a fixed cut descriptor, the algorithm picks an arbitrary assignment fulfilling CD5.
-Since the assignment fixes the vertices of $W$ in $G$, we run the algorithm of Lemma 2 on each $C_{i}$ to find a TM-model connecting the vertices of $W_{i}$ as prescribed by $M_{W}$.
+Since the assignment fixes the vertices of $W$ in $G$, we run the algorithm of Lemma 4 on each $C_{i}$ to find a TM-model connecting the vertices of $T_{i}$ as prescribed by $M_{W}$.
 If any of the runs returns an independent set of size $k$, we output it and stop.
@@
 Output: An independent set of size $k$ in $G$, or a list TM-embedding of $H$ in $G$ respecting $L$, or the output that no such TM-embedding exists.
-1 $\mathcal{M}\leftarrow\emptyset$; for $|W|\leftarrow 0$ to $h+2s$ do
-2       $W\leftarrow$vertex set of size $|W|$; foreach $M_{W}$ a graph on $S\cup W$, $\xi:W\to[t]$, $f_{W}:V(H)\to S\cup W$ do
-3             verify CD1CD4, verify CD5 by a matching instance and pick arbitrary suitable $\tau$; if any of CD1CD5 fail then
+1
+$\mathcal{M}\leftarrow\emptyset$;
+3for $|W|\leftarrow 0$ to $h+2s$ do
+4       $W\leftarrow$vertex set of size $|W|$;
+5      foreach $M_{W}$ a graph on $S\cup W$, $\xi:W\to[t]$, $f_{W}:V(H)\to S\cup W$ do
+6             verify CD1CD4, verify CD5 by a matching instance and pick arbitrary suitable $\tau$;
+7            if any of CD1CD5 fail then
 8                   continue;
 9             end if
-6            $M\leftarrow M_{W}$, $f\leftarrow f_{W}$; for $i\leftarrow 1$ to $t$ do
-7                   $T_{i}\leftarrow\{w\in W:\xi(w)=i\}$, $H_{i}=M_{W}[T_{i}]$; if $H_{i}$ contains no edges then
+10
+11            $M\leftarrow M_{W}$, $f\leftarrow f_{W}$;
+12            for $i\leftarrow 1$ to $t$ do
+13                   $T_{i}\leftarrow\{w\in W:\xi(w)=i\}$, $H_{i}=M_{W}[T_{i}]$;
+14                  if $H_{i}$ contains no edges then
 15                         $O\leftarrow$ empty graph over the vertex set $\tau(T_{i})$;
@@
 18                   end if
+19
 20                  if $O$ is an independent set of size $k$ then
@@
 25                               $f(f_{W}^{-1}(w))\leftarrow\tau(w)$;
-18
+26
 27                         end foreach
-20
+28
 29                   end if
-22
+30
 31             end for
+32
 33            $\mathcal{M}\leftarrow\mathcal{M}\cup\{(M,f)\}$;
 34       end foreach
-26
+35
 36 end for
-28if $\mathcal{M}\neq\emptyset$ then
+37
+38if $\mathcal{M}\neq\emptyset$ then
 39       return $(M,f)\in\mathcal{M}$ with $\max|M|$.
-30else
+40else
 41       return No solution.
@@
 If $H_{i}$ is not empty, a call to Lemma 4 is made in Algorithm 1.
-By 7, the call is valid, and therefore $O$ is either an independent set of size $k$ in $G[C_{i}]$, or the subgraph $M_{i}$ in the desired TM-embedding.
+By 14, the call is valid, and therefore $O$ is either an independent set of size $k$ in $G[C_{i}]$, or the subgraph $M_{i}$ in the desired TM-embedding.
 Since the condition statement before Algorithm 1 explicitly verifies that $O$ is not an independent set of size $k$, in Algorithm 1 the latter case holds.
@@
 Consider an edge between $s\in S$ and $v\in C_{i}$, $i\in[t]$, in $M$.
-By construction of $M$, there exists an edge between $s$ and $w\in W_{i}$ in $M_{W}$ with $\tau(w)=v$.
+By construction of $M$, there exists an edge between $s$ and $w\in T_{i}$ in $M_{W}$ with $\tau(w)=v$.
 By CD5, $N_{M_{W}}(w)\cap S\subseteq N_{G}(\tau(w))$, meaning that $s$ and $v=\tau(w)$ are adjacent in $G$.
@@
 Consider the list TM-embedding $(M^{*},f^{*})$ of $H$ in $G$ of maximum size.
-By 5, there exists a cut descriptor $(W,M_{W},\xi,f_{W})$ with $M_{W}[S]=M^{*}[S]$, and $M^{*}$ hits (touches/avoids) $C_{i}$ if and only if $M_{W}$ hits (touches/avoids) $W_{i}$ for each $i\in[t]$.
+By 12, there exists a cut descriptor $(W,M_{W},\xi,f_{W})$ with $M_{W}[S]=M^{*}[S]$, and $M^{*}$ hits (touches/avoids) $C_{i}$ if and only if $M_{W}$ hits (touches/avoids) $T_{i}$ for each $i\in[t]$.
 Consider the TM-embedding $(M,f)$ constructed by the algorithm at Algorithm 1 with the cut descriptor set to $(W,M_{W},\xi,f_{W})$.
@@
 By construction, $M[S]=M_{W}[S]=M^{*}[S]$.
-Observe also that if $M_{W}$ hits $W_{i}$, then the TM-model $M_{i}$ returned by Lemma 2 spans $C_{i}$, meaning $|M\cap C_{i}|\geq|M^{*}\cap C_{i}|$.
-If $M^{*}$ touches $C_{i}$, then $|M^{*}\cap C_{i}|=|W_{i}|$, and $M$ has at least as many vertices in $C_{i}$.
+Observe also that if $M_{W}$ hits $T_{i}$, then the TM-model $M_{i}$ returned by Lemma 4 spans $C_{i}$, meaning $|M\cap C_{i}|\geq|M^{*}\cap C_{i}|$.
+If $M^{*}$ touches $C_{i}$, then $|M^{*}\cap C_{i}|=|T_{i}|$, and $M$ has at least as many vertices in $C_{i}$.
 If $M^{*}$ avoids $C_{i}$, $|M^{*}\cap C_{i}|=0$.
@@
 Assume a list TM-model of $H$ in $G$ exists.
-Then by 5, there exists also a cut descriptor.
+Then by 12, there exists also a cut descriptor.
 Consider the iteration of the algorithm with this cut descriptor.
@@
 By Lemma 4, up to polynomial factors Algorithm 1 takes time $2^{(h+s+k)^{\mathcal{O}(k)}}$, for each invocation.
-This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 7.
+This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 14.
 Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$.
@@
 We are now ready to show the main result of this section, which we restate for convenience.
-See 1
+See 2
 Proof.
@@
 Output: An independent set of size $k$ in $G$, or a list TM-embedding of $H$ in $G$ respecting $L$, or the output that no such TM-embedding exists.
-1 $S\leftarrow\emptyset$; for $j\leftarrow 1$ to $\infty$ do
-2       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$; if $t\geq k$ then
-3             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$; return $I$.
+1
+[…]
+3for $j\leftarrow 1$ to $\infty$ do
+4       $C_{1}$, …, $C_{t}\leftarrow$ connected components of $G\setminus S$;
+5      if $t\geq k$ then
+6             $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$;
+7            return $I$.
 8       end if
-5      $\kappa_{j}\leftarrow 3\cdot\max\{k+2,10\}\cdot(|V(H)+|E(H)|+|S|)$; if for each $i\in[t]$, $C_{i}$ is $\kappa_{j}$-connected then
+9
+10      $\kappa_{j}\leftarrow 3\cdot\max\{k+2,10\}\cdot(|V(H)+|E(H)|+|S|)$;
+11      if for each $i\in[t]$, $C_{i}$ is $\kappa_{j}$-connected then
 12             break;
 13      else
-8             pick $i$ s.t. $C_{i}$ is not $\kappa_{j}$-connected; find minimum separator $S^{\prime}$ in $G[C_{i}]$; $S\leftarrow S\cup S^{\prime}$;
+14             pick $i$ s.t. $C_{i}$ is not $\kappa_{j}$-connected;
+15            find minimum separator $S^{\prime}$ in $G[C_{i}]$;
+16            $S\leftarrow S\cup S^{\prime}$;
 17       end if
-10
+18
 19 end for
-$O\leftarrow$ invoke Lemma 3 on $G$, $H$, $L$, $k$, $S$; return $O$.
+20
+[…]
+return $O$.
 Algorithm 2 Algorithm of Theorem 2.
@@
 ∎
-3 Proof of Theorem 2: Path cover bellow Gallai-Milgram
-In this section, we prove Theorem 2.
-Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths.
-Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$.
-A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$.
-When we combine this theorem with Theorem 1, we obtain an algorithm solving Path Cover in time time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$.
-Our contribution in Theorem 2 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem.
-We restate Theorem 2 here for the reader’s convenience.
-See 2
-We require two auxiliary results to prove Theorem 2.
-The first result is dealing with small vertex separators of $G$.
-It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
-Lemma 4.
-Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of $G-S$.
-Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
-Proof.
-The construction of $X$ is based on the following observation. We say that a path $P$ in a path cover $\mathcal{P}$ of $G$ is degenerate if $V(P)=C_{i}$ for some $i\in[t]$.
-Claim 13.
-In any minimum path cover $\mathcal{P}$ of $G$, at most $2|S|$ of the sets $C_{1},C_{2},\ldots,C_{t}$ contain vertices of non-degenerate paths of $\mathcal{P}$.
-Proof of Claim 13.
-Notice that if $\mathcal{P}$ has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path by the minimality of $\mathcal{P}$.
-Hence, if $C_{i}$ contains a vertex of a non-degenerate path then $\mathcal{P}$ has a path $P$ with an edge with one endpoint in $S$ and the other in $C_{i}$.
-Each vertex in $S$ is incident to at most two edges in the paths from $\mathcal{P}$.
-Thus, at most $2|S|$ edges are in the path of $\mathcal{P}$ with one endpoint in $S$ and the other in one of $C_{1},C_{2},\ldots,C_{t}$.
-Hence, at most $2|S|$ of the sets can contain vertices of non-degenerate paths.
-∎ To construct $X$, for each vertex $v\in S$, we mark at most $2|S|$ cliques from $\{C_{i}\}_{i\in[t]}$ that contain neighbors of $v$.
-If there are at most $2|S|$ such cliques, then we mark all of them.
-Otherwise, we mark arbitrary $2|S|$ cliques containing neighbors of $v$.
-Then $X$ is defined as the set of indices of the marked cliques.
-It is straightforward that $|X|\leq 2|S|^{2}$ and $X$ can be constructed in polynomial time.
-We claim that there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$.
-In other words, each unmarked clique is covered by a degenerate path.
-Given a path cover $\mathcal{P}$ of $G$, we say that an edge $uv$ of a path $P\in\mathcal{P}$ is bad if $u\in S$ and $v$ is a vertex of an unmarked clique.
-We show that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges.
-The proof is by contradiction.
-We choose $\mathcal{P}$ as a minimum path cover of $G$ that minimizes the total number of bad edges in the paths.
-Suppose that there is $P\in\mathcal{P}$ that contain a bad edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$.
-We have that $C_{i}$ is unmarked and has a neighbor of $u$.
-Then, by the definition of the marking procedure, there are at least $2|S|$ marked cliques with neighbors of $u$.
-Because $C_{i}$ contains a vertex of the non-degenerate path $P$, we obtain that there is a marked clique $C_{j}$ for some $j\in[t]$ with a neighbor $w$ of $u$ such that $C_{j}$ is covered by a degenerate path $P^{\prime}\in\mathcal{P}$ by 13.
-Because $C_{j}$ is a clique, we can assume without loss of generality that $w$ is an end-vertex of $P^{\prime}$.
-Deleting the edge $uv$ separates $P$ into two paths $Q$ and $Q^{\prime}$ having end-vertices $v$ and $u$, respectively.
-We use the fact that $uw\in E(G)$ and construct the path $\hat{Q}$ from $Q^{\prime}$ and $P^{\prime}$ by joining their end-vertices $u$ and $w$ by the edge $uw$.
-Notice that $Q$ and $\hat{Q}$ are vertex-disjoint and $V(Q)\cup V(\hat{Q})=V(P)\cup V(P^{\prime})$.
-This allows us to construct the path cover $\mathcal{P}^{\prime}$ by replacing $P$ and $P^{\prime}$ in $\mathcal{P}$ by $Q$ and $\hat{Q}$.
-Because the number of paths in $\mathcal{P}^{\prime}$ is the same as in $\mathcal{P}$, $\mathcal{P}^{\prime}$ is a minimum path cover.
-However, the number of bad edges in the path of $\mathcal{P}^{\prime}$ is less than the number of bad edges in the paths of $\mathcal{P}$, contradicting the choice of $\mathcal{P}$.
-This proves that there is a minimum path cover $\mathcal{P}$ of $G$ that does not contain paths with bad edges.
-To complete the proof, we remind that if a minimum path cover has no path with an edge $uv$ with $u\in S$ and $v\in C_{i}$ for some $i\in[t]$ then the clique $C_{i}$ is covered by a degenerate path.
-Thus, there is a minimum path cover $\mathcal{P}$ of $G$ such that for each $i\in[t]\setminus X$, there is a path $P\in\mathcal{P}$ satisfying $V(P)=C_{i}$.
-∎
-Our following auxiliary result is a “lighter version” of Theorem 2. It deals with the particular case of $\operatorname{pc}(G)<k^{C}$.
-Lemma 5.
-There is an algorithm with running time $2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$ that, given an $n$-vertex graph $G$, integer $k\geq 0$ and a path cover of $G$ with at most $\mathcal{O}(k^{C})$ paths for constant $C\geq 1$, outputs a path cover $\mathcal{P}$ of $G$ and
-  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover, or
-  • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
-Proof.
-We present an algorithm $\mathcal{A}$.
-The algorithm assumes that $G$ contains at least one edge as, otherwise, $pc(G)=n$ and the minimum-size path cover consists of trivial paths.
-Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$.
-To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$.
-On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching.
-Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial.
-The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 1 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$.
-The subroutine works in
-$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}}\cdot 2^{(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
-running time.
-If the outcome of the subroutine is an independent set $I$ in $G$ of size $p^{\prime}+k$, then the algorithm outputs $\mathcal{P}^{\prime}$ and $I$ as the solution.
-Since $|\mathcal{P}^{\prime}|=p^{\prime}$, $|I|\geq|\mathcal{P}^{\prime}|+k$ is satisfied, so this solution satisfies the lemma statement.
-If the subroutine decides that $H_{ij}$ does not have TM-model in $G$ or the maximum TM-model of $H_{ij}$ in $G$ is not spanning, then $G$ cannot be covered with $i$ paths and $\mathcal{A}$ continues to the next iteration.
-If the subroutine gives a spanning TM-model of $H_{ij}$ in $G$, then the minimum path cover size of $G$ equals $i$ and $\mathcal{A}$ transforms this TM-model into a path cover of $G$ with $i$ paths.
-Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops.
-Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$.
-The proof of the lemma is complete.
-∎
-We are ready to prove the main result of this section.
-Proof of Theorem 2.
-We describe an algorithm that starts from a trivial path cover of $G$ and then gradually transforms it.
-These transformations are indicated as reduction rules and are performed in polynomial time.
-We show that each such rule does not increase the number of paths in the path cover.
-Moreover, each reduction rule necessarily reduces either the total number of paths or the number of special (to be defined later) paths in the path cover.
-A series of consecutive applications of the reduction rules always lead the algorithm to cases when the algorithm can produce a solution satisfying theorem statement in $f(k)\cdot n^{\mathcal{O}(1)}$ time and stop.
-We indicate such cases and the corresponding behavior of the algorithm as solution subroutines.
-The algorithm runtime thus always ends with a solution subroutine run.
-In what follows, we present the reduction rules and solution subroutines.
-Each of the reduction rules reduces the number of paths in a path cover or reduces the number of some special paths.
-Solution subroutines output path covers and independent sets.
-The rules and subroutines are given in the order in which the algorithm checks their applicability.
-That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied.
-When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties.
-In this case, the final solution subroutine calls the algorithms of Lemmas 4 and 5.
-The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 00.
-Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$.
-Let the current set of paths be $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{m}\}$.
-We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
-The first reduction rule is straightforward and comes from the proof of the Gallai-Milgram theorem for undirected graphs.
-Reduction Rule 1.
-If there is an edge $uv$ in $G$ such that $u\in\{s_{i},t_{i}\}$ and $v\in\{s_{j},t_{j}\}$ and $i\neq j$, then join $P_{i}$ and $P_{j}$ into a single path via $uv$.
-The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one.
-Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$.
-When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
-Special and usual paths.
-We call a path $P_{i}$ in the path cover special, if either $s_{i}=t_{i}$, or $P_{i}$ consists of only edge $s_{i}t_{i}$, or its endpoints, are connected by the edge $s_{i}t_{i}$.
-If a path $P_{i}$ is not special, we call it usual.
-We claim that sufficiently many usual paths give a solution to the problem.
-Claim 14.
-If at least $k$ paths are usual in the path cover $\mathcal{P}=P_{1},\ldots,P_{m}$ and 1 is not applicable, then $m\leq\alpha(G)-k$.
-Moreover, an independent set $I$ in $G$ of size $m+k$ is computable in polynomial time.
-Proof of Claim 14.
-We show that $G$ has an independent set $I$ of size $m+k$.
-Initialize $I$ with $\{s_{1},s_{2},\ldots,s_{p}\}$.
-Since 1 is not applicable, $I$ is independent in $G$.
-Let $i_{1},i_{2},\ldots,i_{k}$ be $k$ distinct indices of usual paths.
-Since each $P_{i_{j}}$ is usual, $t_{i_{j}}\neq s_{i_{j}}$ and $s_{i_{j}}$ is not connected with $t_{i_{j}}$.
-Also, 1 is not applicable, so there is no edge between $t_{i_{j}}$ and any vertex in $I$, while $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}\}$ is also an independent set in $G$.
-Put $t_{i_{1}},t_{i_{2}},\ldots,t_{i_{k}}$ in $I$.
-We obtain an independent set of size $m+k$ in $G$, consequently $\alpha(G)-k\geq m$.
-∎ 14 yields the corresponding solution subroutine for the algorithm.
-Solution Subroutine 1.
-If there are at least $k$ usual paths among $\mathcal{P}=P_{1},P_{2},\ldots,P_{m}$, output $\mathcal{P}$ and $I$ as in 14.
-From now on, we assume that $\mathcal{P}$ has less than $k$ usual paths.
-The remaining reduction rules deal primarily with special paths.
-The following rule reduces the total number of paths given that $G$ has an edge between any vertex of a special path and an endpoint of an arbitrary other path.
-Reduction Rule 2.
-If $P_{i}$ is a special path of length at least two and there is an edge in $G$ connecting some $u\in V(P_{i})$ with $s_{j}$ (or $t_{j}$) for some $j\neq i$, then
-  1. 1.
-    Add edge $s_{i}t_{i}$ to $P_{i}$, so it becomes a cycle;
-  2. 2.
-    Remove edge $uv$ from $P_{i}$, where $v$ is any neighbor of $u$ in $P_{i}$, so it becomes an $uv$-path;
-  3. 3.
-    Join $P_{i}$ and $P_{j}$ via $us_{j}$ (or $ut_{j}$).
-2 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, so $\mathcal{P}$ remains a path cover implying that the rule is correct.
-The following reduction rule deals with an edge between two vertices of distinct special paths.
-It is demonstrated in Figure 1a.
-Reduction Rule 3.
-If $P_{i}$ and $P_{j}$ are special paths of length at least two for $i\neq j$, and there is an edge in $G$ connecting $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$, then
-  1. 1.
-    Add edge $s_{j}t_{j}$ to $P_{j}$, so it becomes a cycle;
-  2. 2.
-    Remove edge $u_{j}v_{j}$ from $P_{j}$, so it becomes an $u_{j}v_{j}$-path.
-  3. 3.
-    Apply 2 to $P_{i}$, $P_{j}$ and $u_{i}u_{j}$.
-Again, 3 replaces $P_{i}$ and $P_{j}$ with a path covering $V(P_{i})\cup V(P_{j})$, thus reducing the number of paths in $\mathcal{P}$. We claim the following.
-Claim 15.
-If Reduction Rules 1,2, and 3 are not applicable, then for each pair of distinct special paths $P_{i},P_{j}$ in $\mathcal{P}$, there is no edge between $V(P_{i})$ and $V(P_{j})$ in $G$.
-Proof of Claim 15.
-Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$.
-If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable.
-Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$.
-If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction.
-Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$.
-In this case, 1 is applicable, leading us to the final contradiction.
-∎ The claim is the basis for the next solution subroutine, which the algorithm applies next.
-We explain how it works by proving the following claim.
-Claim 16.
-If 1 and Reduction Rules 1,2, and 3 are not applicable, and there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then $m<\alpha(G)-k$.
-Moreover, an independent set $I$ in $G$ of size more than $m+k$ is computable in polynomial time.
-Proof of Claim 16.
-We construct an independent set $I$ consisting of more than $m+k$ vertices.
-Since 1 is not applicable, there are more than $m-k$ special paths.
-For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$.
-For each other special path, put its arbitrary vertex into $I$.
-By 15, $I$ is an independent set in $G$.
-Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$.
-∎ The solution subroutine based on 16 is given below.
-The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
-Solution Subroutine 2.
-If there are at least $2k$ special paths $P_{i}$ such that $V(P_{i})$ is not a clique in $G$, then output $\mathcal{P}$ and an $I$ as in 16.
-The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$.
-The following reduction rule necessary reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
-a $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$ b $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$$s_{\ell}$$t_{\ell}$$x_{i}$$x_{j}$
-Figure 1: Illustration of transformations performed in (a) 3 and (b) 4..
-Paths before the transformations are horizontal.
-Paths after the transformations are highlighted in bold and have pairwise-distinct colors.
-Dashed lines represent absence of an edge.
-Snake lines represent paths instead of single edges.
-Reduction Rule 4.
-If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$ and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
-  1. 1.
-    Make $u_{i}$ endpoint of $P_{i}$ and $u_{j}$ endpoint of $P_{j}$ similarly to the first two steps of 2;
-  2. 2.
-    Break $P_{\ell}$ into three (or two) parts: the prefix part $s_{\ell}P_{\ell}x_{i}$, the suffix part $x_{j}P_{\ell}t_{\ell}$, and (if exists) the remaining middle part;
-  3. 3.
-    Join $P_{i}$ with the prefix part via $u_{i}x_{i}$;
-  4. 4.
-    Join $P_{j}$ with the suffix part via $u_{j}x_{j}$.
-Figure 1b illustrates the result of application of 4. The following claim addresses its correctness.
-Claim 17.
-If Reduction Rules 1, 2 are not applicable, then 4 reduces the number of special paths in $\mathcal{P}$ and does not increase the number of paths in $\mathcal{P}$.
-Proof of Claim 17.
-The second part of the claim is trivial, since 4 produces three (or two, if $x_{i}x_{j}\in E(P_{\ell})$) paths by transforming three distinct paths $P_{i},P_{j},P_{\ell}$, so $|\mathcal{P}|$ cannot increase.
-To prove the first part, we show that the resulting $v_{i}s_{\ell}$-path and $v_{j}t_{\ell}$-path are usual.
-We denote these two paths by $Q_{i}$ and $Q_{j}$ correspondingly.
-We give the proof only for $Q_{i}$, since the proof for $Q_{j}$ is symmetric.
-We first show that $|V(Q_{i})|\geq 3$.
-Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$.
-The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly.
-Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold.
-But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$.
-Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$.
-This contradiction proves $|V(Q_{i})|\geq 3$.
-It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$.
-Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable.
-This contradiction proves the claim.
-∎
-The list of reduction rules of the algorithm is exhausted.
-Starting from this point, the algorithm exploits the structure of the graph, assuming that no reduction rule is applicable.
-In what follows, we explain these structural properties.
-We also give the last solution subroutine of the algorithm.
-Claim 18.
-If none of the Reduction Rules 1-5 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$.
-Moreover, such a set $S$ could be constructed in polynomial time.
-Proof of Claim 18.
-Since 14 is not applicable, there are less than $k$ usual paths in $\mathcal{P}$.
-We construct $S$ by taking one or zero vertices from each usual path, so $|S|<k$ holds automatically.
-We now explain the choice of the vertex to put in $S$ for each usual path.
-Consider a usual path $P_{i}\in\mathcal{P}$.
-We call a vertex $u\in V(P_{i})$ a connector if there is a special path $P_{j}$ and a vertex $v\in V(P_{j})$ with $uv\in E(G)$.
-If $P_{i}$ has exactly one connector, then we put this connector in $S$.
-Otherwise, i.e.
-when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$.
-We show that at least $m-2k$ special paths that are isolated in $G-S$.
-First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded.
-Assume now that $P_{j}$ is a special path that is not isolated in $G-S$.
-Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$.
-By 15, $P_{i}$ can only be usual.
-Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors.
-As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$.
-Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$.
-Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$.
-∎
-We are ready to present the final solution subroutine which combines all algorithmic results given previously in this section.
-Solution Subroutine 3.
-Perform the following steps:
-  1. 1.
-    Obtain $S$ according to 18 and put $h:=\max\{0,m-4k\}$;
-  2. 2.
-    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in $G-S$;
-  3. 3.
-    Apply the algorithm of Lemma 4 to $G$, $S$ and $V(P_{1}),V(P_{2}),\ldots,V(P_{h})$ and obtain a set $X$ of at most $2k^{2}$ indices, where each $i\in X$ satisfies $i\in[h]$;
-  4. 4.
-    Obtain $G^{\prime}$ and $\mathcal{P}^{\prime}$ by removing $V(P_{j})$ from $G$ and $P_{j}$ from $\mathcal{P}$, for each $j$ such that $j\in[h]$ and $j\notin X$;
-  5. 5.
-    Apply the algorithm of Lemma 5 as a subroutine to $G^{\prime},\mathcal{P}^{\prime}$ and $k^{\prime}:=2k$ and obtain a path cover $\mathcal{S}^{\prime}$ of $G^{\prime}$ and (possibly) an independent set $I^{\prime}$ in $G^{\prime}$;
-  6. 6.
-    Obtain a path cover $\mathcal{S}$ of $G$ by adding all paths in $\mathcal{P}\setminus\mathcal{P}^{\prime}$ to $\mathcal{S}^{\prime}$;
-  7. 7.
-    If $I^{\prime}$ was obtained, remove all vertices of $S$ from $I^{\prime}$ and for each path $P_{j}\in\mathcal{P}\setminus\mathcal{P}^{\prime}$, add one vertex of $P_{j}$ to $I^{\prime}$.
-Denote the resulting set by $I$;
-  8. 8.
-    Output $\mathcal{S}$ and (if obtained) $I$.
-We conclude the proof with a claim certifying the correctness of the solution subroutine.
-In contrast with all previous rules and subroutines, 3 requires superpolynomial (but $\operatorname{{\sf FPT}}$ in $k$) running time.
-Claim 19.
-If Reduction Rules 1-5 and Solution Subroutines 1,2 are not applicable, then 3 works in $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$ time and produces a feasible solution.
-Proof of Claim 19.
-First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 18 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
-We now move on to establish the upper bound on the running time.
-All steps, except for applying Lemma 5, are carried out in polynomial time.
-We have
-$|\mathcal{P}^{\prime}|=|\mathcal{P}|-h+|X|=m-h+|X|\leq 4k+2k^{2}=\mathcal{O}(k^{2}).$
-By applying Lemma 5 to $G^{\prime},k^{\prime}$ and $\mathcal{P}^{\prime}$, we obtain the running time $2^{k^{\mathcal{O}(k^{4})}}\cdot|G|^{\mathcal{O}(1)}$.
-Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$.
-Hence it is also a path cover.
-On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$.
-Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$.
-First, let us consider the scenario where $I$ is not produced.
-This can occur only when the algorithm of Lemma 5 does not produce $I^{\prime}$.
-Then Lemma 5 guarantees that $\mathcal{S^{\prime}}$ is an optimal vertex cover of $G^{\prime}$ and $|\mathcal{S}^{\prime}|=\operatorname{pc}(G^{\prime})$.
-By Lemma 4, removal of $V(P_{j})$ for any $j\in[h]\setminus X$ decreases $\operatorname{pc}(G)$ by one, so $\operatorname{pc}(G^{\prime})=\operatorname{pc}(G)-(h-|X|).$ Since $|\mathcal{P}\setminus\mathcal{P}^{\prime}|=h-|X|$, we have $|\mathcal{S}|=\operatorname{pc}(G)$.
-It means that $\mathcal{S}$ is an optimal path cover of $G$.
-Finally, let us consider the situation where $I$ is produced.
-In this case, the algorithm in Lemma 5 provides $\mathcal{S^{\prime}}$ and $I^{\prime}$ with $|I^{\prime}-|\mathcal{S}^{\prime}|=k^{\prime}=2k$.
-We notice that
-$|I|\geq|I^{\prime}|-|S|+|\mathcal{P}\setminus\mathcal{P}^{\prime}|>|I^{\prime}|-k+(h-|X|).$
-As discussed in the previous paragraph, $|\mathcal{S}|-|\mathcal{S}^{\prime}|=h-|X|$. We conclude that
-$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)+(h-|X|)=k.$
-The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
-This completes the proof of Theorem 2. ∎
 5 Conclusion
@@
 But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$.
-Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [${}^{+}$11, ${}^{+}$19, FGLS10, LMS18].
+Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [CNP+11, FGL+19, FGLS10, LMS18].
 However, all these reductions use gadgets with large independence numbers.

### v3

@@
-Fedor V. Fomin
+Path Cover, Hamiltonicity, and Independence Number: An FPT Perspective
 Abstract
-The classic theorem of Gallai and Milgram (1960) asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$.
-The proof of Gallai–Milgram’s theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths.
-We prove the following algorithmic extension of Gallai–Milgram’s theorem for undirected graphs: determining whether an undirected graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is fixed-parameter tractable ($\mathsf{FPT}$) when parameterized by $k$.
-More precisely, we provide an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time
-$2^{k^{O(k^{4})}}\cdot n^{O(1)},$
-and outputs a path cover $\mathcal{P}$ of $G$. Furthermore, it:
-  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover,
-  • or outputs, together with $\mathcal{P}$, an independent set of size $\lvert\mathcal{P}\rvert+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
-A key subroutine in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path.
-This result is of independent interest—prior to our work, no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most 3.
-Moreover, the techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment.
+The classic theorem of Gallai and Milgram (1960) generalizes several fundamental results in Graph Theory, such as Dilworth’s theorem on posets and Kőnig’s theorem on matchings in bipartite graphs.
+The theorem asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$.
+The proof of the Gallai–Milgram theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths.
+While the Gallai–Milgram theorem is tight—there are graphs where one really needs $\alpha(G)$ paths, not fewer, to cover the vertex set of GG—it was not known prior to our work whether deciding if a graph $G$ could be covered by fewer than $\alpha(G)$ vertex-disjoint paths can be done in polynomial time.
+We resolve this question by proving the following algorithmic extension of the Gallai–Milgram theorem for undirected graphs: There is an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time $2^{2^{\mathcal{O}(k^{4}\log{k})}}\cdot n^{\mathcal{O}(1)}$ and outputs a path cover $\mathcal{P}$ of $G$ together with
+  • a correct conclusion that $\mathcal{P}$ is a minimum-size path cover, or
+  • an independent set of size $\lvert\mathcal{P}\rvert+k$, certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+Thus, for $k\in\mathcal{O}((\log\log n)^{\frac{1}{4}-\varepsilon})$ our algorithm runs in polynomial time and either computes a minimum-size path cover of $G$, or finds a path cover of size at most $\alpha(G)-k$.
+We find the existence of such an algorithm quite surprising for the following reason.
+The problems of computing a path cover and a maximum independent set are both notoriously hard, yet our algorithm either solves one of them or provides meaningful information about the other.
+The proof of our algorithmic extension of the Gallai–Milgram theorem is non-trivial and builds on several novel algorithmic ideas.
+One of the key subroutines in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path.
+This result is of independent interest—prior to our work no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most three.
+Moreover, the algorithmic techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment.
 We show that all these problems are $\mathsf{FPT}$ when parameterized by the independence number of the graph.
-Notably, the independence number parameterization, which describes graph’s density, departs from the typical flow of research in parameterized complexity, which focuses on parameters describing graph’s sparsity, like treewidth or vertex cover.
+Notably, the independence-number parameterization departs from the typical direction of research in parameterized complexity.
+First, $\alpha(G)$ measures a graph’s density, whereas most prior work in the area focuses on parameters describing sparsity, such as treewidth or vertex cover.
+Second, most structural parameters studied in parameterized complexity can be computed exactly or well-approximated in polynomial or even $\mathsf{FPT}$ time, whereas computing $\alpha(G)$ is notoriously difficult from almost any computational perspective.
+The fact that it can nevertheless serve as the basis for efficient parameterization is particularly striking.
 1 Introduction
-The Gallai–Milgram Theorem, published by Tibor Gallai and Arthur Milgram in 1960 [GM60], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s Theorem [Dil50], concerning the covering of graphs by simpler substructures.
+Gallai–Milgram’s theorem, published by Tibor Gallai and Arthur Milgram in 1960 [45], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s and Kőnig’s theorems [29, 68], concerning the covering of graphs.
 In its original form, the theorem states that every (directed or undirected) graph $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the size of a maximum independent set in $G$.
-Beyond its immediate implications, the Gallai–Milgram Theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence.
-This theme can be traced back to the seminal work of Rédei [Réd34], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path.
-Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [NW71] and Chvátal–Erdős [CE72], and have continued in later works (see, e.g., [AFG83, BBVR89, Jac87, Koi94, BAH88, BT07, HJ90, Tho01, FS09]).
-In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among the Gallai–Milgram theorem, Hamiltonicity, and the independence number.
-1.1 Our contribution.
-Known constructive proofs of the Gallai–Milgram theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths.
-Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
+Beyond its immediate implications, Gallai–Milgram’s theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence.
+This theme can be traced back to the seminal work of Rédei [85], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path.
+Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [82] and Chvátal–Erdős [19], and have continued in later works (see, e.g., [3, 5, 58, 67, 6, 8, 57, 88, 43]).
+In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among Gallai–Milgram’s theorem, Hamiltonicity, and the independence number.
+1.1 Our contribution
+Known constructive proofs of Gallai–Milgram’s theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths.
+Our following theorem presents a significant generalization of this algorithmic result for undirected graphs.
 Theorem 1.
@@
   • outputs together with $\mathcal{P}$ an independent set of size $|\mathcal{P}|+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+Loosely speaking, the theorem means that computing a path cover below Gallai–Milgram’s bound, that is, a path cover of size at most $\alpha(G)-k$, is $\operatorname{{\sf FPT}}$ when parameterized by $k$.
+However, formally, such a statement is not correct: A trivial reduction shows that deciding whether $G$ has a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf W}}[1]$-hard.
+Indeed, for an $n$-vertex input graph $G$, construct a graph $G^{\prime}$ by adding a clique of size $n$ whose vertices are made universal to $G$.
+Then $\alpha(G^{\prime})=\alpha(G)$, and $G^{\prime}$ is trivially Hamiltonian.
+Thus, for $k^{\prime}=k-1$, the graph $G^{\prime}$ can be covered by at most $\alpha(G^{\prime})-k^{\prime}$ paths if and only if $\alpha(G)\geq k$.
+Since the Independent Set problem is $\operatorname{{\sf W}}[1]$-complete [30] when parameterized by the solution size, the existence of an $\operatorname{{\sf FPT}}$ algorithm in $k$ for deciding whether $G$ has a path cover of size at most $\alpha(G)-k$ would contradict the basic complexity assumptions of parameterized complexity theory.
+Of course, in situations where the independence number of a graph can be efficiently computed, for example when the input graph is perfect, Theorem 1 implies that the path cover below Gallai–Milgram is $\operatorname{{\sf FPT}}$.
+The interesting feature of Theorem 1 is that it circumvents the intractability of computing $\alpha(G)$—the algorithm does not need to know $\alpha(G)$ at all.
+This is achieved by allowing the computation of an optimal path cover rather than concluding that $G$ has no path cover of size at most $\alpha(G)-k$.
+Another notable feature of the algorithm is that it always outputs a path cover which is either optimal or of size at most $\alpha(G)-k$.
+It is somewhat ironic that the “formal” intractability of the path cover below Gallai–Milgram arises only from the requirement of identifying no-instances, that is, situations where $\alpha(G)-k$ is smaller than the size of a minimum path cover.
 Theorem 1 aligns with a broader trend in parameterized algorithms that seeks algorithmic analogues of classic theorems in extremal graph theory.
-Examples include algorithmic extensions of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem by Gan, Han, and Hu [GHH23].
+Examples include algorithmic extensions of Poljak–Turzík’s theorem [80], Dirac’s theorem [37, 60], and Hajnal–Szemerédi’s theorem [46].
 As with Theorem 1, each of these results provides an $\operatorname{{\sf FPT}}$ algorithm parameterized “above” or “beyond” a known combinatorial bound.
-A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [DF99] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
-Consequently, we cannot simply require $\alpha(G)$ to be part of the input, which makes the existence of an $\operatorname{{\sf FPT}}$ algorithm parameterized below $\alpha(G)$ all the more surprising.
-As a special case of Theorem 1, let $k=\alpha(G)-1$.
-In this scenario, the problem reduces to determining whether all vertices of $G$ can be covered by a single path—i.e., whether $G$ has a Hamiltonian path.
-Consequently, Theorem 1 implies that deciding whether $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$ when parameterized by $\alpha(G)$.
-The construction of such an $\operatorname{{\sf FPT}}$ algorithm for Hamiltonicity is a key component of the proof of Theorem 1.
-Moreover, the techniques developed to achieve this result extend well beyond Hamiltonicity and apply to a broader class of problems.
-Similar to Theorem 1, the situation concerning parameterization by $\alpha(G)$ is unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$.
-For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning.
-First, for an input integer $k$, there are (either exact or approximation) algorithms that either correctly decide that the treewidth of the input graph is more than $k$ or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [Bod96, Kor21, RS86].
-Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$.
-In contrast, this time we deal with the parameter that does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
-Our second theorem provides an $\operatorname{{\sf FPT}}$ algorithm for computing topological minors in a graph, parameterized by its independence number.
+A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [30] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [71, 63].
+One of the key components in the proof of Theorem 1 is an $\operatorname{{\sf FPT}}$ algorithm, parameterized by $\alpha(G)$, for deciding the Hamiltonicity of $G$.
+Moreover, the techniques developed to obtain this result extend well beyond Hamiltonicity and apply to a broader class of problems.
+Our second main result, Theorem 2, provides an $\operatorname{{\sf FPT}}$ algorithm for computing various topological embeddings in a graph, parameterized by its independence number.
+While only a restricted case of this result is needed in the proof of Theorem 1, we believe that the general approach is of independent interest, and therefore present it in full generality.
 Let us recall that a graph $H$ is called a topological minor of a graph $G$ if there exists a subgraph of $G$ that is isomorphic to a subdivision of $H$.
@@
 Definition 1 (TM-model and TM-embedding).
-Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of $M\setminus T$ have degree $2$.
-We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$.
+Let $G$, $H$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of MTM-T have degree two in $M$.
+We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to the graph obtained from $M$ by dissolving in $M$ all vertices of MTM-T.
 Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent.
@@
 For graphs $H$, $G$, and a mapping $L:V(H)\to 2^{V(G)}$, we say that TM-embedding $M$, $f:V(H)\to V(G)$ is a list topological minor embedding (list TM-embedding), if for each $h\in V(H)$, $f(h)\in L(h)$.
-The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
+The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$, that is, solving the Maximum List TM-Embedding problem, parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
 The algorithm, for a given parameter $k$, either computes a TM-embedding of maximum size in $\operatorname{{\sf FPT}}$ time with respect to $k$, or correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$.
-Such algorithms are what Raghavan and Spinrad referred to as robust [RS03].
+Such algorithms are what Raghavan and Spinrad refer to as robust [83].
 More precisely,
 Theorem 2.
-There is an algorithm that, given graphs $H,G$, a list assignment $L:V(H)\to 2^{V(G)}$ and an integer parameter $k\geq 1$, in time $2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$
+There is an algorithm that, given graphs $H,G$, a list assignment $L:V(H)\to 2^{V(G)}$ and an integer parameter $k\geq 1$, in time $2^{|H|^{\mathcal{O}(k)}\cdot k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$
   • either correctly reports that there is no list TM-embedding of $H$ in $G$, or
@@
   • outputs an independent set in $G$ of size $k$.
+Similar to Theorem 1, the question concerning parameterization by $\alpha(G)$ deviates from the standard setup used in structural parameterized complexity, where most algorithmic results rely on the fact that the problem of computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$.
+For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ when parameterized by the treewidth of the graph, for the following reason.
+Given an input integer $k$, there exist algorithms (either exact or approximate) that either correctly decide that the treewidth of the input graph is greater than $k$, or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [12, 69, 86].
+The resulting tree decomposition can then be used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$.
+In contrast, in our setting, the parameter under consideration does not admit an $\operatorname{{\sf FPT}}$ approximation [71, 63], which necessitates a different approach.
+Theorem 2 implies that Maximum List TM-Embedding is $\operatorname{{\sf FPT}}$ with respect to the combined parameter $|H|+\alpha(G)$.
+For any integer $k>\alpha(G)$, the algorithm of Theorem 2 either outputs a maximum-size list TM-embedding or correctly determines that no such embedding exists.
+The fact that $\alpha(G)$ cannot be computed efficiently can be circumvented by the following trick.
+To find a list TM-embedding of $H$ in $G$, we run the algorithm of Theorem 2 for increasing values of the parameter $k$, starting with $k=1$, until the algorithm either finds an embedding or concludes that no appropriate embedding exists.
 To our knowledge, neither list TM-embeddings nor the Maximum List TM-Embedding problem have been studied so far.
-However, this problem is a natural generalization of several vital problems in graph algorithms.
+However, these problems are natural generalizations of several vital problems in graph algorithms.
 In particular, Theorem 2 implies the following corollaries.
@@
 For $H=C_{3}$, we have the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$.
-By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$.
+By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ which implies that the problem of finding a longest path/cycle is $\operatorname{{\sf FPT}}$ parameterized by $\alpha(k)$.
+  • A graph $G$ is Hamilton-connected (also known as Hamiltonian connected) if every two vertices of $G$ are connected by a Hamiltonian path; see, e.g., [16, p.
+61].
+Checking whether a graph is Hamilton-connected thus reduces to computing, for each pair $u,v\in V(G)$, a maximum-size TM-embedding of $H=P_{2}$ with lists $\{u\}$ and $\{v\}$, which can be done in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ by Theorem 2.
   • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths.
 By Gallai–Milgram’s theorem, if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$.
-We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$.
+We can solve the problem by taking $H$ to be disjoint unions of at most $p$ disjoint copies of $K_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$.
 Thus, Theorem 2 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem.
@@
   • Consider the following generalization of Path Cover.
-In the $-\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$.
-The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph.
-For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path.
-$-\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$.
+In the Hamiltonian-$\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$.
+The task is to connect these pairs by disjoint paths that altogether traverse all the vertices of the graph.
+For $\ell=1$, this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path.
+Hamiltonian-$\ell$-Linkage could be encoded as Maximum List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$.
 Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$.
 Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$.
-By Theorem 2, $-\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$.
-By playing with different graphs $H$ in the List TM-embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
+By Theorem 2, Hamiltonian-$\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$.
+By playing with different graphs $H$ in Maximum List TM-Embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
   • Another interesting scenario is the maximization variant of the $T$-Cycle problem.
 In the $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals.
-The task is to decide whether there is a cycle passing through all terminals [BHT12, Kaw08, Wah13] (in [BHT12] authors look for the minimum-length cycle).
+The task is to decide whether there is a cycle passing through all terminals [11, 64, 89] (in [11] authors look for the minimum-length cycle).
 We can reduce the maximization variant to finding a list TM-embedding of maximum size by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution.
 For each guessed ordering, we apply Theorem 2 for $H$ being a cycle on $|T|$ vertices such that each vertex of $H$ is assigned to precisely one terminal vertex, and the cyclic ordering of $H$ corresponds to the guess.
-This brings us to the $f(\alpha(G)+|T|)\cdot|G|^{\mathcal{O}(1)}$ algorithm computing a cycle of maximum length containing all $T$ terminals.
+This brings us to the $f(\alpha(G)+|T|)\cdot|G|^{\mathcal{O}(1)}$ algorithm computing a cycle of maximum length containing all terminals in $T$.
 Similar arguments hold for a more general colored variant of the problem.
 In this variant, we are given a graph $G$, and some of the vertices of $G$ are assigned colors from $1$ to $L$.
-Then, the task is to find a longest cycle containing the maximum number of colors [FGK+23].
+Then, the task is to find a longest cycle containing the maximum number of colors [33].
 Note that an optimal cycle can contain many vertices of the same color.
@@
 This brings us to the algorithm of running time $f(\alpha(G)+L)\cdot|G|^{\mathcal{O}(1)}$.
+We remark that the result of Theorem 2 is tight in the sense that Maximum List TM-Embedding is trivially $\operatorname{{\sf Para-NP}}$-hard when parameterized by either $\alpha(G)$ or $|H|$ only.
+The problem is $\operatorname{{\sf NP}}$-hard for $H=K_{2}$ because of the $\operatorname{{\sf NP}}$-completeness of Hamiltonian Path [47].
+Furthermore, observe that finding a path with $n$ vertices in a bipartite $n$-vertex graph $G$ is equivalent to finding the complement of the $n$-vertex path in the complement of $G$ which is cobipartite.
+Since Hamiltonian Path is $\operatorname{{\sf NP}}$-complete on bipartite graphs [47] and the independence number of a cobipartite graph is at most two, Maximum List TM-Embedding is $\operatorname{{\sf NP}}$-hard for graphs $G$ with $\alpha(G)\leq 2$ when $H$ is a part of the input.
 1.2 Overview of Theorems 1 and 2
 The proof of Theorem 1 uses Theorem 2 as a black box. We start with sketching the proof of Theorem 1.
-Overview of the proof of Theorem 1: To sketch the proof of Theorem 1, we first recall the standard proof of the Gallai–Milgram theorem for undirected graphs.
-This proof hinges on a simple observation: if a path cover of $G$ consists of more than $\alpha(G)$ paths, then there is an edge connecting the endpoints of two different paths.
-Merging these two paths into one reduces the cover by one path.
-Reiterating this step yields a polynomial-time algorithm that produces a path cover of size at most $\alpha(G)$.
-Notably, the algorithm does not need to know $\alpha(G)$ in advance; it merely keeps merging pairs of paths with adjacent endpoints until no such pair remains.
-The procedure can begin from the trivial cover formed by $\lvert V(G)\rvert$ single-vertex paths.
-The procedure above does not, by itself, provide any additional relationship between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ beyond $p\leq\alpha(G)$.
-To address this issue, we introduce the notions of usual and special paths.
+Overview of the proof of Theorem 1: To sketch the proof of Theorem 1, we first recall the standard argument for Gallai–Milgram’s theorem in undirected graphs.
+The key observation is that if a path cover of $G$ has more than $\alpha(G)$ paths, then there must be an edge connecting the endpoints of two distinct paths.
+By merging these paths, the cover size decreases by one.
+Thus, starting from the trivial cover consisting of $\lvert V(G)\rvert$ single-vertex paths, we can iteratively obtain a path cover of size at most $\alpha(G)$.
+This procedure runs in polynomial time.
+Importantly, we do not need to know $\alpha(G)$ in advance; the algorithm simply keeps merging paths with adjacent endpoints until no such pair remains.
+The merging procedure used to prove Gallai–Milgram’s theorem does not provide any additional information beyond $p\leq\alpha(G)$, where $p$ is the size of the constructed path cover $\mathcal{P}$.
+To move beyond this bound, we extract more structural information from a path cover once the Gallai–Milgram merging process can no longer continue.
+For this purpose, we introduce the notions of usual and special paths.
 A path is called usual if its endpoints are distinct and no edge connects them in $G$; otherwise, the path is special.
-We argue that having usual paths in $\mathcal{P}$ is advantageous.
-For instance, if all paths in $\mathcal{P}$ are usual, then the endpoints of these paths form an independent set of size $2p$.
-(Otherwise, we could merge two paths whose endpoints are adjacent.) In particular, this implies $p\leq\alpha(G)/2$.
-Even having $k$ usual paths is sufficient, since their endpoints would then form an independent set of size $p+k$.
-Alongside $\mathcal{P}$, this independent set would constitute a feasible output for the algorithm in Theorem 1.
-While special paths are not helpful for constructing a large enough independent set, they are convenient for merging.
-A special path is either a path of length zero or one, or it can be viewed (along with the edge between its endpoints) as a cycle of length at least three.
-Then, an edge between any two vertices of different special paths is as effective as an edge between the endpoints: it enables us to merge those two paths and cover all their vertices with a single path.
-Our algorithm takes advantage of this property by exhaustively merging special paths until no edges remain between them.
+Intuitively, usual paths are “open-ended”: their endpoints are free, and since no edge ties them together, both endpoints can safely be included in an independent set.
+Usual paths are therefore useful for building independence, as each such path potentially contributes two vertices to the independent set.
+More formally, assume there are at least $k$ usual paths in $\mathcal{P}$ and that no edges connect the endpoints of two distinct paths $P_{i},P_{j}\in\mathcal{P}$.
+Then, if for each special path (of which there are at most $|\mathcal{P}|-k=p-k$) we take one endpoint, and for each usual path we take both endpoints, we obtain an independent set $I$ of size at least $(p-k)+2k\geq p+k$.
+The path cover $\mathcal{P}$ together with the independent set $I$ forms a certificate in Theorem 1 for the case when $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
+Thus, the first step of the proof of Theorem 1 is as follows.
+We start from a trivial path cover $\mathcal{P}$ consisting of single-vertex paths and perform “Gallai–Milgram mergings”.
+That is, we merge paths in $\mathcal{P}$ exhaustively as long as there exists a pair of paths with adjacent endpoints.
+When no such merge is possible, if at least $k$ paths are usual, the algorithm constructs $I$ and outputs a valid solution.
+In the remainder of the proof, the algorithm considers the case when in the constructed path cover $\mathcal{P}$:
+  • (i)
+    No two endpoints of distinct paths are adjacent, and
+  • (ii)
+    fewer than $k$ paths are usual, so that the remaining paths are special.
+Informally, special paths are “closed”: they either consist of a single vertex (a trivial path), or have two vertices, or are of length at least two and their endpoints are connected by an edge, so that the path together with this edge forms a cycle.
+In this situation, an edge between any two vertices of different special paths plays the same role as an edge between their endpoints: it allows us to merge the two paths and cover all their vertices with a single path.
+Therefore, while special paths are not very useful for constructing independent sets, they are extremely convenient for merging.
+Our algorithm exploits this property by exhaustively merging special paths until no edges remain between them.
+As a result, we obtain the following structural property of the path cover:
+  • (iii)
+    if two distinct paths $P_{i},P_{j}\in\mathcal{P}$ are special, then there is no edge between $V(P_{i})$ and $V(P_{j})$ in $G$.
+We perform a similar merging trick in the case when $P_{i}$ is special, $P_{j}$ is usual, and there is an edge from $V(P_{i})$ to an endpoint of $P_{j}$. This yields the fourth property:
+  • (iv)
+    If a path $P_{i}\in\mathcal{P}$ is special, then there is no edge between $V(P_{i})$ and any endpoint of any other path $P_{j}\in\mathcal{P}$.
+Together, properties (i)–(iv) provide additional structural insight into the path cover $\mathcal{P}$ and, consequently, into the graph $G$.
+Armed up with these properties, our algorithm is now able to construct a larger independent set in $G$ than before.
+For each usual path, the algorithm still takes its both endpoints into $I$.
+For each special path $P_{i}\in\mathcal{P}$, the algorithm can now take two arbitrary non-adjacent vertices from $V(P_{i})$ into $I$ provided that such vertices exist.
+If $V(P_{i})$ is instead a clique in $G$, the algorithm takes just a single arbitrary vertex into $I$.
+The resulting set $I$ is independent by properties (i), (iii), (iv) of $\mathcal{P}$.
+If at least $k$ special paths do not induce a clique in $G$, the size of $I$ is at least $p+k$, and this immediately resolves the problem.
+Thus, the only remaining scenario is that,
+  • (v)
+    except for at most $k$ special paths, the vertex set of each of the special paths is a clique in $G$.
 (a)  Refer to caption      (b)  Refer to caption      (c)  Refer to caption
@@
 $P_{1}$ and $P_{3}$ become usual paths, while $P_{2}$ can become special or even disappear completely.
-This property of special paths offers additional insight into the structure of $G$.
-Because there are no edges with endpoints in distinct special paths, and there are fewer than $k$ usual paths in $\mathcal{P}$, selecting one vertex from each special path already forms an independent set of size exceeding $p-k$.
-Moreover, if we choose two non-adjacent vertices from each of $2k$ special paths, we obtain an independent set of size at least $p+k$, immediately resolving the problem.
-Thus, the only remaining scenario is that, except for at most $2k$ special paths, vertex set of each of the special paths is a clique in $G$.
-However, this structural insight alone addresses only part of the problem.
-To progress further, we require a more sophisticated transformation of the paths in $\mathcal{P}$.
+However, this structural insight alone addresses only a part of the problem.
+To progress further, we need a more sophisticated transformation of the paths in $\mathcal{P}$.
 Rather than merging two paths, we now consider two special paths and one usual path such that there are two edges (with no shared endpoints) connecting the special paths to the usual path.
@@
 If all these transformations cannot be applied, we arrive at the following situation.
-There must exist a vertex set $S$ of size smaller than $k$ that separates all usual paths from at least $p-4k$ special paths, each of which induces a clique in $G$.
+There must exist a vertex set $S$ of size smaller than $k$ that separates all usual paths from at least p4kp-4k special paths, each of which induces a clique in $G$.
 In the auxiliary Lemma 1, we demonstrate how this small separator $S$ can be exploited to remove irrelevant cliques from $G$ without affecting the size of an optimal path cover or the independence number.
 Applying this polynomial-time procedure ensures that, in the end, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
-By applying transformations and irrelevant cliques removals, we end up in the situation, where $p=\mathcal{O}(k^{2})$.
+By applying transformations and irrelevant clique removals, we end up in the situation where $p=\mathcal{O}(k^{2})$.
 In this case, we invoke Theorem 2 to compute an optimal path cover of $G$, using $p+k$ as its parameter.
@@
 All steps of the algorithm are performed in polynomial time, except for the call to Theorem 2, resulting in running time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$.
-Overview of the proof of Theorem 2: We focus on outlining the main ideas of the proof specifically for $-\ell$-linkages.
+Overview of the proof of Theorem 2: We focus on outlining the main ideas of the proof specifically for Hamiltonian-$\ell$-linkages.
 The more general extension to arbitrary list TM-embeddings is similar but requires additional technical details.
-To begin, we sketch an $\operatorname{{\sf X}P}$ algorithm that solves $-\ell$-Linkage in time $n^{f(\alpha(G)+\ell)}$.
+To begin, we sketch an $\operatorname{{\sf X}P}$ algorithm that solves Hamiltonian-$\ell$-Linkage in time $n^{f(\alpha(G)+\ell)}$.
 Although this algorithm does not achieve the desired complexity, it is instructive for understanding the proof.
@@
 If its connectivity is low, we find a small vertex cut, delete it, and then recursively process the connected components of the resulting graph.
-Because a Hamiltonian path might traverse each component multiple times, we find it more convenient to work with $-\ell$-linkages, that is, linkages consisting of vertex-disjoint paths connecting specified pairs of vertices and covering all vertices of the graph.
-To handle such cases, one can strengthen Chvátal-Erdős theorem to apply to Hamiltonian-linkages.
-Following the foundational work of Bollobás and Thomason [BT96], there have been numerous advances in understanding how linkages relate to connectivity.
+Because a Hamiltonian path might traverse each component multiple times, we find it more convenient to work with Hamiltonian-$\ell$-linkages, that is, linkages consisting of vertex-disjoint paths connecting specified pairs of vertices and covering all vertices of the graph.
+To handle such cases, one can strengthen Chvátal-Erdős’s theorem to apply to Hamiltonian-linkages.
+Following the foundational work of Bollobás and Thomason [14], there have been numerous advances in understanding how linkages relate to connectivity.
 However, our setting additionally requires that the linkages are Hamiltonian.
-While no existing results directly address this requirement, we adapt the result of Thomas and Wollan [TW05] on linkages in highly connected graphs to show that if a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked.
+While no existing results directly address this requirement, we adapt the result of Thomas and Wollan [87] on linkages in highly connected graphs to show that if a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is Hamiltonian-$\ell$-linked.
 In other words, there exists a path cover of size $\ell$ that links any prescribed set of endpoints.
 The algorithm proceeds as follows.
-If the input graph $G$ is $\Omega(k\cdot\ell)$-connected, then the algorithm correctly reports that $G$ is $-\ell$-linked.
+If the input graph $G$ is $\Omega(k\cdot\ell)$-connected, then the algorithm correctly reports that $G$ is Hamiltonian-$\ell$-linked.
 Otherwise, if the connectivity is lower, there exists a minimum vertex separator $X$ of $G$ with $\mathcal{O}(k\cdot\ell)$ vertices.
-Each connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component.
+Each connected component of GXG-X has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component.
 For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$.
-For every connected component of $G-X$ and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component.
+For every connected component of GXG-X and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component.
 Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices.
@@
 The recursive understanding technique introduced by Chitnis et al.
-[CCH+16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al.
-[LRSZ18].
+[17] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al.
+[75].
 Unfortunately, recursive understanding does not appear to be particularly helpful in our setting for the following reason: in order to apply recursive understanding, one must handle the so-called $(p,q)$-unbreakable graphs as the “base case.” However, for our problem, addressing $(p,q)$-unbreakable graphs is not simpler than handling the general case.
@@
 To prevent the $\operatorname{{\sf X}P}$ running time, the $\operatorname{{\sf FPT}}$ algorithm avoids recursive calls to itself.
-Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either output an independent set of $G$ of size $k$ or find a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of $G-S$ is $\Omega(k\cdot(\ell+|S|))$-connected.
+Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either outputs an independent set of $G$ of size $k$ or finds a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of GSG-S is $\Omega(k\cdot(\ell+|S|))$-connected.
 Set $S$ breaks solution paths into at most $\ell+2|S|$ parts.
-Some parts belong to $G[S]$, while others belong to $G-S$.
+Some parts belong to $G[S]$, while others belong to GSG-S.
 Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time.
@@
 Here is precisely the moment when the property of the decomposition comes into play.
-Each connected component of $G-S$ is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked.
+Each connected component of GSG-S is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked.
 Therefore, for any selection of the paths’ endpoints, the required collection of vertex-disjoint paths spanning all component vertices always exists (unless empty).
-Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of $G-S$ without enumerating the entry-exit endpoints in $G-S$.
+Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of GSG-S without enumerating the entry-exit endpoints in GSG-S.
 The latter task reduces to computing a maximum matching in an auxiliary graph.
@@
 Specifically, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we must be able to construct $\ell$ disjoint paths connecting these endpoints and covering all vertices of the graph, or else find an independent set of size $k$.
-For this purpose, we rely on a combinatorial result by Thomas and Wollan [TW05], which states that in a $10\ell$-connected graph, one can always find $\ell$ vertex-disjoint paths connecting any given pairs of endpoints.
-However, the proof in [TW05] is non-constructive—it builds on certain properties of a minimal counterexample, and it is not obvious how to transform that argument into an algorithm that yields the required paths.
+For this purpose, we rely on a combinatorial result by Thomas and Wollan [87], which states that in a $10\ell$-connected graph, one can always find $\ell$ vertex-disjoint paths connecting any given pairs of endpoints.
+However, the proof in [87] is non-constructive—it builds on certain properties of a minimal counterexample, and it is not obvious how to transform that argument into an algorithm that yields the required paths.
 To circumvent this issue, we make use of the fact that our graphs have bounded independence number.
@@
 In the latter scenario, we can apply Menger’s theorem to connect the endpoints to the clique, thereby constructing the required $\ell$ paths.
-As a side remark, several combinatorial results—including the aforementioned work of Thomas and Wollan [TW05]—guarantee the existence of $\ell$ vertex-disjoint paths between specified pairs of endpoints in a graph with connectivity $\Omega(\ell)$.
+As a side remark, several combinatorial results—including the aforementioned work of Thomas and Wollan [87]—guarantee the existence of $\ell$ vertex-disjoint paths between specified pairs of endpoints in a graph with connectivity $\Omega(\ell)$.
 However, there is a notable shortage of simple and efficient algorithms to compute such vertex-disjoint paths in $\Omega(\ell)$-connected (or even $\ell^{\Omega(1)}$-connected) graphs.
 Our contribution in this direction may therefore be of independent interest.
-Finally, for Theorem 2, we additionally generalize the argument above from $-\ell$-linkages to arbitrary list TM-embeddings.
-Here, once more, the non-recursive structure of the algorithm proves helpful: With the same choice of the set $S$, we enumerate all possible behaviors of the target model on $S$ and between $S$ and the connected components of $G-S$.
-Again, we avoid enumerating concrete terminals and connecting vertices in $G-S$, fixing only the general connection pattern to $S$.
+Finally, for Theorem 2, we additionally generalize the argument above from Hamiltonian-$\ell$-linkages to arbitrary list TM-embeddings.
+Here, once more, the non-recursive structure of the algorithm proves helpful: With the same choice of the set $S$, we enumerate all possible behaviors of the target model on $S$ and between $S$ and the connected components of GSG-S.
+Again, we avoid enumerating concrete terminals and connecting vertices in GSG-S, fixing only the general connection pattern to $S$.
 While considerably more technical, the enumeration can still be done in $\operatorname{{\sf FPT}}$ time.
-Then, similarly to $-\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly-connecting component (or an independent set of size $k$), with arbitrary locations of the terminals.
-1.3 Related Work.
-The interplay among the independence number, connectivity, Hamiltonicity, and path covers has long been a popular theme in graph theory, with historical roots going back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72].
-Surveys by Jackson and Ordaz [JO90], Bondy [Bon95], and Kawarabayashi [Kaw01] offer excellent overviews of this area.
+Then, similarly to Hamiltonian-$\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly connected component (or an independent set of size $k$), with arbitrary locations of the terminals.
+1.3 Related Work
+The interplay among the independence number, connectivity, Hamiltonicity, and path covers has long been a popular theme in graph theory, with historical roots going back to the works of Dilworth [29], Gallai and Milgram [45], Nash-Williams [82], and Chvátal and Erdős [19].
+Surveys by Jackson and Ordaz [59], Bondy [15], and Kawarabayashi [65] offer excellent overviews of this area.
 However, most of the research thus far has been centered on establishing combinatorial conditions for Hamiltonicity (or related properties), rather than on algorithmic extensions.
 Theorem 1 contributes to a growing line of parameterized algorithms that extend classic theorems from extremal graph theory.
-Examples include algorithmic adaptations of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem [GHH23].
+Examples include algorithmic adaptations of Poljak–Turzík’s theorem [80], Dirac’s theorem [37], and Hajnal–Szemerédi’s theorem [46].
 Such algorithmic extensions can be seen as part of a broader trend in parameterized complexity called above (or below) guarantee parameterization, where one observes that: the natural parameterization of a maximization or minimization problem by the solution size is not satisfactory if there is a lower bound for the solution size that is sufficiently large.
-The term above guarantee parameterization was introduced by Mahajan and Raman [MRS09], and it has proved fruitful for studying numerous fundamental problems in parameterized complexity and kernelization [AGK+10, BCDF19, CJM+13, FGL+21, FGSS22b, GP16a, GKLM11, GvIMY12, GP16b, GRSY07, LNR+14, GM22].
-A feature distinguishing Theorem 1 from the previous work is that it provides “tractability below intractable parameter”.
+The term above guarantee parameterization was introduced by Mahajan and Raman [77], and it has proved fruitful for studying numerous fundamental problems in parameterized complexity and kernelization [1, 9, 20, 34, 38, 39, 49, 52, 56, 54, 55, 74, 53].
+A feature distinguishing Theorem 1 from the previous work is that it provides “tractability below an intractable parameter”.
 The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms.
-Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved.
-Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16].
-This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, FLP+20].
-A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph [CNP+11, BCKN13, CKN13, FLPS16, BST23], see also [CFK+15].
+Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [2], substantial progress in algorithmic methodologies has been achieved.
+Notably, a range of powerful algorithmic methods has been developed, including those by Björklund [10], Koutis [70], Williams [90], and Zehavi [91].
+This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [51, 41].
+A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph [23, 13, 22, 40, 4], see also [21].
 Such parameters include treewidth, treedepth, or minimum vertex cover size.
 On the other hand, for more “general” parameters Hamiltonicity becomes intractable.
-For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10] and is paraNP parameterized by $\alpha$-treewidth [DMS24].
-The existence of a Hamiltonian cycle remains NP-complete on planar graphs [GJT76], circle graphs [Dam89] and for several other generalizations of interval graphs [BB86], for chordal bipartite graphs [Mül96] and for split (and therefore also for chordal) graphs [Gol04], but is solvable in linear time in interval graphs [Kei85] and in convex bipartite graphs [Mül96].
-The existence of a Hamiltonian path can be decided in polynomial time for cocomparability graphs [DDKS91] and for circular arc graphs [Dam93].
+For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [35] and is $\operatorname{{\sf Para-NP}}$-hard parameterized by $\alpha$-treewidth (see [24] for the definition) as the problem is $\operatorname{{\sf NP}}$-complete on split graphs [50].
+The existence of a Hamiltonian cycle remains NP-complete on planar graphs [48], circle graphs [25] and for several other generalizations of interval graphs [7], for chordal bipartite graphs [81] and for split (and therefore also for chordal) graphs [50], but is solvable in linear time in interval graphs [66] and in convex bipartite graphs [81].
+The existence of a Hamiltonian path can be decided in polynomial time for cocomparability graphs [27] and for circular arc graphs [26].
 Many of the above mentioned graph classes are hereditary, i.e., closed in the induced subgraph order, and as such can be described by collections of forbidden induced subgraphs.
@@
 For many graphs $H$, the class of $H$-free graphs has nice structural properties (e.g., $P_{4}$-free graphs are exactly the cographs) or their structural properties can be used to design polynomial algorithms for various graph theory problems (as a recent example, cf.
-colouring of $P_{6}$-free graphs in [CSZ19]).
+colouring of $P_{6}$-free graphs in [18]).
 For Hamiltonian-type problems, consider three-vertex graphs $H$.
-For $H=K_{3}$, both Hamiltonian cycle and Hamiltonian path are NP-complete on triangle-free graphs, since they are NP-complete on bipartite graphs [Mül96].
-For $H=3K_{1}$, the edgeless graph on three vertices, both Hamiltonian cycle and Hamiltonian path are polynomial time solvable [DGJ81].
-The remaining two graphs, $P_{3}$ and $K_{1}+K_{2}$, are induced subgraphs of $P_{4}$, thus the corresponding class of $H$-free graphs is a subclass of cographs, and as such a subclass of cocomparability graphs, in which the Hamiltonian path and Hamiltonian cycle problems are solvable in polynomial time [DDKS91].
+For $H=K_{3}$, both Hamiltonian cycle and Hamiltonian path are NP-complete on triangle-free graphs, since they are NP-complete on bipartite graphs [81].
+For $H=3K_{1}$, the edgeless graph on three vertices, both Hamiltonian cycle and Hamiltonian path are polynomial time solvable [31].
+The remaining two graphs, $P_{3}$ and $K_{1}+K_{2}$, are induced subgraphs of $P_{4}$; thus, the corresponding class of $H$-free graphs is a subclass of cographs, and as such a subclass of cocomparability graphs, in which the Hamiltonian path and Hamiltonian cycle problems are solvable in polynomial time [27].
 However, a complete characterization of graphs $H$ for which Hamiltonian path or Hamiltonian cycle problems are solvable in polynomial time (and for which they are NP-complete) is not in sight.
@@
 Theorem 2 answers this complexity question for all edgeless forbidden induced subgraphs.
-In a companion paper Jedličková and Kratochvíl [JK24] explicitly described situations in which a Hamiltonian path for $\alpha(G)=3,4$ and $5$ in which a Hamiltonian path does exist.
+In a companion paper Jedličková and Kratochvíl [62] explicitly described situations for $\alpha(G)=3,4$ and $5$ in which a Hamiltonian path does exist.
 The parameterized complexity of many graph problems remains insufficiently explored for parameters capturing a graph’s density—such as its independence number.
@@
 Only a handful of results ($\operatorname{{\sf FPT}}$ or $\operatorname{{\sf X}P}$) address parameterization by a graph’s independence number.
-These include the $\operatorname{{\sf X}P}$ algorithm of Ovetsky Fradkin and Seymour [FS15] for the edge-disjoint paths problem in directed graphs, the kernelization algorithms of Lochet et al.
-[LLM+20] for various cycle-hitting problems in directed graphs, the subexponential parameterized algorithms of Misra et al.
-[MSSZ23] for digraphs with constant independence number, and the recent work of Maalouly, Steiner, and Wulf [MSW23] on $\operatorname{{\sf FPT}}$ algorithms for exact matching.
+These include the $\operatorname{{\sf X}P}$ algorithm of Ovetsky Fradkin and Seymour [44] for the edge-disjoint paths problem in directed graphs, the kernelization algorithms of Lochet et al.
+[72] for various cycle-hitting problems in directed graphs, the subexponential parameterized algorithms of Misra et al.
+[79] for digraphs with constant independence number, and the recent work of Maalouly, Steiner, and Wulf [76] on $\operatorname{{\sf FPT}}$ algorithms for exact matching.
 2 Preliminaries
@@
 The subgraph of $G$ induced by vertices $S\subseteq V(G)$ will be denoted by $G[S]$.
-We use $G-S$ for $G[V(G)\setminus S]$.
+We use GSG-S for $G[V(G)\setminus S]$.
 The independence number of a graph $G$, denoted by $\alpha(G)$, is the order of the largest edgeless induced subgraph of $G$.
@@
 A cycle is formed by a path of length greater than 1 which connects two adjacent vertices.
-The path (cycle) is Hamiltonian if it contains all vertices of the graph.
+The path (cycle) is Hamiltonian if it contains all the vertices of the graph.
 The path cover number of $G$, denoted by $\mathrm{pc}(G)$, is the smallest number of vertex disjoint paths that cover all vertices of $G$.
@@
 A vertex cut in a graph $G$ is a set $A\subset V(G)$ of vertices such that the graph $G-A=G[V(G)\setminus A]$ is disconnected.
-The vertex connectivity $c_{v}(G)$ of a graph $G$ is the order of a minimum cut in $G$, or $|V(G)|-1$ if $G$ is a complete graph.
+The vertex connectivity $c_{v}(G)$ of a graph $G$ is the order of a minimum cut in $G$, or |V(G)|1|V(G)|-1 if $G$ is a complete graph.
 Since we will not consider edge connectivity, we will often omit the adjective when talking about the connectivity measure, we always have vertex connectivity in mind.
@@
 This is important when creating a longer path by concatenating shorter ones.
-We build upon the following results of Chvátal and Erdős which we will later generalize to Hamiltonian linkages in Theorem 2.
-Proposition 1.
-[CE72] Let $G$ be an $s$-connected graph.
+We build upon the following results of Chvátal and Erdős, which we will later generalize to Hamiltonian linkages in Theorem 2.
+Proposition 1 (Chvátal [19]).
+Let $G$ be an $s$-connected graph.
   1. 1.
@@
     If $\alpha(G)<s$, then $G$ is Hamiltonian connected (i.e., every pair of vertices is joined by a Hamiltonian path).
-3 Proof of Theorem 1: Path cover bellow Gallai-Milgram
+3 Proof of Theorem 1: Path cover below Gallai-Milgram
 In this section, we prove Theorem 1.
@@
 Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$.
-A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$.
+A well-known theorem by Gallai and Milgram [45] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$.
 When we combine this theorem with Theorem 2, we obtain an algorithm solving Path Cover in time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$.
@@
 The first result is dealing with small vertex separators of $G$.
-It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
+It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in GSG-S and the value of $\operatorname{pc}(G)$ simultaneously.
 Lemma 1.
-Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of $G-S$.
+Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of GSG-S.
 Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
@@
 Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$.
-To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$.
-On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching.
+To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 0 to i1i-1.
+On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and iji-j copies of $K_{2}$, that is, $H_{ij}$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching.
 Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial.
@@
 The subroutine works in
-$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}}\cdot 2^{(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
+$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}\cdot(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
 running time.
@@
 Then $\mathcal{A}$ outputs this optimal path cover of $G$ and stops.
-Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$.
+Note that $\mathcal{A}$ necessarily stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$.
 The proof of the lemma is complete.
@@
 That is, the algorithm applies a reduction rule or solution subroutine only if no reduction rule or solution subroutine before it can be applied.
-When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties.
+When we reach the situation where the obtained path cover $\mathcal{P}$ and graph $G$ are irreducible, we deduce that the instance has nice structural properties.
 In this case, the final solution subroutine calls the algorithms of Lemmas 1 and 2.
-The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 00.
+The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 0.
 Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$.
@@
 We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
-The first reduction rule is straightforward and comes from the proof of the Gallai-Milgram theorem for undirected graphs.
+The first reduction rule is straightforward and comes from the proof of Gallai-Milgram’s theorem for undirected graphs.
 Reduction Rule 1.
@@
 The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one.
-Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$.
+Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{m}\}$ induces an independent set in $G$.
 When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
@@
 Proof of Claim 3.
-Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$.
+Assume that the reduction rules are not applicable, but there are special paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$.
 If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable.
@@
 We construct an independent set $I$ consisting of more than $m+k$ vertices.
-Since 1 is not applicable, there are more than $m-k$ special paths.
+Since 1 is not applicable, there are more than mkm-k special paths.
 For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$.
 For each other special path, put its arbitrary vertex into $I$.
-By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎
+By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, |I|>(mk)+2k=m+k|I|>(m-k)+2k=m+k. ∎
 The solution subroutine based on 4 is given below. The algorithm applies it when none of the reduction rules and solution subroutines above are applicable.
@@
 The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$.
-The following reduction rule necessary reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
+The following reduction rule reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
 a $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$ b $s_{i}$$t_{i}$$s_{j}$$t_{j}$$u_{i}$$u_{j}$$v_{i}$$v_{j}$$s_{\ell}$$t_{\ell}$$x_{i}$$x_{j}$
-Figure 2: Illustration of transformations performed in (a) 3 and (b) 4..
+Figure 2: Illustration of transformations performed in (a) 3 and (b) 4.
 Paths before the transformations are horizontal.
@@
 Reduction Rule 4.
-If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$ and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
+If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$, $x_{i}\neq x_{j}$, and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
   1. 1.
@@
 We first show that $|V(Q_{i})|\geq 3$.
-Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$.
+Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subseteq V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$.
 The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly.
 Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold.
-But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$.
+But $u_{i}=v_{i}$ implies that $P_{i}$ has length 0 and $s_{i}=t_{i}=u_{i}$.
 Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$.
@@
 It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$.
-Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable.
+Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$ implying that 2 is applicable.
 This contradiction proves the claim.
@@
 Claim 6.
-If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$.
+If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in GSG-S for at least m2km-2k special paths $P_{i}$.
 Moreover, such a set $S$ could be constructed in polynomial time.
@@
 when $P_{i}$ has zero or more than one connectors, we do not put any vertex of $P_{i}$ in $S$.
-We show that at least $m-2k$ special paths are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded.
-Assume now that $P_{j}$ is a special path that is not isolated in $G-S$.
+We show that at least m2km-2k special paths are isolated in GSG-S. First note that there are at least mkm-k special paths in $\mathcal{P}$, since the number of usual paths is bounded.
+Assume now that $P_{j}$ is a special path that is not isolated in GSG-S.
 Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$.
@@
 As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$.
-Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$.
-Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$.
+Hence, each usual path $P_{i}$ gives at most one non-isolated special path in GSG-S.
+Thus, at most $k$ special paths in GSG-S are not isolated, implying that at least (mk)k=m2k(m-k)-k=m-2k special paths are isolated in GSG-S.
 ∎
@@
   2. 2.
-    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in $G-S$;
+    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in GSG-S;
   3. 3.
@@
 Proof of Claim 7.
-First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 6 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
+First note that at least m4km-4k complete connected components are guaranteed in GSG-S by 6 ($S$ separates m2km-2k special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
 We now move on to establish the upper bound on the running time. All steps, except for applying Lemma 2, are carried out in polynomial time. We have
@@
 Hence it is also a path cover.
-On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$.
+On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in GSG-S.
 Now we prove that either $\mathcal{S}$ is optimal or $|I|-|\mathcal{S}|\geq k$.
@@
 As discussed in the previous paragraph, $|\mathcal{S}|-|\mathcal{S}^{\prime}|=h-|X|$. We conclude that
-$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)+(h-|X|)=k.$
+$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)-(h-|X|)=k.$
 The feasibility of $\mathcal{S}$ and $I$ is proved. ∎
@@
 The proof consists of three main steps, deferred to their respective subsections.
-The first step is the proof of the stronger version of the theorem for highly connected graphs which we call Spanning Lemma.
-It shows that for any fixed injection $f:V(H)\to V(G)$ the desired TM-embedding respecting $f$ or an independent set can be found; moreover, the found TM-embedding spans $V(G)$.
-The second step is the so-called Merging Lemma, that combines the solutions from several highly-connected components, provided that the size of the separator between them is small.
-Finally, to obtain the proof of Theorem 2 itself, we show an iterative procedure that either finds the desired separator, or an independent set.
+The first step is the proof of the stronger version of the theorem for highly connected graphs, which we call Spanning Lemma.
+It shows that for any fixed injection $f:V(H)\to V(G)$, the desired TM-embedding respecting $f$ or an independent set can be found; moreover, the found TM-embedding spans $V(G)$.
+The second step is the so-called Merging Lemma, which combines the solutions from several highly-connected components, provided that the size of the separator between them is small.
+Finally, to obtain the proof of Theorem 2 itself, we show an iterative procedure that either finds the desired separator or an independent set.
 4.1 Spanning Lemma
@@
 Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$.
-A graph is called $-\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
+A graph is called Hamiltonian-$\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
 We need a procedure constructing the linkage for computing a TM-embedding of $H$.
-For that we need a constructive version of the following theorem of Thomas and Wollan [TW05], which proof is non-constructive.
-Proposition 2 (Corollary 1.3, [TW05]).
+For that we need a constructive version of the following theorem of Thomas and Wollan [87] whose original proof is non-constructive.
+Proposition 2 (Corollary 1.3, [87]).
 If $G$ is $10\ell$-connected then $G$ is $\ell$-linked.
@@
 For simplicity, our algorithm first ensures that $s_{1},t_{1},s_{2},t_{2},\ldots,s_{\ell},t_{\ell}$ are $2\ell$ distinct vertices, i.e., $|T|=2\ell$.
-For each vertex $v\in T$ that appears in $p>1$ input pairs, the algorithm replaces $v$ by $p$ true twins, i.e., vertices that contain the neighborhood of $v$ and are also adjacent to each other.
+For each vertex $v\in T$ that appears in $p>1$ input pairs, the algorithm replaces $v$ by $p$ true twins, i.e., by vertices with the same neighborhood as $v$ that are also adjacent to each other.
 Then, each repeating occurrence of $v$ in the input pairs is replaced by a distinct copy.
@@
 Let $C$ be a given clique of size $2\ell$ in $G$.
-Then a family of internally vertex-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$, where for each $i\in[\ell]$, $P_{i}$ connects $s_{i}$ with $t_{i}$, can be constructed in polynomial time.
+Then a family of vertex-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$, where for each $i\in[\ell]$, $P_{i}$ connects $s_{i}$ with $t_{i}$, can be constructed in polynomial time.
 Proof of Claim 8.
-Since $G$ is $2\ell$-connected, by Menger’s theorem [Men27, Die17], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$.
-These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [FF56] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time.
+Since $G$ is $2\ell$-connected, by Menger’s theorem [78, 28], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$.
+These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [42] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time.
 Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$.
@@
 It is left to explain how the algorithm constructs the clique $C$, given that $G$ has at least $(k+2\ell)^{k}$ vertices.
-The essential tool here is the classical Ramsey’s theorem [Ram30], which guarantees a large clique or independent set in a large enough graph.
-For completeness, we present the corresponding polynomial-time subroutine based on the Ramsey number bound given by Erdős and Szekeres in 1935 [ES09].
+The essential tool here is the classical Ramsey’s theorem [84], which guarantees a large clique or independent set in a large enough graph.
+For completeness, we present the corresponding folklore polynomial-time subroutine based on the Ramsey number bound given by Erdős and Szekeres in 1935 [32].
 Claim 10.
@@
 Proof of Claim 10.
-Note that $G$ necessary has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [ES09].
+Note that $G$ necessarily has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [32].
 The algorithm is recursive.
-Since $r,s>0$, $|V(G)|>1$ and $G$ has at least one vertex, denote it by $v$.
+Since $r,s>0$, $|V(G)|\geq 1$ and $G$ has at least one vertex, denote it by $v$.
 If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops.
@@
 Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$.
-If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size $s-1$.
-In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and $s-1>0$.
-If its output is a clique of size $s-1$, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$.
+If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size s1s-1.
+In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and s1>0s-1>0.
+If its output is a clique of size s1s-1, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$.
 Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required.
@@
 From now on, every vertex of $H$ has at least one incident edge, while $G$ is ($\max\{k+2,10\}\cdot|E(H)|$)-connected.
-In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ vertex pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$.
+In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$.
 To see that the reformulation is equivalent, put $\ell=|E(H)|$ and $(s_{i},t_{i})=(f(x_{i}),f(y_{i}))$, where $x_{i},y_{i}$ are the endpoints of the $i$-th edge of $H$.
@@
 If we show how this can be done exhaustively and in polynomial time, it is clearly sufficient to find the desired TM-embedding.
-We describe the enlargement process, akin to the method outlined in [JK23], with a key distinction.
-Unlike the scenario in [JK23], we lack assurance that $G$ satisfies $\alpha(G)<k$, making the procedure potentially prone to failure.
+We describe the enlargement process, akin to the method outlined in [61], with a key distinction.
+Unlike the scenario in [61], we lack assurance that $G$ satisfies $\alpha(G)<k$, making the procedure potentially prone to failure.
 Therefore, we ensure that (a) the enlargement is completed within polynomial time and (b) if the enlargement fails, an independent set of size $k$ is identified in $G$.
@@
 Each path goes from $x$ to $S$, and no two paths share any common vertex (including endpoints in $S$) but $x$.
-One can find such paths by utilizing the Ford-Fulkerson algorithm [FF56] to find a flow of size at most $(k+2)\ell$ in $\mathcal{O}(k\ell\cdot|G|)$ running time.
+One can find such paths by utilizing the Ford-Fulkerson algorithm [42] to find a flow of size at most $(k+2)\ell$ in $\mathcal{O}(k\ell\cdot|G|)$ running time.
 If $|S|\leq(k+2)\ell$, then two of these paths end in $s_{1}$ and $v$, where $v$ is the neighbor of $s_{1}$ in $P_{1}$.
@@
 Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$.
-The algorithm takes arbitrary $k$ of them.
+The algorithm arbitrary takes $k$ such paths.
 Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$.
@@
 The correctness of the algorithm follows from the discussion. The proof of the lemma is complete. ∎
+Finally, we note that Lemma 4 admits an existential variant as a direct corollary: If $G$ is sufficiently well-connected compared to its independence number $\alpha(G)$ and the size of $H$, then a spanning TM-embedding of $H$ exists in $G$ for any choice of $f$.
+As this may be of independent interest, we state formally the following.
+Proposition 3.
+Let $H$ and $G$ be non-empty graphs, and let $f\colon V(H)\rightarrow V(G)$ be an injective mapping.
+Let $G$ additionally be $(\max\{\alpha(G)+3,10\}\cdot h)$-connected, where $h=|V(H)|+|E(H)|$.
+There exists a subgraph $M$ of $G$ such that $(M,f)$ is a TM-embedding of $H$ in $G$ spanning $V(G)$.
 4.2 Merging Lemma
@@
 Lemma 5 (Merging Lemma).
-Let $G$ be a graph, let $S\subseteq V(G)$ be a vertex subset in $G$, and let $C_{1}$, …, $C_{t}$ be the connected components of $G-S$.
+Let $G$ be a graph, let $S\subseteq V(G)$ be a vertex subset in $G$, and let $C_{1}$, …, $C_{t}$ be the connected components of GSG-S.
 Let $H$ be a graph, let $k$ be a parameter, and $L:V(H)\to 2^{V(G)}$ be a list assignment.
@@
 Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$.
-Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)\in C_{i}\}$.
+Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)=i\}$.
 We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@
 From CD2, every connector in $w$ is adjacent to a vertex of $S$ in $M_{W}$.
-Since $f_{W}$ is injective, number of terminals in $W$ is at most $|V(H)|$, therefore it remains to bound the number of connectors in $W$ by $|E(H)|+2s$.
+Since $f_{W}$ is injective, the number of terminals in $W$ is at most $|V(H)|$, therefore it remains to bound the number of connectors in $W$ by $|E(H)|+2s$.
 Let $w\in W$ be a connector that is adjacent to a terminal in $S$, the number of such connectors is at most $\sum_{s\in S\cap T_{W}}|N_{M_{W}}(s)\cap W|\leq|E(H)|$, since by CD1 every $s\in S\cap T_{W}$ has degree $\deg_{H}(h)$, where $h$ is the unique pre-image of $s$ with respect to $f_{W}$, and no edge of $H$ is counted twice.
@@
 First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$.
-We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$.
+We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$; here the dissolving a vertex $x$ of degree two with the neighbors $y$ and $z$ deletes $x$ and adds the edge $yz$.
 We set $f_{W}$ as the respective mapping of the terminals obtained from $f$.
@@
 Moreover, $f\equiv f_{W}$ on $S$; let $s$ be such that $f(h)=f_{W}(h)=s$, then $s\in L(h)$ since $(M,f)$ is a list TM-embedding respecting $L$.
-Finally, for CD5, the assigment $\tau$ is given directly by construction as an identity assignment on $W\subset V(G)$.
+Finally, for CD5, the assignment $\tau$ is given directly by construction as an identity assignment on $W\subset V(G)$.
 For each $w\in W$, $\tau(w)\in C_{\xi(w)}$ by the construction of $\xi$.
@@
 1
-$\mathcal{M}\leftarrow\emptyset$;
+[…]
 3for $|W|\leftarrow 0$ to $h+2s$ do
@@
 22       else
-23                         $M_{i}\leftarrow O$; $M\leftarrow M$ where $M[T_{i}]$ is replaced by $M_{i}$;
+23          $_{i}\leftarrow O; MM$$\leftarrow M$ where $M[T_{i}]$ is replaced by $M_{i}$;
 24          foreach $w\in(T_{i}\cap\operatorname{Im}f_{W})$ do
@@
 This holds since $|V(H_{i})|+|E(H_{i})|\leq 3h+3s$, see the proof of 14.
-Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $w^{t}$ choices for $\xi$, and at most $|V(H)|^{s+w}$ choices for $f_{W}$.
+Moving to Algorithm 1, let $w=h+2s$ be the maximum size of $W$; there are at most $t^{w}$ choices for $\xi$, and at most ${(s+w)}^{|V(H)|}$ choices for $f_{W}$.
 We may assume $t\leq k$, since otherwise, an independent set of size $k$ can be found trivially by taking an arbitrary vertex from each $C_{i}$.
@@
 Since every non-terminal has degree two in $M_{W}$, there are at most $(s+w)^{2}$ choices for each non-terminal, and at most $(s+w)^{2(s+w)}$ choices in total.
-This part can also be upper-bounded by $(s+h)^{\mathcal{O}(s+h)}=2^{(s+h)^{1+o(1)}}$.
+In total, Algorithm 1 performs at most $(s+h+k)^{\mathcal{O}(s+h)}=2^{(s+h+k)^{1+o(1)}}$ iterations.
 The overall running time is therefore at most $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$.
@@
 6    $I\leftarrow\{v_{1},\ldots,v_{k}\}$, where $v_{i}\in C_{i}$;
-7            return $I$.
+7    return II.
 8   end if
@@
 We iteratively construct the vertex set $S$ suitable for Lemma 5, then invoke Lemma 5.
-We start by setting $S=\emptyset$, then iteratively find a separator in one of the connected components of $G-S$, and add it to $S$.
-This happens until the connected components of $G-S$ are all $\kappa$-connected, where $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$.
+We start by setting $S=\emptyset$, then iteratively find a separator in one of the connected components of GSG-S, and add it to $S$.
+This happens until the connected components of GSG-S are all $\kappa$-connected, where $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$.
 If, at some point, there are at least $k$ connected components, we return the independent set constructed by taking a vertex from each connected component.
@@
 It is left to verify that the call to Lemma 5 is valid.
-For that, we need that every connected component of $G-S$ is $\kappa$-connected, for $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$.
+For that, we need that every connected component of GSG-S is $\kappa$-connected, for $\kappa=3\cdot\max\{k+2,10\}\cdot(|V(H)|+|E(H)|+|S|)$.
 However, this is exactly the condition at Algorithm 2, and this is the only option for the algorithm to break out of the main loop while reaching Algorithm 2.
@@
 First, we observe that the main loop performs at most $k$ iterations before it stops either at Algorithm 2 or at Algorithm 2.
-Specifically, at the start of the $j$-th iteration, the number of connected components in $G-S$ is at least $j$, which implies that at iteration $k$, the loop necessarily stops at Algorithm 2.
+Specifically, at the start of the $j$-th iteration, the number of connected components in GSG-S is at least $j$, which implies that at iteration $k$, the loop necessarily stops at Algorithm 2.
 This holds since each iteration increases the number of connected components by at least one, since a separator of one of the former components is added into $S$, while the remaining components are unchanged.
@@
 We now upper-bound the running time of Algorithm 1 by the running time bound of Lemma 5, since clearly up to Algorithm 2 the algorithm takes polynomial time.
-By Lemma 5, the total running time is thus at most $2^{(hk^{\mathcal{O}(k)})^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}=2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$.
+By Lemma 5, the total running time is thus at most $2^{(hk^{\mathcal{O}(k)})^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}=2^{|H|^{\mathcal{O}(k)}\cdot k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$.
 ∎
-5 Conclusion
-In conclusion, this paper shows that Hamiltonian Path/Cycle and Path Cover are $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$.
-Furthermore, Path Cover is $\operatorname{{\sf FPT}}$ parameterized “below” the Gallai–Milgram bound.
+5 Open questions
 We conclude with several open research questions.
@@
 But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$.
-Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [CNP+11, FGL+19, FGLS10, LMS18].
+Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [23, 36, 35, 73].
 However, all these reductions use gadgets with large independence numbers.
@@
 The “simplest” question here is whether there is a polynomial-time algorithm for computing a path cover of size $\alpha(G)-1$?
+Acknowledgements
+Fedor V.
+Fomin and Petr A.
+Golovach were supported by the Research Council of Norway via the BWCA project (grant no.
+314528), Fedor V.
+Fomin was supported by the European Research Council (ERC) via grant NewPC (reference 101199930).
+[Uncaptioned image]
+Petr A. Golovach was supported by the Research Council of Norway via the Extreme-Algorithms project (grant no 355137).
+Nikola Jedličková and Jan Kratochvíl were supported by GAUK 370122.
+In the later stages of preparing this article, Nikola Jedličková was supported by the project PRIMUS/24/SCI/008 and Charles University Research Centre program No.
+UNCE/24/SCI/022.
+Danil Sagunov was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement 075-15-2025-344 dated 29/04/2025 for Saint Petersburg Leonhard Euler International Mathematical Institute at PDMI RAS).